### v1

Abstract
Effective code optimization in compilers plays a central role in computer and software engineering. While compilers can be made to automatically search the optimization space without the need for user interventions, this is not a standard practice since the search is slow and cumbersome. Here we present CodeZero, an artificial intelligence agent trained extensively on large data to produce effective optimization strategies instantly for each program in a single trial of the agent. To overcome the huge range of possible test programs, we prepare a large dataset of training programs that emphasize quality, naturalness, and diversity. To tackle the vast space of possible optimizations, we adapt deep reinforcement learning to train the agent in a sample-efficient manner through interacting with a world model of the compiler environment. Evaluation on both benchmark suites and production-level code optimization problems demonstrates our agent’s supercompiler performances and zero-shot generalization abilities, outperforming built-in optimization options designed by compiler experts. Our methodology kindles the great potential of artificial intelligence for engineering and paves the way for scaling machine learning techniques in the realm of code optimization.
1 Introduction
Entering the post-Moore’s law era, code optimization is crucial for computer and software engineering, which plays an important role in realizing the full potential of slow-growing hardware. Developers typically rely on a compiler’s ability to transform input programs into semantically equivalent but more efficient versions, improving metrics like execution time, code size, and power consumption. For example, standard optimization options -O1, -O2, and -O3 aim to reduce execution time, while the -Os and -Oz options are crafted to reduce code size. Still, it is not common for users to explore beyond these conventional compiler options. Given the vast diversity of programs and platforms, coupled with the increasing number of optimization passes integrated into compiler frameworks, these off-the-shelf optimization strategies predefined heuristically by compiler experts may struggle to guarantee near-optimal performance across ever-changing scenarios triantafyllis2003compiler ; georgiou2018less .
Automatic code optimization is therefore crucial in compilers. Autotuning bodin1998iterative improves code by systematically searching the optimization space through iterative executing and profiling optimization strategies. This search technique can yield remarkable performance gains but must be rerun for each new program with thousands of compilations, which is too time-consuming to be practical for all but a few specialized use cases. Meanwhile, machine learning techniques hold the capabilities to generalize the optimization strategy of one program to other similar ones, thereby facilitating faster code optimization. A direct method is to use supervised learning to predict good optimization strategies of input programs calder1997evidence ; fursin2008milepost ; zanella2020yacos , which is impractical due to prohibitive computations to construct labeled training data by search. Another more promising routine, reinforcement learning, that successfully discovered faster sorting algorithm (AlphaDev fawzi2022discovering ) and matrix multiplication algorithm (AlphaTensor mankowitz2023faster ), can explore the optimization space from feedback on optimization metrics without requiring optimal labeled data. For both techniques, broad generalization across different programs, even out of the training samples, arises as a major bottleneck. The community has noted that in a range of machine learning applications devlin2018bert ; brown2020language ; radford2021learning ; kirillov2023segment , training high-capacity models on large-scale datasets has yielded unprecedented performances. For example, large language models like GPT-4 have demonstrated impressive zero-shot generalization abilities ouyang2022training ; openai2023gpt4 . However, in the area of code optimization, the current practice is to learn optimization heuristics in a per-program manner huang2019autophase ; shahzad2022reinforcement or from a small training set with hundreds of programs mammadli2020static ; jain2022poset , lagging far behind the new era of solving challenging problems by scaling up machine learning models.
Figure 1: CodeZero agent performs code optimization using a learned policy in a compiler environment. Guided by a learned policy, at each step, CodeZero analyzes the Intermediate Representation (IR) of the program, selects to apply an optimization pass, and receives a reward based on the improvement in optimization metrics. Through sequential interactions, the agent aims to maximize cumulative rewards and enhance the final IR’s performance.
In this study, we focus on the LLVM lattner2004llvm phase ordering problem, a longstanding challenge for compiler research, and propose CodeZero, a reinforcement-learned code optimization agent, capable of generating a sequence of optimization passes tailored to a particular input program. The problem is formulated such that a code optimization agent, upon encountering a program, selects a series of optimization passes, and receives feedback based on the outcomes of applying these transformations, from the compiler environment. To tackle the huge range of possible test programs, we aim to enhance the generalization ability of trained agents by assembling a large-scale training dataset of natural programs. Several previous works rely on randomly generated programs due to data scarcity but the significant distribution shift from random-generated and human-written programs can even hurt the generalization of learned agents when tested on real-world scenarios cummins2022compilergym . Our training data not only includes real-world programs sourced from GitHub da2021anghabench but also incorporates complex algorithmic solutions of competitive programming li2022competition and diverse programs generated by large language models (LLM) tihanyi2023formai . To tackle the vast space of possible optimizations ($\sim 10^{73}$ sequences), we employ a state-of-the-art model-based reinforcement learning method hafner2023mastering to train the code optimization agent sample-efficiently. This method not only learns a predictive world model of compilers to reduce the amount of real compiler executions but also benefits generalization by learning representations that better capture the structure of the compiler state transitions anand2021procedural . After training on massive programs via trial-and-error, the code optimization agent can generalize in a zero-shot way to previously unseen programs with superior optimization performance against off-the-shelf compiler heuristics.
We demonstrate the effectiveness of our trained CodeZero agent on a range of domains cummins2022compilergym from benchmark suites encompassing fundamental algorithms that are ubiquitously employed in everyday applications, to production-level open-source programs, including object files from C++ TensorFlow abadi2016tensorflow and OpenCV culjak2012brief library. On six test datasets, our agent can produce optimization sequences in a single trial, yielding more efficient code size reduction compared to the -Oz flag. Detailed analysis underscores the effectiveness of both the newly introduced program datasets and the model-based reinforcement learning technique within the realm of code optimization. We posit that the agents developed through our approach could be integrated into the existing toolkits of optimization strategy in compilers alongside other manually designed heuristics, such as -Oz or -O3.
2 CodeZero Agent
2.1 Code Optimization as Decision Making
As illustrated in Figure 1, compilers consist of three main components: the front end which translates the source code into an intermediate representation (IR), the middle end, and the back end which converts IR to the binary code. The middle end is responsible for language- and platform-agnostic optimizations over the IR, implemented as passes to either collect information about the program or apply a transform on it, like function inlining, loop unrolling, and dead code elimination. For example, the LLVM-opt tool has more than 100 optimization passes available. A specific order of applying these passes forms an optimization sequence for an input program. This is critical as the right selection and ordering of passes can significantly boost the program’s performance. Despite that compiler developers have provided standard optimization sequences at various levels, e.g. -O2, -O3, -Oz, these preset sequences may not always yield optimal results, especially for emerging programs written in new frameworks such as TensorFlow. Particularly due to the increasing number of optimization passes, it is an open challenge to determine the most effective sequence for each program.
This problem, known as phase ordering, can be naturally formulated as a partially observable Markov decision process (POMDP) sutton2018reinforcement . In this formulation, a code optimization agent determines the optimization sequence of an input program through a series of interactions with the compiler environment, guided by its policy $\pi$. The process starts at the initial state $s_{0}$, representing the IR of the program to be optimized, which is randomly sampled from all IRs of interest. It is critical that aiming to capture the most important characteristics of the target optimization, the agent only receives partially observable information $o_{0}$ of the state. The observation space can vary widely, ranging from manually designed features (e.g. the number of basic blocks) huang2019autophase to more complex tree-based or graph-based program representations cummins2020programl ; alon2018general , and even raw text strings of IR cummins2017end . At each time step $t=0,1,2,\dots$, the agent takes an action $a_{t}$ based on its policy, corresponding to selecting an optimization pass. Following this action, the agent receives an immediate reward $r_{t+1}$, reflecting changes in the optimization metrics, and an observation $o_{t+1}$ of the next state $s_{t+1}=p(s_{t},a_{t})$ which represents the IR transformed by the compiler using the selected pass. This process can be terminated either when the agent finds no positive gains can be achieved or a maximum number of steps is reached. The goal of the agent is to learn an optimization policy $\pi(a_{t}\mid o_{\leq t})$ that effectively maximizes the cumulative rewards, thereby optimizing the performance metrics of the final IR.
2.2 Large-Scale Data Preparation
To ensure that our CodeZero agent can effectively generalize to unseen situations, a concept known as zero-shot generalization, we have identified three critical factors in preparing our training dataset. Firstly, the dataset must reflect naturalness. Training data should be within the distribution of human-written programs, otherwise, overfitting to programs that deviate significantly from this, such as those generated by tools like Csmith yang2011finding and llvm-stress lattner2004llvm , could provide no benefits or even hurt the generalization to real-world scenarios. Secondly, diversity in the dataset is essential. It should encompass a wide range of human-written program styles and structures, ensuring globally comprehensive coverage of possible scenarios. Lastly, the pursuit of high-quality training data is imperative. Existing large-scale program collections da2021anghabench ; cummins2020programl have been proposed to serve as training data cummins2022compilergym but are proven unfruitful in our preliminary experiments. Instead, we focus on data that features moderate lengths of IR, complex algorithmic logic, and potential room for optimizations. This allows the agent to flexibly explore and understand the local transformation structure within the IR space. Collectively, these three properties – naturalness, diversity, and high quality – contribute to aligning the visited IR distribution during training with that will be encountered in real-world applications.
We construct our training dataset containing hundreds of thousands of programs by combining three distinct single-source program datasets, CodeContests li2022competition , FormAI tihanyi2023formai , and AnghaBench da2021anghabench . CodeContests comprises human-written solutions to competitive programming problems with complex and optimizable algorithmic logic. FormAI is a large collection of AI-generated C programs with various functionality types and coding styles, aiming to enrich the dataset’s diversity. AnghaBench is a collection of real-world C programs extracted from GitHub. As shown by visualization in Figure 7, our training data has a broad coverage of evaluation programs in a variety of domains. Future expansions of the dataset can be conducted following the aforementioned principles.
2.3 Agent Training with Model-based RL
Figure 2: Design overview of CodeZero. (a) In the model-based training loop, the CodeZero agent interacts with the compiler environment across diverse training programs, learns a world model from historical experience, and updates its policy efficiently through model simulations. (b) Deployment of the trained agent is capable of zero-shot generalization to unseen programs, delivering effective optimizations. (c) The compiler environment is set up with the input program’s IR, $s_{0}$. At each step, upon receiving an optimization pass $a_{t}$, the environment executes the pass internally, resulting in a transformed IR $s_{t+1}$, and provides the agent with relevant IR features $o_{t+1}$ and immediate rewards $r_{t+1}$ based on optimization metric improvements. (d) The world model simulates the compiler environment, initiated with an observation, maintaining its internal states, and predicting future observations and rewards in response to input actions. (e) The policy shares with the world model a generalizable representation that captures the environment’s structure.
While existing work on reinforcement learning for code optimization prevalently focuses on model-free RL methods, model-based RL can offer advantages in terms of both sample efficiency sutton2018reinforcement and generalization anand2021procedural . Executing and profiling extensive optimization sequences, especially for the runtime metric, can be time-consuming. This is further compounded when constructing complex observations, such as control-data flow graphs brauckmann2020compiler ; cummins2020programl . Model-based RL addresses these challenges by learning a world model to approximate state transitions and reward signals of the environment. This allows the agent to learn its policy by simulating trajectories based on model predictions, rather than relying solely on trial-and-error interactions in the real compiler environment. This approach thus improves sample efficiency. Moreover, as the policy can share the representation with the world model, model learning can act as an auxiliary task and thus aid in learning representations that better capture the structure of the environment and manifest in better generalization of the policy mazoure2021cross .
We train the CodeZero agent by adapting an advanced model-based RL method, Dreamer hafner2023mastering , as depicted in Figure 2. This approach involves learning a predictive world model $(\hat{p}_{\theta},\hat{r}_{\theta})$ of the compiler environment by approximating the underlying transition dynamics $p(o_{t+1}|o_{\leq t},a_{\leq t})$ and reward function $r(o_{\leq t},a_{\leq t})$. Through imaginary rollouts using this world model over a horizon $H$, the policy $\pi_{\psi}(a_{t}|o_{\leq t})$ can be learned using the REINFORCE policy gradients williams1992simple with an entropy regularizer $\mathrm{H}$. The training objective is formulated as
$\mathcal{L}\left(\psi\right)\doteq\mathbb{E}_{\hat{p}_{\theta},\hat{r}_{\theta},\pi_{\psi}}\left[\sum_{t=1}^{H}-\left(V_{t}-v_{\xi}\left(o_{\leq t}\right)\right)\log\pi_{\psi}\left(a_{t}\mid o_{\leq t}\right)-\eta\mathrm{H}\left[\pi_{\psi}\left(o_{\leq t}\right)\right]\right],$ (1)
where $V_{t}$ is the imagined return estimation and the value prediction $v_{\xi}$ serves as a baseline for variance reduction greensmith2004variance . The value prediction is learned separately by regressing the cumulative rewards that the agent should expect to receive from the current state. All components are realized as deep neural networks.
After extensive training on a wide range of programs, our agent demonstrates the capability to generalize “zero-shot” — that is, to be effective without further training — to new, unseen programs across various domains.
3 Results
3.1 Evaluation Benchmarks
Our experiments focus on code size reduction, which benefits applications targeting low-resource hardware such as embedded systems. This focus is driven by the practical advantages of code size as a metric: it is both cost-effective and convenient to construct extensive compilable training and test datasets and to evaluate the optimization performance for code size.
We evaluate our method on benchmarks from the CompilerGym platform cummins2022compilergym : benchmark suites including cBench fursin2007midatasets , CHStone hara2008chstone , MiBench guthaus2001mibench , and NASA Parallel Benchmarks (NPB) bailey1995parallel , in addition to kernels from open source projects including BLAS lawson1979basic , Linux, OpenCV culjak2012brief , and TensorFlow abadi2016tensorflow . We disregard benchmarks generated by synthetic program generators lattner2004llvm ; yang2011finding since they do not align with real-world scenarios.
Figure 3: Code size reduction in terms of IR instruction count over LLVM -Oz under different methods. Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
Table 1: Top performances of the zero-shot CodeZero agent on individual programs. The agent’s optimization sequences are streamlined, omitting passes that do not contribute to performance enhancement. Full IRs of these programs before and after optimizations are shown in the Supplementary Information.
Dataset Program Pass Sequences by CodeZero Agent Code size
O0 CodeZero Oz (Reduction)
cBench sha -sroa -gvn -instcombine -simplifycfg -licm -early-cse -simplifycfg -indvars -gvn -simplifycfg -memcpyopt -reassociate 799 349 500 (1.43$\times$)
bzip2 -sroa -gvn -simplifycfg -instcombine -simplifycfg -early-cse -simplifycfg -reassociate -memcpyopt -jump-threading -functionattrs -licm -gvn -simplifycfg -reassociate -early-cse -jump-threading -early-cse -instcombine -simplifycfg 28748 13565 15946 (1.18$\times$)
qsort -sroa -gvn -simplifycfg -instcombine -simplifycfg -early-cse -gvn -indvars -gvn -simplifycfg 638 280 315 (1.12$\times$)
OpenCV #41 -lowerinvoke -simplifycfg -globalopt 28 18 28 (1.56$\times$)
#9 -sroa -gvn -sroa -early-cse -simplifycfg -loop-deletion -lowerinvoke -simplifycfg -sroa -early-cse -instcombine -simplifycfg -early-cse -memcpyopt -early-cse -simplifycfg -jump-threading -early-cse -reassociate -instcombine 9510 6341 9269 (1.18$\times$)
TensorFlow #17 -sroa -gvn -simplifycfg -lowerinvoke -simplifycfg -sroa -early-cse -indvars -simplifycfg 5512 4247 5450 (1.28$\times$)
#6 -gvn -simplifycfg -lowerinvoke -simplifycfg -instcombine -early-cse -gvn -jump-threading -instcombine 16173 12643 16076 (1.27$\times$)
3.2 Code Optimization Results
Figure 3 reports CodeZero’s performance in reducing code size, measured in terms of IR instruction count, compared to the expert-designed -Oz flag in LLVM. Remarkably, CodeZero, without any specific training on in-domain programs, outperforms -Oz across all but two test benchmark datasets in a single trial. The marginal performance variations between different methods on the BLAS and Linux datasets suggest that they are already highly optimized. CodeZero also demonstrates the ability to match or even slightly surpass a competitive autotuning method, random search, within a similar time budget (in the order of seconds), while the latter blindly aggregates hundreds of trials. This notable performance, coupled with the significant advantage over a single random trial, underscores the effectiveness of our agent’s policy learning. Moreover, CodeZero’s zero-shot generalization capability either matches or exceeds the performance of in-domain counterparts, which are trained using the train subset of each benchmark dataset. This is particularly significant in cases like the NPB dataset, which contains only 22 training programs. In such scenarios, where data sparsity presents a challenge for in-domain agents, the zero-shot CodeZero notably achieves an extra 3% reduction in code size. Notably, CodeZero’s pre-trained policy can generalize to IRs compiled from novel programming languages beyond C and C++, as evidenced in its performances on the BLAS and NPB datasets of Fortran programs. Further, in an AI-generated benchmark dataset of Objective-C, another language supported by LLVM-Clang apart from C/C++, CodeZero successfully improves upon LLVM -Oz, achieving an average code size reduction of 1.027$\times$, and reaching up to 2.87$\times$ in certain instances.
Figure 4: A comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model. The learned world model accurately captures the variations of program features and optimization metrics.
3.3 Program Case Study
We analyze the internal behavior of our CodeZero agent to optimize IR code via recording the sequence of passes chosen by the policy on individual programs. Table 1 presents the agent’s top performance outcomes on various benchmark datasets. We observe that the agent indeed produces a specialized optimization strategy tailored for each program. Additionally, these results also highlight certain passes, such as -sroa (scalar replacement of aggregates), -gvn (global value numbering), and -simplifycfg, as particularly effective in code size optimization.
In Figure 4, we display a predicted optimization trajectory for an unseen program from cBench, as forecasted by our learned compiler world model. The model successfully forecasts numeric features of future IR, including the counts of branches and blocks, alongside future rewards that signify optimization outcomes. This instance exemplifies the capability of our learned compiler world model to serve as a viable alternative for a real compiler environment in training code optimization agents.
3.4 Effect of Training Dataset
Figure 5: Evaluations of CodeZero with different datasets (a-b) and RL algorithms (c-d). (a) Performance of CodeZero trained on various datasets, with “Mixture” representing a combined dataset of CodeContests, FormAI, and AnghaBench. (b) Performance of CodeZero when trained on the full CodeContests dataset versus a subset of 100 randomly selected training programs from CodeContests. (c) Learning curves of various RL algorithms, measured by the geometric mean of code size reduction on the CodeContests validation set. A Gaussian filter ($\sigma=2.0$) is applied to enhance the visualization of trends. (d) Zero-shot generalization capabilities of different RL algorithms on various test benchmark datasets.
In our quest to understand the impact of training datasets on generalization, we explored how different datasets from various domains and sizes influence performance. As shown in Figure 5a, we discovered that the CodeContests dataset stands out as the most effective for generalizing to our test benchmarks. Although comparable outcomes were observed with other datasets like FormAI and AnghaBench, agents trained on these datasets notably lag in performance when tested on OpenCV and TensorFlow benchmarks. Moreover, combining these datasets does not yield performance enhancement. Consequently, we present in Figure 3 the results based on the agent trained exclusively with the CodeContests dataset. These results underscore the significance of high-quality training data, a principle that is increasingly recognized as vital in other scalable machine learning applications, such as large language models. In Figure 5b, we compare the agent trained on the full CodeContests dataset versus the one trained on a 100-program subset sampled from the same dataset. This subset is comparable in size to the datasets used in previous research mammadli2020static ; jain2022poset that applies reinforcement learning to code optimization. The suboptimal performance on this smaller dataset highlights its inadequacy for training a strong zero-shot generalizable agent and emphasizes the significance of a larger and more diverse training set prepared by this study.
3.5 Sample Efficiency and Zero-Shot Generalization
We further evaluate the sample efficiency and zero-shot generalization abilities of CodeZero based on the world model algorithm, Dreamer, against a range of model-free counterparts. These include notable algorithms such as DQN mnih2015human , A2C mnih2016asynchronous , APEX horgan2018distributed , IMPALA espeholt2018impala , and PPO schulman2017proximal (see Appendix B.4 for more details). Figure 5c shows that while PPO is the most competitive among the model-free baselines, Dreamer outperforms it by learning an order of magnitude faster, requiring fewer interactions with the compiler environment to achieve excellent performance. Additionally, as Figure 5d indicates, Dreamer excels in generalizing to unseen test benchmarks, outstripping its counterparts. These findings verify our hypothesis that a world model-based agent holds a stronger capability to capture the environment’s structure and can effectively enhance zero-shot generalization.
4 Discussion
We tackled the major challenge of zero-shot generalization that arises when applying artificial intelligence techniques to improve code optimization. We introduce the CodeZero agent that leverages the strengths of a diverse, high-quality training program dataset, as well as the sample efficiency and generalization capabilities of world model-based reinforcement learning. Our results have shown that CodeZero, with its zero-shot generalization ability, attains supercompiler code optimization performance, in the challenging phase ordering problem for code size reduction. Our study provides an AI-centric framework and serves as a meaningful step towards scaling machine learning techniques for code optimization. There is a substantial scope for further exploration, including expansion of the training dataset, scaling up the compiler world models, optimizing multiple objectives like execution time, and enriching feature and action spaces with deeper expert knowledge or large language models.
5 Methods
In this section, we provide the mathematical and computational details of our approach.
5.1 LLVM Phase Ordering POMDP
We formulate the phase ordering problem of LLVM as a partially observable Markov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\mu,\mathcal{O},\phi)$. The state space $\mathcal{S}$ covers all possible Intermediate Representations (IRs), the action space $\mathcal{A}$ comprises individual compiler optimization passes, and the reward function $r$ is defined by the metric being optimized. The transition dynamics $p:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ represents the outcome of applied IR transformations. The initial state distribution $\mu\in\Delta(\mathcal{S})$ captures all IRs of interest, which can be approximated via uniform sampling from the training dataset. The observation function $\phi:\mathcal{S}\mapsto\mathcal{O}$ maps the underlying IR into the observation space. Complex observation spaces offer comprehensive program information, while expert-designed features embed problem-specific knowledge, potentially enhancing the optimization policy’s generalization by eliminating irrelevant details. The code optimization agent with a policy $\pi$ interacts with the compiler environment according to the protocol described in Section 2.1. We fix the horizon of interactions to 45 steps.
Following Autophase huang2019autophase and CompilerGym cummins2022compilergym , the action space consists of 42 optimization passes out of over a hundred in LLVM, which effectively prunes the vast optimization space while not sacrificing the performance of the learned policy. The observation space is a concatenation of two numeric features: the 56-dimension Autophase feature summarizing the statistics of the IR, such as the number of specific basic blocks, branches, and instructions, and a 42-dimension histogram of the agent’s previous actions. The reward function is defined as the normalized change of the optimization metric $C(s)$:
$r_{t+1}=\frac{C(s_{t})-C(s_{t+1})}{C(s_{0})-C(s_{\mathrm{b}})},$ (2)
where a lower $C$ indicates better performance and $C(s_{\mathrm{b}})$ stands for the performance of a baseline policy, such as built-in -Oz or -O3. A total reward greater than 1 means that the optimization sequence performs better than the baseline policy.
5.2 Training Dataset
We construct our training datasets, by considering three datasets from distinct domains. The CodeContests dataset li2022competition consists of over 13,000 competitive problems, each on average having hundreds of solutions in three languages. We subsample up to ten C++ solutions for each training problem, resulting in 110,240 programs, as part of our training data, and sample one solution for each of 100 test problems from CodeContests as our validation data. We also include the full Form-AI dataset tihanyi2023formai with 112,000 AI-generated programs and the collection of the largest 9,998 single-function programs and 15,264 multiple-function programs from the AnghaBench da2021anghabench into our training data.
5.3 Model-based RL Method
We utilize a deep model-based RL method, DreamerV3 hafner2023mastering , to model the compiler environment upon which it learns an effective policy that maps observations to actions with parameterized deep neural networks.
World Model
The world model simulating the compiler environment is formulated as a latent dynamics model with the following four components:
$&\text{Representation model:}&&z_{t}\sim q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})\\ &\text{Transition model:}&&\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})\\ &\text{Image decoder:}&&\hat{o}_{t}\sim p_{\theta}(\hat{o}_{t}\mid z_{t})\\ &\text{Reward decoder:}&&\hat{r}_{t}\sim p_{\theta}(\hat{r}_{t}\mid z_{t})\end{aligned}$ (3)
The representation model estimates the latent state $z_{t}$ based on the previous state $z_{t-1}$, the previous action $a_{t-1}$ and the current observation $o_{t}$, while the transition model predicts it directly from $z_{t-1}$ and $a_{t-1}$. The overall models are jointly learned by minimizing the negative evidence lower bound (ELBO) jordan1999introduction ; kingma2013auto :
$\mathcal{L}_{\text{model}}(\theta)\doteq\;\mathbb{E}_{q_{\theta}\left(z_{1:T}\mid a_{1:T},o_{1:T}\right)}$ $\Big{[}\sum_{t=1}^{T}\Big{(}{-\ln p_{\theta}(o_{t}\mid z_{t})-\ln p_{\theta}(r_{t}\mid z_{t})}$ (4)
$\,\,\,\,{+\,\text{KL}\left[q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})\,\|\,p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})\right]}\Big{)}\Big{]}.$
In practice, we employ a reward smoothing technique lee2023dreamsmooth to mitigate the sparsity and the long tail distribution of rewards during an episode. This is achieved through the following equation:
$r_{t}^{\prime}\leftarrow\alpha r_{t-1}^{\prime}+(1-\alpha)r_{t},\;\;t=1,2,\dots$ (5)
with $\alpha\in[0,1)$. Consequently, we train a reward decoder $p_{\theta}(\hat{r}_{t}^{\prime}\mid z_{t})$ to predict the smoothed rewards.
Actor-Critic Learning
The actor and critic neural networks are parameterized on top of the latent representations:
[…]
The actor and critic are jointly trained on the same imagined trajectories $\{\hat{z}_{\tau},\hat{a}_{\tau},\hat{r}_{\tau}\}$ with horizon $H$, generated by the interactions between the transition model and reward model in Eq. (3) and the actor: starting at the latent state $\hat{z}_{t}=z_{t}$, at each step $\tau=t,t+1,t+2,\dots$, the policy takes an action $\hat{a}_{\tau}\sim\pi_{\psi}\left(\hat{a}_{\tau}\mid\hat{z}_{\tau}\right)$, and transits to the next latent state $\hat{z}_{\tau+1}\sim p_{\theta}(\hat{z}_{\tau+1}\mid z_{\tau},a_{\tau})$ with a reward $\hat{r}_{\tau+1}\sim p_{\theta}(\hat{r}_{\tau+1}\mid\hat{z}_{\tau+1})$. The critic is trained to predict the $\lambda$-return sutton2018reinforcement through a discrete regression loss bellemare2017distributional ; hafner2023mastering :
$\mathcal{L}_{\text{critic}}(\xi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum^{t+H}_{\tau=t}-\log v_{\xi}(V_{\tau}^{\lambda}\mid\hat{z}_{\tau})\right],$ (7)
$V_{\tau}^{\lambda}\doteq\hat{r}_{\tau}+\gamma\begin{cases}(1-\lambda)v_{\xi}(\hat{z}_{\tau+1})+\lambda V_{\tau+1}^{\lambda}&\text{if}\ \tau<t+H\\ v_{\xi}(\hat{z}_{\tau+1})&\text{if}\ \tau=t+H.\end{cases}$ (8)
The actor, meanwhile, is trained to maximize the imagined return through the REINFORCE policy gradient williams1992simple :
$\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum^{t+H}_{\tau=t}\left(-\left(V_{\tau}^{\lambda}-v_{\xi}\left(\hat{z}_{\tau}\right)\right)\log\pi_{\psi}\left(\hat{a}_{\tau}\mid\hat{z}_{\tau}\right)-\eta\,\text{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\right)\right],$ (9)
where $\text{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]$ is an entropy regularization which encourages exploration, and $\eta$ is a hyperparameter that adjusts the regularization strength.
Appendix A Related Work
In this section, we will elaborate a more extensive review of the development of machine learning methods for code optimization in compilers.
One of the key challenges for compilation is to determine which code transformations to apply, how to apply them (e.g., using suitable parameters), and in what order. This requires effectively searching and evaluating a massive number of possible options, known as iterative compilation bodin1998iterative or autotuning datta2008stencil . However, this search-based approach only finds a good optimization for one specific program and does not generalize into a compiler heuristic. This limitation underscores the importance of integrating machine learning techniques.
Pioneering work has delved into supervised machine learning, adopting two main approaches. The first one learns a predictive model that can directly predict the best option. It involves iteratively compiling training programs to identify the most effective compilation strategy for each, which then serves as the labels of training data. An early example calder1997evidence used a neural network for branch prediction, and one more well-known work is MilepostGCC fursin2008milepost , a practical attempt to integrate machine learning into a production compiler, GCC. It employs models trained on a large dataset of programs distributed over the Internet. The second approach aims to learn a cost or performance function capable of estimating the quality of various compiler options, which enables evaluation of a range of possible options without the need to compile and profile each one stephenson2003meta ; luk2009qilin .
Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the requirement for collecting optimal labeled data kulkarni2012mitigating . This technique has been applied to optimize individual compilation heuristics, such as inlining trofin2021mlgo , loop transformation haj2020neurovectorizer ; brauckmann2021reinforcement , and graph partitioning mirhoseini2017device . Several works relevant to us have explored the full optimization pipeline, i.e. the LLVM phase ordering problem, including AutoPhase huang2019autophase , CORL mammadli2020static , and POSET-RL jain2022poset . These methods predominantly utilize model-free RL algorithms while our work pioneers the use of an advanced model-based RL approach to reduce real compiler interactions. Model-based RL sutton1991dyna learns a simulation model of the compiler environment, which is related in spirit to the aforementioned approach of supervised performance models. However, it goes a step further by also learning a policy capable of directly determining the best optimization option, thereby eliminating the necessity for a guided search process.
Both machine learning techniques require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering. The most prevalent feature vectors are based on the frequencies of various types of instructions within the programs magni2014automatic ; huang2019autophase , designed by expert intuitions. Numerous studies have aimed to reduce the cost of feature design. Following the success of word2vec embeddings in natural language processing mikolov2013distributed , methods like code2vec alon2019code2vec , inst2vec ben2018neural , and IR2vec venkatakeerthy2020ir2vec represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR). The surge in deep learning has enabled feeding raw information such as AST alon2018general , control-data flow graphs (CFG) brauckmann2020compiler ; cummins2020programl , and code token sequences cummins2017end into powerful deep neural networks, capable of learning useful representations end-to-end.
Platforms that expose the compiler as a playground for AI experiments have significantly reduced the entry barriers to intelligent compiler research. OpenTuner ansel2014opentuner and YaCoS filho2018yet serve as autotuning frameworks with a range of compiler optimization techniques. Our experiments utilize CompilerGym cummins2022compilergym , which offers user-friendly interfaces for researchers to interact with compilers in a reinforcement learning manner. We are optimistic that the future release of our trained code optimization agents, in conjunction with these platforms, can have a democratizing effect on applying AI techniques to compiler optimizations.
Appendix B Implementation Details
B.1 Compiler Environment
Our experiments are conducted on the CompilerGym platform cummins2022compilergym , version 0.2.5, with LLVM-10.0.0 integration.
B.2 Features and Actions
Table 2: Descriptions of 56-dimension Autophase features (adapted from cummins2022compilergym ).
Index Name Description
0 BBNumArgsHi Number of BB where total args for phi nodes is gt 5
1 BBNumArgsLo Number of BB where total args for phi nodes is [1, 5]
2 onePred Number of basic blocks with 1 predecessor
3 onePredOneSuc Number of basic blocks with 1 predecessor and 1 successor
4 onePredTwoSuc Number of basic blocks with 1 predecessor and 2 successors
5 oneSuccessor Number of basic blocks with 1 successor
6 twoPred Number of basic blocks with 2 predecessors
7 twoPredOneSuc Number of basic blocks with 2 predecessors and 1 successor
8 twoEach Number of basic blocks with 2 predecessors and successors
9 twoSuccessor Number of basic blocks with 2 successors
10 morePreds Number of basic blocks with gt. 2 predecessors
11 BB03Phi Number of basic blocks with Phi node count in range (0, 3]
12 BBHiPhi Number of basic blocks with more than 3 Phi nodes
13 BBNoPhi Number of basic blocks with no Phi nodes
14 BeginPhi Number of Phi-nodes at beginning of BB
15 BranchCount Number of branches
16 returnInt Number of calls that return an int
17 CriticalCount Number of critical edges
18 NumEdges Number of edges
19 const32Bit Number of occurrences of 32-bit integer constants
20 const64Bit Number of occurrences of 64-bit integer constants
21 numConstZeroes Number of occurrences of constant 0
22 numConstOnes Number of occurrences of constant 1
23 UncondBranches Number of unconditional branches
24 binaryConstArg Binary operations with a constant operand
25 NumAShrInst Number of AShr instructions
26 NumAddInst Number of Add instructions
27 NumAllocaInst Number of Alloca instructions
28 NumAndInst Number of And instructions
29 BlockMid Number of basic blocks with instructions between [15, 500]
30 BlockLow Number of basic blocks with less than 15 instructions
31 NumBitCastInst Number of BitCast instructions
32 NumBrInst Number of Br instructions
33 NumCallInst Number of Call instructions
34 NumGetElementPtrInst Number of GetElementPtr instructions
35 NumICmpInst Number of ICmp instructions
36 NumLShrInst Number of LShr instructions
37 NumLoadInst Number of Load instructions
38 NumMulInst Number of Mul instructions
39 NumOrInst Number of Or instructions
40 NumPHIInst Number of PHI instructions
41 NumRetInst Number of Ret instructions
42 NumSExtInst Number of SExt instructions
43 NumSelectInst Number of Select instructions
44 NumShlInst Number of Shl instructions
45 NumStoreInst Number of Store instructions
46 NumSubInst Number of Sub instructions
47 NumTruncInst Number of Trunc instructions
48 NumXorInst Number of Xor instructions
49 NumZExtInst Number of ZExt instructions
50 TotalBlocks Number of basic blocks
51 TotalInsts Number of instructions (of all types)
52 TotalMemInst Number of memory instructions
53 TotalFuncs Number of non-external functions
54 ArgsPhi Total arguments to Phi nodes
55 testUnary Number of Unary operations
We extract program features following the approach of Autophase huang2019autophase . As described in Section 5.1, we use a 56-dimension Autophase feature vector concatenated with a 42-dimension action histogram vector as the observation. The Autophase feature contains various statistics of the IR code, with each dimension thoroughly explained in Table 2. The action histogram vector contains the counts of actions already taken by the agent within the current episode. Both vectors are normalized to ensure that their values fall within a reasonable range. Specifically, each element of the Autophase vector is divided by the total instruction count of the program, whereas the action histogram vector is normalized by the total action count per episode, set as 45.
Our action space is derived from Autophase huang2019autophase , comprising 45 LLVM optimization passes. However, CompilerGym excludes 3 actions due to updates in the latest LLVM version. Thus, we utilize a total of 42 actions. Table 3 presents the names of the LLVM optimization passes corresponding to these 42 actions.
Table 3: A list of LLVM transformation passes selected as actions.
Index Name Index Name Index Name
0 -adce 14 -instcombine 28 -lowerinvoke
1 -break-crit-edges 15 -ipsccp 29 -lowerswitch
2 -constmerge 16 -jump-threading 30 -mem2reg
3 -correlated-propagation 17 -lcssa 31 -memcpyopt
4 -deadargelim 18 -licm 32 -partial-inliner
5 -dse 19 -loop-deletion 33 -prune-eh
6 -early-cse 20 -loop-idiom 34 -reassociate
7 -functionattrs 21 -loop-reduce 35 -sccp
8 -functionattrs 22 -loop-rotate 36 -simplifycfg
9 -globaldce 23 -loop-simplify 37 -sink
10 -globalopt 24 -loop-unroll 38 -sroa
11 -gvn 25 -loop-unswitch 39 -strip
12 -indvars 26 -lower-expect 40 -strip-nondebug
13 -inline 27 -loweratomic 41 -tailcallelim
B.3 Hyperparameters
The hyperparameters for our DreamerV3 agent implementation are outlined in Table 4. For hyperparameters not specified, we use the same value as the original DreamerV3 hafner2023mastering .
Table 4: Hyperparameters of DreamerV3 in our experiments.
Hyperparameter Value
Architecture RSSM recurrent units 1024
RSSM number of latents 32
RSSM classes per latent 32
MLP layers 4
MLP hidden units 400
Activation LayerNorm + SiLU
Training Random exploration 500 environment steps
Replay buffer capacity $2\times 10^{6}$
Reward smoothing $\alpha$ lee2023dreamsmooth 0.6
Training frequency Every 5 environment steps
Batch size 50
Batch length $T$ 50
Imagination horizon $H$ 15
Discount $\gamma$ 0.99
$\lambda$-target discount 0.95
World model loss scales 100.0 for Autophase
10.0 for action histogram
1.0 for reward
5.0 for discount
0.1 for KL
Actor entropy regularization $\eta$ $3\times 10^{-4}$
KL balancing 0.8
Optimizer Adam
World model learning rate $1\times 10^{-4}$
Actor-critic learning rate $3\times 10^{-5}$
Weight decay $1\times 10^{-5}$
Gradient clipping 100
B.4 Model-Free Baselines
We use RLlib liang2018rllib to train and test model-free reinforcement learning algorithms including PPO schulman2017proximal , A2C mnih2016asynchronous , IMPALA espeholt2018impala , APEX horgan2018distributed , and DQN mnih2015human . All the experiments are conducted with 10 CPUs and an RTX-3090 GPU over a minimum duration of 10 hours. The training process uses 5 rollout workers to interact with the environment and use 4 evaluation workers to evaluate the checkpoint on the validation set. We use default hyperparameters of algorithms in RLlib following the CompilerGym platform cummins2022compilergym , except that we have carefully tuned the hyperparameters for our strongest baseline, PPO, as listed in Table 5.
Table 5: Hyperparameters for the PPO baseline, well tuned on our dataset to be deviating from the default value in RLlib.
Hyperparameters Value
PPO gamma 1.0
use_gae True
lambda_ 1.0
train_batch_size 9000
lr 5e-5
kl_coeff 0.2
kl_target 0.01
vf_loss_coeff 1.0
num_sgd_iter 30
sgd_minibatch_size 128
clip_param 0.3
vf_clip_param 10.0
weight_decay 1e-6
Appendix C Benchmarks
CompilerGym Benchmarks
In our study, we select eight benchmarks for zero-shot test and in-domain training: BLAS, cBench, CHStone, Linux, MiBench, NPB, OpenCV, and TensorFlow. These benchmarks are part of the built-in datasets provided by CompilerGym version 0.2.5. For benchmarks with a total number of programs more than 100, we use the first 50 programs as the test set, the following 50 programs as the validation set, and all of the rest as the training set. These training and validation sets are only used for in-domain training. The datasets comprising fewer than 100 programs are not applicable for in-domain training; instead, all their programs are allocated to the test set. The number of programs in each dataset after division is detailed in Table 6.
Table 6: Dataset division of 8 CompilerGym benchmarks.
Dataset Training Split Validation Split Test Split
BLAS 200 50 50
cBench 23 N/A N/A
CHStone 12 N/A N/A
Linux 13,794 50 50
MiBench 40 N/A N/A
NPB 22 50 50
OpenCV 342 50 50
TensorFlow 1,885 50 50
AI-Generated Benchmarks
To further test the generalization ability of our CodeZero agent on different programming languages, we borrow the method from FormAI tihanyi2023formai and generate a dataset containing 50 unique Objective-C programs using GPT-3.5. We use the same prompt as FormAI, except that we add an instruction to ask GPT to generate programs that can be directly compiled under Clang version 10.0.0 and do not use ARC (Automatic Reference Counting), to improve the compilation pass rate of generated programs. We compile the generated programs using Clang without including any third-party libraries, and all programs that cannot pass compilation are discarded.
Appendix D Extended Experimental Results
D.1 Learning Curves
The zero-shot test performance of our CodeZero agents during training is shown in Figure 6. Note that we validate and test the agent every 10000 environment steps and report the test performance from the checkpoint that achieved the best validation results for comparison among various methods.
Figure 6: Zero-shot test performance of the CodeZero agent during training. We report mean and standard deviation across three runs.
D.2 Quantitative Results
Quantitative results corresponding to Figure 3 in the main text are provided in Table 7.
Table 7: Quantitative results for code size reduction, corresponding to Figure 3. We report mean and standard deviation across three runs.
-O0 Random Policy
Random Search
CodeZero
(In-domain)
CodeZero
(Zero-shot)
BLAS geomean 0.931 0.821$\pm$0.018 0.960$\pm$0.004 0.993$\pm$0.005 0.991$\pm$0.000
min 0.707 0.432$\pm$0.106 0.723$\pm$0.076 0.831$\pm$0.148 0.913$\pm$0.000
max 1.000 1.002$\pm$0.004 1.031$\pm$0.010 1.025$\pm$0.005 1.016$\pm$0.000
cBench geomean 0.481 0.737$\pm$0.015 0.858$\pm$0.020 N/A 1.036$\pm$0.002
min 0.299 0.401$\pm$0.097 0.560$\pm$0.030 N/A 0.962$\pm$0.001
max 0.626 1.150$\pm$0.113 1.298$\pm$0.083 N/A 1.395$\pm$0.054
CHStone geomean 0.487 0.701$\pm$0.043 1.037$\pm$0.006 N/A 1.094$\pm$0.005
min 0.402 0.278$\pm$0.108 0.889$\pm$0.044 N/A 1.000$\pm$0.006
max 0.655 1.038$\pm$0.033 1.304$\pm$0.038 N/A 1.378$\pm$0.021
Linux geomean 0.988 0.981$\pm$0.004 1.001$\pm$0.000 0.993$\pm$0.000 0.986$\pm$0.005
min 0.615 0.740$\pm$0.126 1.000$\pm$0.000 0.700$\pm$0.001 0.642$\pm$0.041
max 1.011 1.001$\pm$0.002 1.011$\pm$0.000 1.011$\pm$0.000 1.009$\pm$0.003
Mibench geomean 0.389 0.812$\pm$0.022 1.005$\pm$0.001 N/A 1.006$\pm$0.002
min 0.278 0.352$\pm$0.074 0.858$\pm$0.010 N/A 0.879$\pm$0.003
max 0.760 1.392$\pm$0.074 1.603$\pm$0.015 N/A 1.588$\pm$0.000
NPB geomean 0.530 0.833$\pm$0.011 1.074$\pm$0.010 1.075$\pm$0.017 1.108$\pm$0.001
min 0.191 0.414$\pm$0.043 0.805$\pm$0.104 0.810$\pm$0.082 0.886$\pm$0.000
max 1.066 1.848$\pm$0.193 2.141$\pm$0.062 2.315$\pm$0.028 2.343$\pm$0.012
OpenCV geomean 0.981 0.949$\pm$0.011 1.080$\pm$0.001 1.087$\pm$0.007 1.092$\pm$0.000
min 0.833 0.603$\pm$0.133 0.888$\pm$0.021 0.898$\pm$0.001 0.897$\pm$0.000
max 1.370 1.409$\pm$0.167 1.571$\pm$0.022 1.635$\pm$0.186 1.556$\pm$0.000
TensorFlow geomean 0.983 0.912$\pm$0.009 1.006$\pm$0.003 1.032$\pm$0.000 1.032$\pm$0.001
min 0.927 0.618$\pm$0.046 0.877$\pm$0.009 0.970$\pm$0.010 0.968$\pm$0.011
max 1.010 1.235$\pm$0.034 1.267$\pm$0.007 1.289$\pm$0.008 1.282$\pm$0.002
D.3 Data Distribution Visualization
In Figure 7, we visualize the distribution of our training and test datasets. To accurately represent the dynamic behavior of programs, we randomly select 1000 action sequences, each with a length of 45, from our action space. These sequences are subsequently executed on each program, with the resulting Autophase features concatenated to form a feature vector with dimensions of $1000\times 45\times 56$ for every program. These comprehensive feature vectors are finally processed using t-SNE van2008visualizing for dimensionality reduction and visualization.
Figure 7 illustrates that our training data (denoted as triangles) has a broad coverage of test programs (denoted as circles). This contrasts with the Csmith dataset (denoted as crosses) employed in CompilerGym experiments cummins2022compilergym , which shows a significant deviation from real-world applications. Nonetheless, our visualization can still not perfectly capture the transferability across datasets. For instance, empirical evidence suggests that CodeContests are the most effective in generalizing to OpenCV and TensorFlow, while the visual analysis does not directly imply this.
Figure 7: t-SNE van2008visualizing visualization of programs from training and test datasets.

### v2

@@
-
+CompilerDream: Learning a Compiler World Model for General Code Optimization
 Abstract
-Effective code optimization in compilers plays a central role in computer and software engineering.
-While compilers can be made to automatically search the optimization space without the need for user interventions, this is not a standard practice since the search is slow and cumbersome.
-Here we present CodeZero, an artificial intelligence agent trained extensively on large data to produce effective optimization strategies instantly for each program in a single trial of the agent.
-To overcome the huge range of possible test programs, we prepare a large dataset of training programs that emphasize quality, naturalness, and diversity.
-To tackle the vast space of possible optimizations, we adapt deep reinforcement learning to train the agent in a sample-efficient manner through interacting with a world model of the compiler environment.
-Evaluation on both benchmark suites and production-level code optimization problems demonstrates our agent’s supercompiler performances and zero-shot generalization abilities, outperforming built-in optimization options designed by compiler experts.
-Our methodology kindles the great potential of artificial intelligence for engineering and paves the way for scaling machine learning techniques in the realm of code optimization.
+Effective code optimization in compilers is crucial for computer and software engineering.
+The success of these optimizations primarily depends on the selection and ordering of the optimization passes applied to the code.
+While most compilers rely on a fixed sequence of optimization passes, current methods to find the optimal sequence either employ impractically slow search algorithms or learning methods that struggle to generalize to code unseen during training.
+We introduce CompilerDream, a model-based reinforcement learning approach to general code optimization.
+CompilerDream comprises a compiler world model that accurately simulates the intrinsic properties of optimization passes and an agent trained on this model to produce effective optimization strategies.
+By training on a large-scale program dataset, CompilerDream is equipped to serve as a general code optimizer across various application scenarios and source-code languages.
+Our extensive experiments first highlight CompilerDream’s strong optimization capabilities for autotuning, where it leads the CompilerGym leaderboard.
+More importantly, the zero-shot generalization ability of large-scale trained compiler world model and agent, excels across diverse datasets, surpassing LLVM’s built-in optimizations and other state-of-the-art methods in both settings of value prediction and end-to-end code optimization.
 1 Introduction
-Entering the post-Moore’s law era, code optimization is crucial for computer and software engineering, which plays an important role in realizing the full potential of slow-growing hardware.
-Developers typically rely on a compiler’s ability to transform input programs into semantically equivalent but more efficient versions, improving metrics like execution time, code size, and power consumption.
-For example, standard optimization options -O1, -O2, and -O3 aim to reduce execution time, while the -Os and -Oz options are crafted to reduce code size.
-Still, it is not common for users to explore beyond these conventional compiler options.
-Given the vast diversity of programs and platforms, coupled with the increasing number of optimization passes integrated into compiler frameworks, these off-the-shelf optimization strategies predefined heuristically by compiler experts may struggle to guarantee near-optimal performance across ever-changing scenarios triantafyllis2003compiler ; georgiou2018less .
-Automatic code optimization is therefore crucial in compilers.
-Autotuning bodin1998iterative improves code by systematically searching the optimization space through iterative executing and profiling optimization strategies.
-This search technique can yield remarkable performance gains but must be rerun for each new program with thousands of compilations, which is too time-consuming to be practical for all but a few specialized use cases.
-Meanwhile, machine learning techniques hold the capabilities to generalize the optimization strategy of one program to other similar ones, thereby facilitating faster code optimization.
-A direct method is to use supervised learning to predict good optimization strategies of input programs calder1997evidence ; fursin2008milepost ; zanella2020yacos , which is impractical due to prohibitive computations to construct labeled training data by search.
-Another more promising routine, reinforcement learning, that successfully discovered faster sorting algorithm (AlphaDev fawzi2022discovering ) and matrix multiplication algorithm (AlphaTensor mankowitz2023faster ), can explore the optimization space from feedback on optimization metrics without requiring optimal labeled data.
-For both techniques, broad generalization across different programs, even out of the training samples, arises as a major bottleneck.
-The community has noted that in a range of machine learning applications devlin2018bert ; brown2020language ; radford2021learning ; kirillov2023segment , training high-capacity models on large-scale datasets has yielded unprecedented performances.
-For example, large language models like GPT-4 have demonstrated impressive zero-shot generalization abilities ouyang2022training ; openai2023gpt4 .
-However, in the area of code optimization, the current practice is to learn optimization heuristics in a per-program manner huang2019autophase ; shahzad2022reinforcement or from a small training set with hundreds of programs mammadli2020static ; jain2022poset , lagging far behind the new era of solving challenging problems by scaling up machine learning models.
-Figure 1: CodeZero agent performs code optimization using a learned policy in a compiler environment.
-Guided by a learned policy, at each step, CodeZero analyzes the Intermediate Representation (IR) of the program, selects to apply an optimization pass, and receives a reward based on the improvement in optimization metrics.
+Code optimization plays an important role in realizing the full potential of software and hardware.
+Developers desire a universal solution to transform input programs into semantically equivalent but more efficient versions without manual effort.
+To achieve this, compilers usually employ a front-end that translates the source code of any language into an intermediate representation (IR), a back-end that converts IR to the binary code optimizer, and a middle-end optimizer that performs language- and platform-agnostic optimizations over the IR of the original code (Figure 1).
+The optimizer is typically implemented as a series of passes applying transforms on the code.
+The performance of the optimization primarily depends on the selection and order of the optimization passes.
+Standard compilers use a few fixed sets of optimization sequences to enhance specific aspects of program performance.
+For example, -O1, -O2, and -O3 improve for execution speed, and -Os and -Oz for reducing program size.
+Obviously, given the vast diversity of programs and platforms, these off-the-shelf optimization strategies predefined by compiler experts are not optimal for most circumstances, and we can significantly improve the performance of a specific program if we can find a better optimization pass sequence compared to that provided by compilers (Triantafyllis et al.
+2003; Georgiou et al.
+2018).
+Figure 1: CompilerDream agent performs code optimization using a learned policy in a compiler environment.
+Guided by a learned policy, at each step, CompilerDream analyzes the Intermediate Representation (IR) of the program, selects to apply an optimization pass, and receives a reward based on the improvement in optimization metrics.
 Through sequential interactions, the agent aims to maximize cumulative rewards and enhance the final IR’s performance.
-In this study, we focus on the LLVM lattner2004llvm phase ordering problem, a longstanding challenge for compiler research, and propose CodeZero, a reinforcement-learned code optimization agent, capable of generating a sequence of optimization passes tailored to a particular input program.
-The problem is formulated such that a code optimization agent, upon encountering a program, selects a series of optimization passes, and receives feedback based on the outcomes of applying these transformations, from the compiler environment.
-To tackle the huge range of possible test programs, we aim to enhance the generalization ability of trained agents by assembling a large-scale training dataset of natural programs.
-Several previous works rely on randomly generated programs due to data scarcity but the significant distribution shift from random-generated and human-written programs can even hurt the generalization of learned agents when tested on real-world scenarios cummins2022compilergym .
-Our training data not only includes real-world programs sourced from GitHub da2021anghabench but also incorporates complex algorithmic solutions of competitive programming li2022competition and diverse programs generated by large language models (LLM) tihanyi2023formai .
-To tackle the vast space of possible optimizations ($\sim 10^{73}$ sequences), we employ a state-of-the-art model-based reinforcement learning method hafner2023mastering to train the code optimization agent sample-efficiently.
-This method not only learns a predictive world model of compilers to reduce the amount of real compiler executions but also benefits generalization by learning representations that better capture the structure of the compiler state transitions anand2021procedural .
-After training on massive programs via trial-and-error, the code optimization agent can generalize in a zero-shot way to previously unseen programs with superior optimization performance against off-the-shelf compiler heuristics.
-We demonstrate the effectiveness of our trained CodeZero agent on a range of domains cummins2022compilergym from benchmark suites encompassing fundamental algorithms that are ubiquitously employed in everyday applications, to production-level open-source programs, including object files from C++ TensorFlow abadi2016tensorflow and OpenCV culjak2012brief library.
-On six test datasets, our agent can produce optimization sequences in a single trial, yielding more efficient code size reduction compared to the -Oz flag.
-Detailed analysis underscores the effectiveness of both the newly introduced program datasets and the model-based reinforcement learning technique within the realm of code optimization.
-We posit that the agents developed through our approach could be integrated into the existing toolkits of optimization strategy in compilers alongside other manually designed heuristics, such as -Oz or -O3.
-2 CodeZero Agent
-2.1 Code Optimization as Decision Making
-As illustrated in Figure 1, compilers consist of three main components: the front end which translates the source code into an intermediate representation (IR), the middle end, and the back end which converts IR to the binary code.
-The middle end is responsible for language- and platform-agnostic optimizations over the IR, implemented as passes to either collect information about the program or apply a transform on it, like function inlining, loop unrolling, and dead code elimination.
-For example, the LLVM-opt tool has more than 100 optimization passes available.
-A specific order of applying these passes forms an optimization sequence for an input program.
-This is critical as the right selection and ordering of passes can significantly boost the program’s performance.
-Despite that compiler developers have provided standard optimization sequences at various levels, e.g.
--O2, -O3, -Oz, these preset sequences may not always yield optimal results, especially for emerging programs written in new frameworks such as TensorFlow.
-Particularly due to the increasing number of optimization passes, it is an open challenge to determine the most effective sequence for each program.
-This problem, known as phase ordering, can be naturally formulated as a partially observable Markov decision process (POMDP) sutton2018reinforcement .
-In this formulation, a code optimization agent determines the optimization sequence of an input program through a series of interactions with the compiler environment, guided by its policy $\pi$.
-The process starts at the initial state $s_{0}$, representing the IR of the program to be optimized, which is randomly sampled from all IRs of interest.
-It is critical that aiming to capture the most important characteristics of the target optimization, the agent only receives partially observable information $o_{0}$ of the state.
-The observation space can vary widely, ranging from manually designed features (e.g.
-the number of basic blocks) huang2019autophase to more complex tree-based or graph-based program representations cummins2020programl ; alon2018general , and even raw text strings of IR cummins2017end .
-At each time step $t=0,1,2,\dots$, the agent takes an action $a_{t}$ based on its policy, corresponding to selecting an optimization pass.
-Following this action, the agent receives an immediate reward $r_{t+1}$, reflecting changes in the optimization metrics, and an observation $o_{t+1}$ of the next state $s_{t+1}=p(s_{t},a_{t})$ which represents the IR transformed by the compiler using the selected pass.
-This process can be terminated either when the agent finds no positive gains can be achieved or a maximum number of steps is reached.
-The goal of the agent is to learn an optimization policy $\pi(a_{t}\mid o_{\leq t})$ that effectively maximizes the cumulative rewards, thereby optimizing the performance metrics of the final IR.
-2.2 Large-Scale Data Preparation
-To ensure that our CodeZero agent can effectively generalize to unseen situations, a concept known as zero-shot generalization, we have identified three critical factors in preparing our training dataset.
-Firstly, the dataset must reflect naturalness.
-Training data should be within the distribution of human-written programs, otherwise, overfitting to programs that deviate significantly from this, such as those generated by tools like Csmith yang2011finding and llvm-stress lattner2004llvm , could provide no benefits or even hurt the generalization to real-world scenarios.
-Secondly, diversity in the dataset is essential.
-It should encompass a wide range of human-written program styles and structures, ensuring globally comprehensive coverage of possible scenarios.
-Lastly, the pursuit of high-quality training data is imperative.
-Existing large-scale program collections da2021anghabench ; cummins2020programl have been proposed to serve as training data cummins2022compilergym but are proven unfruitful in our preliminary experiments.
-Instead, we focus on data that features moderate lengths of IR, complex algorithmic logic, and potential room for optimizations.
-This allows the agent to flexibly explore and understand the local transformation structure within the IR space.
-Collectively, these three properties – naturalness, diversity, and high quality – contribute to aligning the visited IR distribution during training with that will be encountered in real-world applications.
-We construct our training dataset containing hundreds of thousands of programs by combining three distinct single-source program datasets, CodeContests li2022competition , FormAI tihanyi2023formai , and AnghaBench da2021anghabench .
-CodeContests comprises human-written solutions to competitive programming problems with complex and optimizable algorithmic logic.
-FormAI is a large collection of AI-generated C programs with various functionality types and coding styles, aiming to enrich the dataset’s diversity.
-AnghaBench is a collection of real-world C programs extracted from GitHub.
-As shown by visualization in Figure 7, our training data has a broad coverage of evaluation programs in a variety of domains.
-Future expansions of the dataset can be conducted following the aforementioned principles.
-2.3 Agent Training with Model-based RL
-Figure 2: Design overview of CodeZero.
-(a) In the model-based training loop, the CodeZero agent interacts with the compiler environment across diverse training programs, learns a world model from historical experience, and updates its policy efficiently through model simulations.
-(b) Deployment of the trained agent is capable of zero-shot generalization to unseen programs, delivering effective optimizations.
+Automatically optimizing the pass sequence is thus crucial in improving the compiler optimization performance.
+To be practical, such an algorithm must produce a satisfactory pass sequence within a reasonable time and handle a wide variety of programs.
+However, current research often fails to meet these requirements simultaneously.
+Search-based methods (Bodin et al.
+1998) ensure near-optimality but require thousands of compilations for each new program, making it too time-consuming for regular use.
+In contrast, machine learning methods can either predict the optimization sequences directly or estimate the outcomes of applying an optimization sequence to guide a search, thereby avoiding the time-consuming process of interacting with the compiler to validate the optimization outcome.
+However, in addition to potentially sacrificing some optimality compared to searching, these learning-based methods face a significant bottleneck in broad generalization across diverse programs that may be out of the training samples.
+A range of machine learning applications (Devlin et al.
+2019; Brown et al.
+2020; Radford et al.
+2021; Kirillov et al.
+2023) have witnessed that training high-capacity models on large-scale datasets can yield unprecedented performance.
+Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al.
+2022).
+However, the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al.
+2020; Shahzad et al.
+2022) or from relatively small training sets comprising only a few hundred programs (Mammadli, Jannesari, and Wolf 2020; Jain et al.
+2022).
+Additionally, the predominant use of model-free reinforcement learning methods with limited capacity may also hinder generalization.
+In this study, we focus on the LLVM (Lattner and Adve 2004) phase ordering problem, a longstanding challenge for compiler research, and propose CompilerDream, a world model-based reinforcement learning approach capable of generating an optimization pass sequence tailored to any particular input program.
+Our model-based method (Hafner et al.
+2023) not only learns an accurate predictive world model of compilers to simulate real compiler executions, whose number can be greatly reduced, but also benefits generalization by learning representations that better capture the structure of the compiler state transitions (Anand et al.
+2022).
+Trained on a carefully curated large-scale dataset of natural programs, the world model accurately captures the dynamics of the compiler optimization process, enabling the optimization agent to generalize to unseen programs and predict superior optimization sequences.
+We demonstrate the effectiveness of CompilerDream on a range of domains (Cummins et al.
+2022), from benchmark suites encompassing fundamental algorithms to production-level open-source programs, including object files from C++ TensorFlow (Abadi et al.
+2016) and OpenCV (Culjak et al.
+2012) library.
+Even without considering generalization, CompilerDream’s strong optimization capabilities place it at the top of the CompilerGym leaderboard for autotuning.
+For general code optimization, the large-scale trained world models of CompilerDream can accurately predict the outcomes of pass sequences on unseen programs, a setting known as value prediction.
+Beyond selecting optimization sequences based on value prediction, the agent can also directly produce optimization sequences in a single trial, yielding more efficient code size reduction compared to both the -Oz flag and state-of-the-art methods.
+The main contributions of this work are three-fold:
+  • We introduce CompilerDream, the first model-based reinforcement learning approach for code optimization.
+  • We first leverage a large-scale dataset, CodeContests, to achieve strong generalization on unseen programs.
+  • Extensive experiments spanning autotuning, value prediction, and end-to-end code optimization, demonstrate the superior performance of CompilerDream.
+2 Related Work
+One of the key challenges for compilation is to determine which code transformations to apply, how to apply them (e.g., using suitable parameters), and in what order.
+This requires effectively searching and evaluating a massive number of possible options, known as iterative compilation (Bodin et al.
+1998) or autotuning (Datta et al.
+2008).
+However, this search-based approach only finds a good optimization for one specific program and does not generalize into a compiler strategy.
+This limitation underscores the importance of integrating machine learning techniques.
+Supervised Learning
+Pioneering work has delved into supervised machine learning, adopting two main approaches (Leather and Cummins 2020).
+The first approach requires an extensive search on each training programs to identify the most effective optimization sequence, which then serves as the data labels.
+An early example (Calder et al.
+1997) used a neural network for branch prediction, and one more well-known work is MilepostGCC (Fursin et al.
+2008), a practical attempt to integrate machine learning into a production compiler, GCC.
+It employs models trained on a large dataset of programs distributed over the Internet.
+The second approach aims to learn a cost or performance function capable of estimating the quality of various compiler options, which enables evaluation of a range of possible options without the need to compile and profile each one (Stephenson et al.
+2003; Luk, Hong, and Kim 2009).
+Reinforcement Learning
+Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the requirement for collecting optimal labeled data (Kulkarni and Cavazos 2012).
+This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al.
+2021), loop transformation (Haj-Ali et al.
+2020; Brauckmann, Goens, and Castrillón 2021), and graph partitioning (Mirhoseini et al.
+2017).
+Several works relevant to us have explored the full optimization pipeline, i.e.
+the LLVM phase ordering problem, including AutoPhase (Haj-Ali et al.
+2020), CORL (Mammadli, Jannesari, and Wolf 2020), and POSET-RL (Jain et al.
+2022).
+Model-based Reinforcement Learning
+While existing work on reinforcement learning for code optimization predominantly focuses on model-free RL methods, our work pioneers the use of an advanced model-based RL approach.
+Model-based RL (Sutton and Barto 2018) can offer advantages in terms of both sample efficiency and generalization.
+Executing and profiling extensive optimization sequences, especially for the runtime metric, can be time-consuming.
+This is further compounded when constructing complex observations, such as control-data flow graphs (Brauckmann et al.
+2020; Cummins et al.
+2021).
+Model-based RL addresses these challenges by learning a world model that approximates state transitions and reward signals of the compiler.
+This is related in spirit to the aforementioned approach of supervised performance models but goes further by also learning a policy through simulating trajectories based on model predictions.
+This approach thus improves sample efficiency not only by eliminating the necessity for a guided search process during deployment but also by reducing trial-and-error interactions with the real compiler environment during training.
+Moreover, as the policy can share the representation with the world model, model learning can act as an auxiliary task and thus aid in learning representations that better capture the structure of the environment and manifest in better generalization of the policy (Anand et al.
+2022; Mazoure et al.
+2022; Ma et al.
+2024).
+Figure 2: Design overview of CompilerDream.
+(a) In the model-based training loop, the CompilerDream agent interacts with the compiler environment across diverse training programs, learns a world model from historical experience, and updates its policy efficiently through model simulations.
+(b) The trained agent can be deployed and is capable of zero-shot generalization to unseen programs, delivering effective optimizations.
 (c) The compiler environment is set up with the input program’s IR, $s_{0}$.
 At each step, upon receiving an optimization pass $a_{t}$, the environment executes the pass internally, resulting in a transformed IR $s_{t+1}$, and provides the agent with relevant IR features $o_{t+1}$ and immediate rewards $r_{t+1}$ based on optimization metric improvements.
-(d) The world model simulates the compiler environment, initiated with an observation, maintaining its internal states, and predicting future observations and rewards in response to input actions.
+(d) The world model simulates the compiler environment, initiated with an observation, maintaining its internal neural states, and predicting future observations and rewards in response to input actions.
 (e) The policy shares with the world model a generalizable representation that captures the environment’s structure.
-While existing work on reinforcement learning for code optimization prevalently focuses on model-free RL methods, model-based RL can offer advantages in terms of both sample efficiency sutton2018reinforcement and generalization anand2021procedural .
-Executing and profiling extensive optimization sequences, especially for the runtime metric, can be time-consuming.
-This is further compounded when constructing complex observations, such as control-data flow graphs brauckmann2020compiler ; cummins2020programl .
-Model-based RL addresses these challenges by learning a world model to approximate state transitions and reward signals of the environment.
-This allows the agent to learn its policy by simulating trajectories based on model predictions, rather than relying solely on trial-and-error interactions in the real compiler environment.
-This approach thus improves sample efficiency.
-Moreover, as the policy can share the representation with the world model, model learning can act as an auxiliary task and thus aid in learning representations that better capture the structure of the environment and manifest in better generalization of the policy mazoure2021cross .
-We train the CodeZero agent by adapting an advanced model-based RL method, Dreamer hafner2023mastering , as depicted in Figure 2.
-This approach involves learning a predictive world model $(\hat{p}_{\theta},\hat{r}_{\theta})$ of the compiler environment by approximating the underlying transition dynamics $p(o_{t+1}|o_{\leq t},a_{\leq t})$ and reward function $r(o_{\leq t},a_{\leq t})$.
-Through imaginary rollouts using this world model over a horizon $H$, the policy $\pi_{\psi}(a_{t}|o_{\leq t})$ can be learned using the REINFORCE policy gradients williams1992simple with an entropy regularizer $\mathrm{H}$.
-The training objective is formulated as
-$\mathcal{L}\left(\psi\right)\doteq\mathbb{E}_{\hat{p}_{\theta},\hat{r}_{\theta},\pi_{\psi}}\left[\sum_{t=1}^{H}-\left(V_{t}-v_{\xi}\left(o_{\leq t}\right)\right)\log\pi_{\psi}\left(a_{t}\mid o_{\leq t}\right)-\eta\mathrm{H}\left[\pi_{\psi}\left(o_{\leq t}\right)\right]\right],$ (1)
-where $V_{t}$ is the imagined return estimation and the value prediction $v_{\xi}$ serves as a baseline for variance reduction greensmith2004variance .
-The value prediction is learned separately by regressing the cumulative rewards that the agent should expect to receive from the current state.
-All components are realized as deep neural networks.
-After extensive training on a wide range of programs, our agent demonstrates the capability to generalize “zero-shot” — that is, to be effective without further training — to new, unseen programs across various domains.
-3 Results
-3.1 Evaluation Benchmarks
-Our experiments focus on code size reduction, which benefits applications targeting low-resource hardware such as embedded systems.
-This focus is driven by the practical advantages of code size as a metric: it is both cost-effective and convenient to construct extensive compilable training and test datasets and to evaluate the optimization performance for code size.
-We evaluate our method on benchmarks from the CompilerGym platform cummins2022compilergym : benchmark suites including cBench fursin2007midatasets , CHStone hara2008chstone , MiBench guthaus2001mibench , and NASA Parallel Benchmarks (NPB) bailey1995parallel , in addition to kernels from open source projects including BLAS lawson1979basic , Linux, OpenCV culjak2012brief , and TensorFlow abadi2016tensorflow .
-We disregard benchmarks generated by synthetic program generators lattner2004llvm ; yang2011finding since they do not align with real-world scenarios.
-Figure 3: Code size reduction in terms of IR instruction count over LLVM -Oz under different methods. Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
-Table 1: Top performances of the zero-shot CodeZero agent on individual programs.
-The agent’s optimization sequences are streamlined, omitting passes that do not contribute to performance enhancement.
-Full IRs of these programs before and after optimizations are shown in the Supplementary Information.
-Dataset Program Pass Sequences by CodeZero Agent Code size
-O0 CodeZero Oz (Reduction)
-cBench sha -sroa -gvn -instcombine -simplifycfg -licm -early-cse -simplifycfg -indvars -gvn -simplifycfg -memcpyopt -reassociate 799 349 500 (1.43$\times$)
-bzip2 -sroa -gvn -simplifycfg -instcombine -simplifycfg -early-cse -simplifycfg -reassociate -memcpyopt -jump-threading -functionattrs -licm -gvn -simplifycfg -reassociate -early-cse -jump-threading -early-cse -instcombine -simplifycfg 28748 13565 15946 (1.18$\times$)
-qsort -sroa -gvn -simplifycfg -instcombine -simplifycfg -early-cse -gvn -indvars -gvn -simplifycfg 638 280 315 (1.12$\times$)
-OpenCV #41 -lowerinvoke -simplifycfg -globalopt 28 18 28 (1.56$\times$)
-#9 -sroa -gvn -sroa -early-cse -simplifycfg -loop-deletion -lowerinvoke -simplifycfg -sroa -early-cse -instcombine -simplifycfg -early-cse -memcpyopt -early-cse -simplifycfg -jump-threading -early-cse -reassociate -instcombine 9510 6341 9269 (1.18$\times$)
-TensorFlow #17 -sroa -gvn -simplifycfg -lowerinvoke -simplifycfg -sroa -early-cse -indvars -simplifycfg 5512 4247 5450 (1.28$\times$)
-#6 -gvn -simplifycfg -lowerinvoke -simplifycfg -instcombine -early-cse -gvn -jump-threading -instcombine 16173 12643 16076 (1.27$\times$)
-3.2 Code Optimization Results
-Figure 3 reports CodeZero’s performance in reducing code size, measured in terms of IR instruction count, compared to the expert-designed -Oz flag in LLVM.
-Remarkably, CodeZero, without any specific training on in-domain programs, outperforms -Oz across all but two test benchmark datasets in a single trial.
-The marginal performance variations between different methods on the BLAS and Linux datasets suggest that they are already highly optimized.
-CodeZero also demonstrates the ability to match or even slightly surpass a competitive autotuning method, random search, within a similar time budget (in the order of seconds), while the latter blindly aggregates hundreds of trials.
-This notable performance, coupled with the significant advantage over a single random trial, underscores the effectiveness of our agent’s policy learning.
-Moreover, CodeZero’s zero-shot generalization capability either matches or exceeds the performance of in-domain counterparts, which are trained using the train subset of each benchmark dataset.
-This is particularly significant in cases like the NPB dataset, which contains only 22 training programs.
-In such scenarios, where data sparsity presents a challenge for in-domain agents, the zero-shot CodeZero notably achieves an extra 3% reduction in code size.
-Notably, CodeZero’s pre-trained policy can generalize to IRs compiled from novel programming languages beyond C and C++, as evidenced in its performances on the BLAS and NPB datasets of Fortran programs.
-Further, in an AI-generated benchmark dataset of Objective-C, another language supported by LLVM-Clang apart from C/C++, CodeZero successfully improves upon LLVM -Oz, achieving an average code size reduction of 1.027$\times$, and reaching up to 2.87$\times$ in certain instances.
-Figure 4: A comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model.
-The learned world model accurately captures the variations of program features and optimization metrics.
-3.3 Program Case Study
-We analyze the internal behavior of our CodeZero agent to optimize IR code via recording the sequence of passes chosen by the policy on individual programs.
-Table 1 presents the agent’s top performance outcomes on various benchmark datasets.
-We observe that the agent indeed produces a specialized optimization strategy tailored for each program.
-Additionally, these results also highlight certain passes, such as -sroa (scalar replacement of aggregates), -gvn (global value numbering), and -simplifycfg, as particularly effective in code size optimization.
-In Figure 4, we display a predicted optimization trajectory for an unseen program from cBench, as forecasted by our learned compiler world model.
-The model successfully forecasts numeric features of future IR, including the counts of branches and blocks, alongside future rewards that signify optimization outcomes.
-This instance exemplifies the capability of our learned compiler world model to serve as a viable alternative for a real compiler environment in training code optimization agents.
-3.4 Effect of Training Dataset
-Figure 5: Evaluations of CodeZero with different datasets (a-b) and RL algorithms (c-d).
-(a) Performance of CodeZero trained on various datasets, with “Mixture” representing a combined dataset of CodeContests, FormAI, and AnghaBench.
-(b) Performance of CodeZero when trained on the full CodeContests dataset versus a subset of 100 randomly selected training programs from CodeContests.
-(c) Learning curves of various RL algorithms, measured by the geometric mean of code size reduction on the CodeContests validation set.
-A Gaussian filter ($\sigma=2.0$) is applied to enhance the visualization of trends.
-(d) Zero-shot generalization capabilities of different RL algorithms on various test benchmark datasets.
-In our quest to understand the impact of training datasets on generalization, we explored how different datasets from various domains and sizes influence performance.
-As shown in Figure 5a, we discovered that the CodeContests dataset stands out as the most effective for generalizing to our test benchmarks.
-Although comparable outcomes were observed with other datasets like FormAI and AnghaBench, agents trained on these datasets notably lag in performance when tested on OpenCV and TensorFlow benchmarks.
-Moreover, combining these datasets does not yield performance enhancement.
-Consequently, we present in Figure 3 the results based on the agent trained exclusively with the CodeContests dataset.
-These results underscore the significance of high-quality training data, a principle that is increasingly recognized as vital in other scalable machine learning applications, such as large language models.
-In Figure 5b, we compare the agent trained on the full CodeContests dataset versus the one trained on a 100-program subset sampled from the same dataset.
-This subset is comparable in size to the datasets used in previous research mammadli2020static ; jain2022poset that applies reinforcement learning to code optimization.
-The suboptimal performance on this smaller dataset highlights its inadequacy for training a strong zero-shot generalizable agent and emphasizes the significance of a larger and more diverse training set prepared by this study.
-3.5 Sample Efficiency and Zero-Shot Generalization
-We further evaluate the sample efficiency and zero-shot generalization abilities of CodeZero based on the world model algorithm, Dreamer, against a range of model-free counterparts.
-These include notable algorithms such as DQN mnih2015human , A2C mnih2016asynchronous , APEX horgan2018distributed , IMPALA espeholt2018impala , and PPO schulman2017proximal (see Appendix B.4 for more details).
-Figure 5c shows that while PPO is the most competitive among the model-free baselines, Dreamer outperforms it by learning an order of magnitude faster, requiring fewer interactions with the compiler environment to achieve excellent performance.
-Additionally, as Figure 5d indicates, Dreamer excels in generalizing to unseen test benchmarks, outstripping its counterparts.
-These findings verify our hypothesis that a world model-based agent holds a stronger capability to capture the environment’s structure and can effectively enhance zero-shot generalization.
-4 Discussion
-We tackled the major challenge of zero-shot generalization that arises when applying artificial intelligence techniques to improve code optimization.
-We introduce the CodeZero agent that leverages the strengths of a diverse, high-quality training program dataset, as well as the sample efficiency and generalization capabilities of world model-based reinforcement learning.
-Our results have shown that CodeZero, with its zero-shot generalization ability, attains supercompiler code optimization performance, in the challenging phase ordering problem for code size reduction.
-Our study provides an AI-centric framework and serves as a meaningful step towards scaling machine learning techniques for code optimization.
-There is a substantial scope for further exploration, including expansion of the training dataset, scaling up the compiler world models, optimizing multiple objectives like execution time, and enriching feature and action spaces with deeper expert knowledge or large language models.
-5 Methods
-In this section, we provide the mathematical and computational details of our approach.
-5.1 LLVM Phase Ordering POMDP
-We formulate the phase ordering problem of LLVM as a partially observable Markov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\mu,\mathcal{O},\phi)$.
+3 Method
+In this section, we first present the mathematical formulation of the code optimization problem (Section 3.1).
+Then we introduce the CompilerDream approach for both compiler world model learning (Section 3.2) and optimization agent learning (Section 3.3).
+Finally, we elaborate on considerations of curating large-scale training dataset (Section 3.4).
+3.1 Phase Ordering Decision Process
+As illustrated in Figure 1, one key problem of compiler optimization is to find the optimal sequence of optimization passes for a given program, which is also referred to as the phase ordering problem.
+It can be naturally formulated as a partially observable Markov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\mu,\mathcal{O},\phi)$ (Sutton and Barto 2018).
 The state space $\mathcal{S}$ covers all possible Intermediate Representations (IRs), the action space $\mathcal{A}$ comprises individual compiler optimization passes, and the reward function $r$ is defined by the metric being optimized.
@@
 The observation function $\phi:\mathcal{S}\mapsto\mathcal{O}$ maps the underlying IR into the observation space.
-Complex observation spaces offer comprehensive program information, while expert-designed features embed problem-specific knowledge, potentially enhancing the optimization policy’s generalization by eliminating irrelevant details.
-The code optimization agent with a policy $\pi$ interacts with the compiler environment according to the protocol described in Section 2.1.
-We fix the horizon of interactions to 45 steps.
-Following Autophase huang2019autophase and CompilerGym cummins2022compilergym , the action space consists of 42 optimization passes out of over a hundred in LLVM, which effectively prunes the vast optimization space while not sacrificing the performance of the learned policy.
-The observation space is a concatenation of two numeric features: the 56-dimension Autophase feature summarizing the statistics of the IR, such as the number of specific basic blocks, branches, and instructions, and a 42-dimension histogram of the agent’s previous actions.
-The reward function is defined as the normalized change of the optimization metric $C(s)$:
-$r_{t+1}=\frac{C(s_{t})-C(s_{t+1})}{C(s_{0})-C(s_{\mathrm{b}})},$ (2)
-where a lower $C$ indicates better performance and $C(s_{\mathrm{b}})$ stands for the performance of a baseline policy, such as built-in -Oz or -O3.
-A total reward greater than 1 means that the optimization sequence performs better than the baseline policy.
-5.2 Training Dataset
-We construct our training datasets, by considering three datasets from distinct domains.
-The CodeContests dataset li2022competition consists of over 13,000 competitive problems, each on average having hundreds of solutions in three languages.
-We subsample up to ten C++ solutions for each training problem, resulting in 110,240 programs, as part of our training data, and sample one solution for each of 100 test problems from CodeContests as our validation data.
-We also include the full Form-AI dataset tihanyi2023formai with 112,000 AI-generated programs and the collection of the largest 9,998 single-function programs and 15,264 multiple-function programs from the AnghaBench da2021anghabench into our training data.
-5.3 Model-based RL Method
-We utilize a deep model-based RL method, DreamerV3 hafner2023mastering , to model the compiler environment upon which it learns an effective policy that maps observations to actions with parameterized deep neural networks.
-World Model
-The world model simulating the compiler environment is formulated as a latent dynamics model with the following four components:
-$&\text{Representation model:}&&z_{t}\sim q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})\\ &\text{Transition model:}&&\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})\\ &\text{Image decoder:}&&\hat{o}_{t}\sim p_{\theta}(\hat{o}_{t}\mid z_{t})\\ &\text{Reward decoder:}&&\hat{r}_{t}\sim p_{\theta}(\hat{r}_{t}\mid z_{t})\end{aligned}$ (3)
-The representation model estimates the latent state $z_{t}$ based on the previous state $z_{t-1}$, the previous action $a_{t-1}$ and the current observation $o_{t}$, while the transition model predicts it directly from $z_{t-1}$ and $a_{t-1}$.
-The overall models are jointly learned by minimizing the negative evidence lower bound (ELBO) jordan1999introduction ; kingma2013auto :
-$\mathcal{L}_{\text{model}}(\theta)\doteq\;\mathbb{E}_{q_{\theta}\left(z_{1:T}\mid a_{1:T},o_{1:T}\right)}$ $\Big{[}\sum_{t=1}^{T}\Big{(}{-\ln p_{\theta}(o_{t}\mid z_{t})-\ln p_{\theta}(r_{t}\mid z_{t})}$ (4)
-$\,\,\,\,{+\,\text{KL}\left[q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})\,\|\,p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})\right]}\Big{)}\Big{]}.$
-In practice, we employ a reward smoothing technique lee2023dreamsmooth to mitigate the sparsity and the long tail distribution of rewards during an episode.
-This is achieved through the following equation:
+Complex observation spaces offer comprehensive program information, while expert-designed features (Haj-Ali et al.
+2020) embed problem-specific knowledge, potentially enhancing the optimization policy’s generalization by eliminating irrelevant details.
+A code optimization agent guided by policy $\pi$ determines the optimization sequence of an input program through a series of interactions with the compiler environment.
+At each time step $t=0,1,2,\dots$, the agent applies an action $a_{t}$ to current IR based on observation $o_{t}$ and its policy $\pi(a_{t}\mid o_{\leq t})$, which transforms the current state $s_{t}$ to the next state $s_{t+1}=p(s_{t},a_{t})$ and will receive a reward $r_{t}$.
+The goal of the agent is to learn an optimization policy $\pi$ that effectively maximizes the cumulative rewards $G=\sum_{t}r_{t}$, thereby optimizing the performance metrics of the final IR.
+3.2 Learning a Compiler World Model
+We develop our CompilerDream approach by adapting an advanced model-based RL method, Dreamer (Hafner et al.
+2023), as depicted in Figure 2.
+This approach involves learning a predictive world model $(\hat{p}_{\theta},\hat{r}_{\theta})$ of the compiler environment parameterized by $\theta$ that approximates the underlying transition dynamics $p(o_{t+1}|o_{\leq t},a_{\leq t})$ and reward function $r(o_{\leq t},a_{\leq t})$.
+The compiler world model simulating the compiler environment is formulated with the following four components:
+Representation model: $z_{t}\sim q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})$ (1)
+Transition model: $\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})$
+Observation decoder: $\hat{o}_{t}\sim p_{\theta}(\hat{o}_{t}\mid z_{t})$
+Reward decoder: $\hat{r}_{t}\sim p_{\theta}(\hat{r}_{t}\mid z_{t}).$
+The representation model estimates a neural compiler state $z_{t}$ from the current observation $o_{t}$ of real compiler state, the previous state $z_{t-1}$ and the previous optimization action $a_{t-1}$.
+A representation loss $\mathcal{L}_{\mathrm{repr}}$ is used to train the neural compiler states to accurately reconstruct the observation and reward by two decoders:
+$\mathcal{L}_{\mathrm{repr}}(\theta)\doteq-\ln p_{\theta}(o_{t}|z_{t})-\ln p_{\theta}(r_{t}|z_{t}).$ (2)
+The transition model captures the dynamics of the compiler world model, predicting future neural state ${\hat{z}}_{t}$ directly from $z_{t-1}$ and $a_{t-1}$.
+A prediction loss minimizes the difference between the estimated state $z_{t}$ and the predicted state ${\hat{z}}_{t}$, simultaneously enhancing the transition model’s accuracy in predicting future states and making the state produced by the representation model easier to predict:
+$\mathcal{L}_{\mathrm{pred}}(\theta)\doteq\mathrm{KL}\left[q_{\theta}(z_{t}|z_{t-1},a_{t-1},o_{t})\,\|\,p_{\theta}(\hat{z}_{t}|z_{t-1},a_{t-1})\right].$ (3)
+The overall models are jointly learned by minimizing the sum of the representation loss and the prediction loss.
+Compiler simulation
+We can simulate the behavior of the real compiler using our trained compiler world model.
+Specifically, a simulated compiler optimization trajectory $\{\hat{z}_{\tau},\hat{a}_{\tau},\hat{r}_{\tau},\hat{o}_{\tau}\}$ with horizon $H$ can be generated by the interactions between the world model and an optimization agent: starting at a neural state $\hat{z}_{t}\sim q_{\theta}(z_{t}|o_{t})$, at each step $\tau=t,t+1,t+2,\dots$, the agent takes an action $\hat{a}_{\tau}\sim\pi_{\psi}\left(\hat{a}_{\tau}|\hat{z}_{\tau}\right)$, and transits to the next latent state $\hat{z}_{\tau+1}\sim p_{\theta}(\hat{z}_{\tau+1}|z_{\tau},a_{\tau})$ with a reward $\hat{r}_{\tau+1}\sim p_{\theta}(\hat{r}_{\tau+1}|\hat{z}_{\tau+1})$.
+Predicted real compiler state $\hat{o}_{\tau+1}$ can be optionally reconstructed by the observation decoder.
+Reward smoothing
+During an optimization sequence, not all passes can effectively transform the code, and the improvements gradually saturate.
+We employ a reward smoothing technique (Lee, Abbeel, and Lee 2024) to mitigate the sparsity and the long tail distribution of rewards during an episode.
+This is achieved through the following:
 $r_{t}^{\prime}\leftarrow\alpha r_{t-1}^{\prime}+(1-\alpha)r_{t},\;\;t=1,2,\dots$ (4)
 with $\alpha\in[0,1)$. Consequently, we train a reward decoder $p_{\theta}(\hat{r}_{t}^{\prime}\mid z_{t})$ to predict the smoothed rewards.
-Actor-Critic Learning
-The actor and critic neural networks are parameterized on top of the latent representations:
+3.3 Learning a Code Optimization Agent
+We then train a code optimization agent on top of the world model, which consists of an actor and a critic neural networks parameterized on top of the neural compiler state:
 […]
-The actor and critic are jointly trained on the same imagined trajectories $\{\hat{z}_{\tau},\hat{a}_{\tau},\hat{r}_{\tau}\}$ with horizon $H$, generated by the interactions between the transition model and reward model in Eq.
-(3) and the actor: starting at the latent state $\hat{z}_{t}=z_{t}$, at each step $\tau=t,t+1,t+2,\dots$, the policy takes an action $\hat{a}_{\tau}\sim\pi_{\psi}\left(\hat{a}_{\tau}\mid\hat{z}_{\tau}\right)$, and transits to the next latent state $\hat{z}_{\tau+1}\sim p_{\theta}(\hat{z}_{\tau+1}\mid z_{\tau},a_{\tau})$ with a reward $\hat{r}_{\tau+1}\sim p_{\theta}(\hat{r}_{\tau+1}\mid\hat{z}_{\tau+1})$.
-The critic is trained to predict the $\lambda$-return sutton2018reinforcement through a discrete regression loss bellemare2017distributional ; hafner2023mastering :
+The critic evaluates the cumulative optimization return that the actor can achieve starting from a given neural state, also known as value or return, denoted as $v_{\xi}(\hat{z}_{t})$.
+This is trained by minimizing the difference between the predicted value $v_{\xi}(\hat{z}_{t})$ and the Monte-Carlo return or more advanced bootstrapped $\lambda$-return (Sutton and Barto 2018):
 $\mathcal{L}_{\text{critic}}(\xi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum^{t+H}_{\tau=t}-\log v_{\xi}(V_{\tau}^{\lambda}\mid\hat{z}_{\tau})\right],$ (6)
-$V_{\tau}^{\lambda}\doteq\hat{r}_{\tau}+\gamma\begin{cases}(1-\lambda)v_{\xi}(\hat{z}_{\tau+1})+\lambda V_{\tau+1}^{\lambda}&\text{if}\ \tau<t+H\\ v_{\xi}(\hat{z}_{\tau+1})&\text{if}\ \tau=t+H.\end{cases}$ (8)
-The actor, meanwhile, is trained to maximize the imagined return through the REINFORCE policy gradient williams1992simple :
-$\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum^{t+H}_{\tau=t}\left(-\left(V_{\tau}^{\lambda}-v_{\xi}\left(\hat{z}_{\tau}\right)\right)\log\pi_{\psi}\left(\hat{a}_{\tau}\mid\hat{z}_{\tau}\right)-\eta\,\text{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\right)\right],$ (9)
-where $\text{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]$ is an entropy regularization which encourages exploration, and $\eta$ is a hyperparameter that adjusts the regularization strength.
-Appendix A Related Work
-In this section, we will elaborate a more extensive review of the development of machine learning methods for code optimization in compilers.
-One of the key challenges for compilation is to determine which code transformations to apply, how to apply them (e.g., using suitable parameters), and in what order.
-This requires effectively searching and evaluating a massive number of possible options, known as iterative compilation bodin1998iterative or autotuning datta2008stencil .
-However, this search-based approach only finds a good optimization for one specific program and does not generalize into a compiler heuristic.
-This limitation underscores the importance of integrating machine learning techniques.
-Pioneering work has delved into supervised machine learning, adopting two main approaches.
-The first one learns a predictive model that can directly predict the best option.
-It involves iteratively compiling training programs to identify the most effective compilation strategy for each, which then serves as the labels of training data.
-An early example calder1997evidence used a neural network for branch prediction, and one more well-known work is MilepostGCC fursin2008milepost , a practical attempt to integrate machine learning into a production compiler, GCC.
-It employs models trained on a large dataset of programs distributed over the Internet.
-The second approach aims to learn a cost or performance function capable of estimating the quality of various compiler options, which enables evaluation of a range of possible options without the need to compile and profile each one stephenson2003meta ; luk2009qilin .
-Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the requirement for collecting optimal labeled data kulkarni2012mitigating .
-This technique has been applied to optimize individual compilation heuristics, such as inlining trofin2021mlgo , loop transformation haj2020neurovectorizer ; brauckmann2021reinforcement , and graph partitioning mirhoseini2017device .
-Several works relevant to us have explored the full optimization pipeline, i.e.
-the LLVM phase ordering problem, including AutoPhase huang2019autophase , CORL mammadli2020static , and POSET-RL jain2022poset .
-These methods predominantly utilize model-free RL algorithms while our work pioneers the use of an advanced model-based RL approach to reduce real compiler interactions.
-Model-based RL sutton1991dyna learns a simulation model of the compiler environment, which is related in spirit to the aforementioned approach of supervised performance models.
-However, it goes a step further by also learning a policy capable of directly determining the best optimization option, thereby eliminating the necessity for a guided search process.
-Both machine learning techniques require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering.
-The most prevalent feature vectors are based on the frequencies of various types of instructions within the programs magni2014automatic ; huang2019autophase , designed by expert intuitions.
+$V_{\tau}^{\lambda}\doteq\hat{r}_{\tau}+\gamma\begin{cases}(1-\lambda)v_{\xi}(\hat{z}_{\tau+1})+\lambda V_{\tau+1}^{\lambda}&\text{if}\ \tau<t+H\\ v_{\xi}(\hat{z}_{\tau+1})&\text{if}\ \tau=t+H.\end{cases}$
+The actor produces a policy $\pi_{\psi}$ that predicts the best optimization pass to choose, represented as an action distribution, which is trained to maximize the simulated return through the REINFORCE policy gradient (Williams 1992):
+$(\displaystyle\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big{[}\sum^{t+H}_{\tau=t}\Big{(}$$-\left(V_{\tau}^{\lambda}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}|\hat{z}_{\tau})$
+$-\eta\,\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\Big{)}\Big{]},$ (7)
+where $\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]$ is an entropy regularization which encourages exploration, and $\eta$ is a hyperparameter that adjusts the regularization strength.
+Both the actor and critic are completely trained on world model-simulated compiler optimization trajectories.
+3.4 Data Curation
+To facilitate that our CompilerDream method can effectively generalize to unseen situations, a concept known as zero-shot generalization, we have identified three critical factors in preparing our training dataset.
+First, the dataset should reflect naturalness.
+During our preliminary experiments, we found that those generated code datasets like Csmith (Yang et al.
+2011) and llvm-stress (Lattner and Adve 2004) provide no benefits or even hurt the generalization to real-world scenarios.
+Second, the dataset should be large to prevent the agent from overfitting to a small number of codes and failing to generalize.
+Last, the code must exhibit high quality from an optimization perspective.
+We need data with complex algorithmic logic and potential for optimizations to help the world model and the agent better understand the problem’s intricacies.
+Datasets like AnghaBench (Da Silva et al.
+2021) which consist of millions of human-written codes collected from GitHub, are often too simple to allow for significant optimization improvements.
+Therefore, we choose to construct our training datasets on top of the CodeContests dataset released with AlphaCode (Li et al.
+2022), which consists of over 13,000 problems of coding competition, and each problem, on average, has hundreds of solutions in multiple languages.
+We subsample up to ten C++ solutions for each training problem, resulting in 110,240 programs, as our training data, and sample one solution for each of 100 test problems as our validation data.
+4 Experiments
+We evaluate CompilerDream in three different settings and compare its performance with prior state-of-the-art methods.
+We first demonstrate that, without considering generalization, CompilerDream excels as a powerful autotuning method, leading the CompilerGym leaderboard (Section 4.2).
+Then we show that CompilerDream can learn a general compiler world model capable of accurately estimating and selecting pass sequences (Section 4.3) and finally learn a general optimization agent that directly generates a pass sequence tailored to any specific input program.
+(Section 4.4)
+4.1 Evaluation
+Our experiments focus on code size reduction, which benefits applications targeting low-resource hardware such as embedded systems.
+This focus is driven by the practical advantages of code size as a metric: it is cost-effective to construct extensive compilable training and test datasets and to evaluate the optimization performance for code size.
+Metric
+To be more robust to outliners, we evaluate the code size optimization results by the geometric mean of IR instruction count reduction of program $s$ in a dataset $\mathcal{D}$:
+$R_{\mathcal{D}}(\pi)=\left(\prod_{s\in\mathcal{D}}\frac{C(s_{Oz})}{C(s_{t}^{\pi})}\right)^{\frac{1}{|\mathcal{D}|}}$ (8)
+where $C$ counts the number of IR-instructions, $s_{Oz}$ is the IR optimized by LLVM -Oz flag, and $s_{t}^{\pi}$ is the final IR produced by the algorithm $\pi$.
+A value above $1$ indicates that the algorithm performs better than -Oz option on average.
+Benchmarks
+We evaluate our method mainly on benchmarks from the CompilerGym platform (Cummins et al.
+2022): benchmark suites including cBench (Fursin et al.
+2007), CHStone (Hara et al.
+2008), MiBench (Guthaus et al.
+2001), and NASA Parallel Benchmarks (NPB) (Bailey et al.
+1995), in addition to kernels from open source projects including BLAS (Lawson et al.
+1979), Linux, OpenCV (Culjak et al.
+2012), and TensorFlow (Abadi et al.
+2016).
+We disregard benchmarks generated by synthetic program generators (Lattner and Adve 2004; Yang et al.
+2011) since they do not align with real-world scenarios.
+We follow the standard data splits of CompilerGym.
+Additionally, we evaluate CompilerDream on a large-scale dataset named FormAI (Tihanyi et al.
+2023).
+FormAI comprises a vast collection of AI-generated C programs with diverse functionalities and coding styles.
+We filtered out codes that failed to compile into CompilerGym benchmarks, resulting in a test set of 109,016 programs.
+Inspired by FormAI, we also constructed a dataset of 50 Objective-C programs generated by a Large Language Model to further evaluate CompilerDream’s ability to generalize to different programming languages.
+All details on benchmarks can be found in Appendix C.
+Figure 3: Generel code optimization results: Code size reduction in terms of IR instruction count over LLVM -Oz under different methods.
+Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
+4.2 Autotuning: CompilerGym Leaderboard
+Method Walltime
+Code size
+Reduction
+CompilerDream
+   + Guided Search
+60.8s 1.073$\times$
+PPO + Guided Search 69.8s 1.070$\times$
+CompilerDream 2.9s 1.068$\times$
+Random Search ($t$=10800) 10,512.4s 1.062$\times$
+Random Search ($t$=3600) 3,630.8s 1.061$\times$
+Greedy Search 169.2s 1.055$\times$
+GATv2 + DD-PPO 258.1s 1.047$\times$
+Table 1: Autotuning results: CompilerDream’s code size reduction and time consumption on cBench dataset compared with other methods from CompilerGym leaderboard (Facebook 2022).
+Results for methods other than ours are taken directly from the leaderboard, where learning-based methods report only inference time as wall time.
+We first use CompilerDream as an autotuning method to verify that its training process has sufficient capability to discover high-quality optimization sequences, supporting our goal of training an effective general code optimization agent.
+Implementation
+We target the task of CompilerGym leaderboard (Facebook 2022), optimizing pass sequences across 23 cBench programs.
+CompilerDream is trained for 25 hours, averaging about one hour per program—a time budget comparable to search methods.
+After training, besides evaluating single-trial optimization performance (more in Section 4.4), we test CompilerDream+Guided Search, following the leaderboard’s top method.
+This approach uses the agent’s policy $\pi_{\psi}(a_{t}|s_{t})$ to guide random search and limits search time to 1 minute per program (details in Appendix B.7).
+Results
+As shown in Table 1, CompilerDream’s policy achieves an average 1.068×1.068$\times$1.068 × code size reduction on cBench with a single trial after training, surpassing random search methods that require significantly more time—even when accounting for CompilerDream’s training time.
+This highlights the effectiveness of CompilerDream’s model-based RL training process in finding superior optimization passes, making it a promising component in building a general code optimizer.
+Moreover, when combined with guided search, CompilerDream outperforms the top method on the leaderboard while using less wall time.
+4.3 General Value Prediction
+We then demonstrate that the large-scale trained world model in CompilerDream can generalize to accurately simulate the optimization process of unseen input programs, thereby supporting the training of a general optimization agent.
+To illustrate this, we utilize our trained world model to estimate the outcome of applying an optimization sequence, a task named value prediction.
+Implementation
+We adopt the state-of-the-art Coreset-NVP method (Liang et al.
+2023), which evaluates pass sequences within a fixed candidate set (the coreset with 625 sequence prefixes) and selects the best prefix via compiler validation.
+Coreset-NVP employs a neural network to directly predict the normalized value of sequences, which is learned via supervised learning.
+Our Coreset-CompilerDream method leverages a world model trained on CodeContests, to simulate the optimization process and accumulate predicted rewards as the value.
+Both our method and the baseline method Coreset-NVP are tested on four datasets orthogonal to their training sets.
+The number of compiler validation calls is set the same across methods to ensure a fair comparison.
+Additional implementation details are available in Appendix B.8.
+Method Dataset
+cBench CHStone MiBench NPB
+Coreset-NVP 1.028 1.101 1.003 1.085
+Coreset
+-CompilerDream
+1.038 1.101 1.017 1.140
+Coreset-Oracle 1.041 1.106 1.020 1.159
+Table 2: General value prediction results: The geometric mean of code size reduction achieved by the best sequences selected by different methods across 4 datasets.
+Results of Coreset-NVP and Coreset-Oracle is reported from Coreset-NVP’s original paper (Liang et al.
+2023).
+Results
+Table 2 shows the geometric mean reduction achieved by different methods.
+Despite the test datasets being distinct from the training set, CompilerDream outperforms Coreset-NVP across three datasets, highlighting its strong compiler modeling and generalization capabilities.
+On MiBench, CompilerDream performs nearly optimally compared to the Oracle baseline, which uses a brute-force search and serves as the problem’s upper bound.
+Both our method and Coreset-NVP achieve an average reduction of $1.101$ on CHStone, close to the upper bound of $1.06$.
+4.4 General Code Optimization
+We finally aim to demonstrate that CompilerDream can learn a general optimization agent, capable of end-to-end generating effective optimization sequences for a wide variety of programs unseen during training.
+This scenario represents a typical daily use case where the optimization algorithm has limited time to compute the best solution for each of the diverse programs.
+Consequently, all methods generate only one optimization sequence per program in the test datasets unless otherwise specified.
+Figure 4: Histograms of large-scale evaluations comparing CompilerDream and PPO on the FormAI dataset.
+Figure 5: A comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model.
+The learned world model accurately captures the variations of program features and optimization metrics.
+Figure 6: Analysis.
+Evaluations of different RL algorithms: (a) Learning curves of various RL algorithms, measured by the geometric mean of code size reduction on the CodeContests validation set.
+A Gaussian filter ($\sigma=2.0$) is applied to enhance the visualization of trends.
+(b) Zero-shot generalization capabilities of different RL algorithms on various test benchmark datasets.
+Effect of training dataset: (c) Test performance of CompilerDream trained on CodeContests and Csmith.
+Bars indicate the standard deviation.
+Implementation
+We compare CompilerDream with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method.
+The random search conducts hundreds of trials within a time budget similar to CompilerDream’s.
+The -Oz flag represents LLVM’s highest level of code size optimization while the -O0 flag represents no optimization.
+The learning-based method using Proximal Policy Optimization (PPO) (Schulman et al.
+2017) and CompilerDream are trained on the large-scale CodeContests dataset and zero-shot generalized to unseen test programs.
+Following the state-of-the-art approach, all methods operate within a limited action space.
+Results
+Figure 3 shows CompilerDream’s performance in reducing code size, measured by the geometric mean of IR instruction count reduction.
+Without specific in-domain training, CompilerDream outperforms -Oz in all but two test benchmarks in a single trial and surpasses PPO in all but the BLAS datasets.
+It also exceeds random search within a similar time budget on most datasets, except Linux.
+The minimal performance differences on the BLAS and Linux datasets suggest they are already highly optimized.
+Moreover, CompilerDream’s zero-shot generalization matches or exceeds in-domain training.
+This advantage is especially notable in the NPB dataset, where data sparsity challenges in-domain agents, yet CompilerDream achieves an additional $3\%$ code size reduction.
+CompilerDream also generalizes well to new languages, evidenced in its performances on the BLAS and NPB datasets of Fortran programs.
+Further, on the AI-generated Objective-C dataset, CompilerDream achieves an average code size reduction of 1.027$\times$, reaching up to 2.87$\times$ in certain instances.
+On the large-scale FormAI test set (Figure 4), CompilerDream outperforms PPO, matching or outperforming -Oz on more programs and achieving higher optimization levels, demonstrating its superior and consistent performance.
+4.5 Analysis
+Comparison with model-free methods
+We further evaluate the sample efficiency and zero-shot generalization abilities of CompilerDream based on the world model algorithm, Dreamer, against a range of model-free counterparts, including PPO (Schulman et al.
+2017), DQN (Mnih et al.
+2015), A2C (Mnih et al.
+2016), APEX (Horgan et al.
+2018), and IMPALA (Espeholt et al.
+2018; Schulman et al.
+2017).
+Figure 6a illustrates that while PPO is the strongest model-free baseline, CompilerDream outperforms it, learning an order of magnitude faster with fewer compiler interactions.
+As Figure 6b demonstrates, CompilerDream also excels in generalizing to unseen benchmarks, supporting our hypothesis that world model-based agents better capture the environment’s structure and enhance zero-shot generalization.
+Effect of training dataset
+To assess the impact of the CodeContests dataset on the generalization ability, we compared it with the commonly used Csmith (Yang et al.
+2011) dataset, a large LLVM IR dataset generated by rules.
+We train CompilerDream on both datasets and the results shown in Figure 6 indicate that CompilerDream trained on CodeContests significantly outperforms the Csmith-trained version across all five test datasets.
+This advantage is particularly evident on the manually curated cBench (Fursin et al.
+2007) and CHStone (Hara et al.
+2008) datasets, demonstrating that the CodeContests-trained model can generalize more effectively to human-written programs.
+Program showcase
+In Figure 5, we display a predicted optimization trajectory for an unseen program from cBench, as forecasted by our learned compiler world model.
+The model successfully forecasts numeric features of future IR, including the counts of branches and blocks, alongside future rewards that signify optimization outcomes.
+This instance exemplifies the capability of our learned compiler world model to serve as a viable alternative for a real compiler environment in training code optimization agents.
+5 Discussion
+We aim to tackle the major challenge of zero-shot generalization that arises when applying learning-based code optimization.
+We thus introduce the CompilerDream approach that leverages the sample efficiency and generalization capabilities of world model-based reinforcement learning, as well as the strengths of a diverse, high-quality training program dataset.
+Our results have shown that CompilerDream, with its powerful generalization ability, attains optimization performance that outperforms existing methods, in multiple problem scenarios under the challenging area of phase ordering.
+There is substantial scope for further exploration, including expansion of the training dataset, scaling up the compiler world model, optimizing multiple objectives like execution time, and enriching feature and action spaces with deeper expert knowledge or large language models.
+Appendix A Extended Related Work
+Here we provide an overview of several topics in compiler optimization that our work did not focus on but are nonetheless important:
+Feature Extraction
+Machine learning techniques introduced in Section 2 require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering.
+The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni, Dubach, and O’Boyle 2014; Haj-Ali et al.
+2020), designed by expert intuitions.
 Numerous studies have aimed to reduce the cost of feature design.
-Following the success of word2vec embeddings in natural language processing mikolov2013distributed , methods like code2vec alon2019code2vec , inst2vec ben2018neural , and IR2vec venkatakeerthy2020ir2vec represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR).
-The surge in deep learning has enabled feeding raw information such as AST alon2018general , control-data flow graphs (CFG) brauckmann2020compiler ; cummins2020programl , and code token sequences cummins2017end into powerful deep neural networks, capable of learning useful representations end-to-end.
+Following the success of word2vec embeddings in natural language processing (Mikolov et al.
+2013), methods like code2vec (Alon et al.
+2019), inst2vec (Ben-Nun, Jakobovits, and Hoefler 2018), and IR2vec (VenkataKeerthy et al.
+2020) represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR).
+The surge in deep learning has enabled feeding raw information such as AST (Alon et al.
+2018), control-data flow graphs (CFG) (Brauckmann et al.
+2020; Cummins et al.
+2021), and code token sequences (Cummins et al.
+2017) into powerful deep neural networks, capable of learning useful representations end-to-end.
+Large Language Model
+With the rise of large language models (LLMs), recent work has introduced LLMs into compiler optimization problems, demonstrating remarkable effectiveness in predicting optimal optimization sequences and comprehending the optimization process (Cummins et al.
+2023).
+LLMs can directly analyze the intermediate representation of code in textual form, enabling a better understanding of the code’s intricate details.
+However, due to limitations on input length, it has to split programs into shorter functions for training and prediction, preventing it from predicting on code of arbitrary length.
+We also want to emphasize that optimizing individual functions is quite different from optimizing an entire program.
+Many optimization passes, such as –inline and –mergefunc, which handle function inlining and merging, are either ineffective or less critical when there’s only one function to optimize.
+Frameworks
 Platforms that expose the compiler as a playground for AI experiments have significantly reduced the entry barriers to intelligent compiler research.
-OpenTuner ansel2014opentuner and YaCoS filho2018yet serve as autotuning frameworks with a range of compiler optimization techniques.
-Our experiments utilize CompilerGym cummins2022compilergym , which offers user-friendly interfaces for researchers to interact with compilers in a reinforcement learning manner.
-We are optimistic that the future release of our trained code optimization agents, in conjunction with these platforms, can have a democratizing effect on applying AI techniques to compiler optimizations.
+OpenTuner (Ansel et al.
+2014), and YaCoS (Filho, Rodriguez, and da Silva 2018) serve as autotuning frameworks with a range of compiler optimization techniques.
+Our experiments utilize CompilerGym (Cummins et al.
+2022), which offers user-friendly interfaces for researchers to interact with compilers in a reinforcement learning manner.
+We are optimistic that the future release of our compiler world model and code optimization agents, in conjunction with these platforms, can have a democratizing effect on applying AI techniques to compiler optimizations.
 Appendix B Implementation Details
 B.1 Compiler Environment
-Our experiments are conducted on the CompilerGym platform cummins2022compilergym , version 0.2.5, with LLVM-10.0.0 integration.
+Our experiments are conducted on the CompilerGym platform (Cummins et al. 2022), version 0.2.5, with LLVM-10.0.0 integration.
 B.2 Features and Actions
-Table 2: Descriptions of 56-dimension Autophase features (adapted from cummins2022compilergym ).
 Index Name Description
@@
 55 testUnary Number of Unary operations
-We extract program features following the approach of Autophase huang2019autophase .
-As described in Section 5.1, we use a 56-dimension Autophase feature vector concatenated with a 42-dimension action histogram vector as the observation.
+Table 3: Descriptions of 56-dimension Autophase features (adapted from (Cummins et al. 2022)).
+We extract program features following the approach of Autophase (Haj-Ali et al.
+2020).
+We use a 56-dimension Autophase feature vector concatenated with a 42-dimension action histogram vector as the observation.
 The Autophase feature contains various statistics of the IR code, with each dimension thoroughly explained in Table 3.
@@
 Specifically, each element of the Autophase vector is divided by the total instruction count of the program, whereas the action histogram vector is normalized by the total action count per episode, set as 45.
-Our action space is derived from Autophase huang2019autophase , comprising 45 LLVM optimization passes.
+We choose to use this feature extraction method because it is both efficient and effective.
+In our preliminary experiments, we find that more complex feature extraction methods, such as ProGraML (Cummins et al.
+2021) and inst2vec (Ben-Nun, Jakobovits, and Hoefler 2018), significantly slow down the training and inference of CompilerDream, with only marginal improvements in performance and generalization.
+We employ two distinct action spaces in our experiments to align with the baseline methods.
+In Sections 4.2 and Section 4.3, we utilize the full action space comprising all 124 optimization passes in LLVM.
+In Section 4.4 and Section 4.5, CompilerDream and the baseline methods are trained and tested on an action space derived from Autophase (Haj-Ali et al.
+2020).
+The original version of this action space included 45 LLVM optimization passes.
 However, CompilerGym excludes 3 actions due to updates in the latest LLVM version.
@@
 Table 4 presents the names of the LLVM optimization passes corresponding to these 42 actions.
-Table 3: A list of LLVM transformation passes selected as actions.
+Using this limited action space is a prevailing practice, and it has been proved to be effective in prior works(Haj-Ali et al.
+2020; Cummins et al.
+2022).
+It also makes the training easier since the model can focus on the most effective optimization passes.
 Index Name Index Name Index Name
@@
 13 -inline 27 -loweratomic 41 -tailcallelim
-B.3 Hyperparameters
-The hyperparameters for our DreamerV3 agent implementation are outlined in Table 4. For hyperparameters not specified, we use the same value as the original DreamerV3 hafner2023mastering .
-Table 4: Hyperparameters of DreamerV3 in our experiments.
+Table 4: A list of LLVM transformation passes selected in Autophase’s action space.
+B.3 Reward Function
+The reward function is defined as the normalized change of the optimization metric $C(s)$:
+$r_{t+1}=\frac{C(s_{t})-C(s_{t+1})}{C(s_{0})-C(s_{\mathrm{b}})},$ (9)
+where a lower $C$ indicates better performance.
+In line with other works focused on optimizing code size, we define $C(s)$ as the IR instruction count at compiler state $s$.
+The $C(s_{\mathrm{b}})$ stands for the performance of a baseline policy, specifically the built-in -Oz flag in our experiments.
+A total reward greater than 1 means that the optimization sequence performs better than the baseline policy.
+B.4 Hyperparameters
+The hyperparameters for our DreamerV3 agent implementation are outlined in Table 5.
+For hyperparameters not specified, we use the same value as the original DreamerV3 (Hafner et al.
+2023).
+Most of the listed hyperparameter values are directly taken from DreamerV2 (Hafner et al.
+2021) or previous works, with minimal tuning.
+However, the loss scales are carefully tuned, as the observation loss is relatively small compared to the reward loss in our setting, and the balance between these losses has a significant impact on CompilerDream’s performance.
+All experiments share the same set of hyperparameters unless otherwise specified.
 Hyperparameter Value
@@
 Replay buffer capacity $2\times 10^{6}$
-Reward smoothing $\alpha$ lee2023dreamsmooth 0.6
+Reward smoothing $\alpha$ (Lee, Abbeel, and Lee 2024) 0.6
 Training frequency Every 5 environment steps
@@
 Gradient clipping 100
-B.4 Model-Free Baselines
-We use RLlib liang2018rllib to train and test model-free reinforcement learning algorithms including PPO schulman2017proximal , A2C mnih2016asynchronous , IMPALA espeholt2018impala , APEX horgan2018distributed , and DQN mnih2015human .
-All the experiments are conducted with 10 CPUs and an RTX-3090 GPU over a minimum duration of 10 hours.
-The training process uses 5 rollout workers to interact with the environment and use 4 evaluation workers to evaluate the checkpoint on the validation set.
-We use default hyperparameters of algorithms in RLlib following the CompilerGym platform cummins2022compilergym , except that we have carefully tuned the hyperparameters for our strongest baseline, PPO, as listed in Table 5.
-Table 5: Hyperparameters for the PPO baseline, well tuned on our dataset to be deviating from the default value in RLlib.
+Table 5: Hyperparameters of DreamerV3 in our experiments.
+B.5 Training Model-free Baselines
+We use RLlib (Liang et al.
+2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al.
+2017) (Haj-Ali et al.
+2020), A2C (Mnih et al.
+2016), IMPALA (Espeholt et al.
+2018), APEX (Horgan et al.
+2018), and DQN (Mnih et al.
+2015).
+All the experiments are conducted with 10 CPUs and an RTX-3090 GPU.
+For the PPO baseline in Section 4.4, we trained it for around 7 hours since we observed overfit to CodeContests when training for a longer time.
+Other model-free methods are trained for at least 10 hours.
+The training process uses 5 workers to interact with the environment and use 4 evaluation workers to evaluate the checkpoint on the validation set.
+We use default hyperparameters of algorithms in RLlib following the CompilerGym platform (Cummins et al.
+2022), except that we have carefully tuned the hyperparameters for PPO, as listed in Table 6.
+We explored roughly 2 or 3 values for each hyperparameter listed and selected the set of hyperparameters that yielded the best performance on the validation set.
+The results of the PPO baseline in Section 4.5 and Figure 6 are exactly the same as those in Section 4.4 and Figure 3.
 Hyperparameters Value
@@
 weight_decay 1e-6
+Table 6: Hyperparameters for the PPO baseline, well-tuned on our dataset to be deviating from the default value in RLlib.
+B.6 Autotuning
+In Section 4.2, we extensively trained CompilerDream on the cBench dataset to evaluate its effectiveness as an autotuning method in comparison to other approaches on CompilerGym’s leaderboard.
+Similar to other studies, we trained CompilerDream using all programs in cBench, excluding the one called ghostscript due to its considerable size, which would significantly hinder the training process.
+The action space is set to the full set of 124 actions in LLVM.
+All other training details remained unchanged.
+B.7 Guided Search
+The guided search strategy we use in Section 4.2 followed the design in the PPO+Guided Search baseline.
+During the search, actions are sampled from the agent’s learned policy $\pi_{\psi}(a_{t}|s_{t})$ up to 45 steps, at which point a new episode begins.
+Code size reductions after applying each action are all recorded to calculate the maximum.
+To enhance exploration, we add a uniform distribution to the policy, implemented by a $5\%$ chance of selecting a uniformly random action instead of sampling from the policy.
+At each step, the code size reduction achieved so far is recorded for later comparison, and the elapsed time is monitored to ensure that the search is completed within the 1-minute time limit per benchmark.
+Compared to the guided search in PPO+Guided Search (in Table 1), we use a horizon of 45 steps instead of 200 steps.
+Additionally, PPO+Guided Search includes an extra search of 500 steps on the best sequence found, which we do not include, resulting in our wall time being a few seconds shorter than theirs.
+B.8 Value Estimation
+Here, we provide additional details about the experiment in Section 4.3.
+The baseline method, Coreset-NVP (Liang et al.
+2023), trains a value prediction model that scores sequences from a fixed set of optimization sequences, referred to as the core set, which is obtained through an extensive search on their training set.
+For each program to optimize, it enumerates the first 45 passes in the top-scored sequences and validates their effectiveness through actual compilation.
+Specifically, sequences are processed in descending order of scores, executing each pass within a sequence.
+The compiler state (i.e., the IR) resets to its initial state after all passes in a sequence have been enumerated, before moving on to the next sequence.
+The code size reduction achieved after each pass is recorded, and the pass, along with its corresponding sequence prefix that yields the highest code size reduction, is selected.
+The core set in Coreset-NVP contains 50 sequences comprising a total of 625 passes, resulting in an average sequence length of 12.5 passes.
+Thus, roughly 3 or 4 sequences are attempted per benchmark, given the limit of 45 passes.
+These passes correspond to the optimization actions defined in Section 3.1, with a total of 79 different optimization passes included.
+For our experiment, we set the action space to the full set of all 124 optimization passes available in LLVM when training our world model.
+Focusing on predicting the target value of cumulative reward after executing each optimization pass in the core set, we replace CompilerDream’s optimization agent with a policy that sequentially outputs the sequences from the core set, allowing the world model to focus on the problem dynamics driven by these sequences.
+This setup is also comparable to the baseline, as Coreset-NVP employs a value predictor that outputs a 50-dimension vector corresponding to the normalized value of each sequence in the core set.
+That prediction model is obtained through supervised training.
+We train our world model using the CodeContests dataset introduced in Section 3.4, consistent with the experiments in Section 4.4.
+During training, we increase the weight of the reward loss to 100.0, while keeping the other hyperparameters unchanged.
+To align with the methodology used in Coreset-NVP, we validate the top 45 sequence prefixes of each program with the highest predicted values through actual compilation and select the best-performing sequence.
+The world model is trained using three different random seeds, and the average results are reported in Table 2.
+B.9 Hardware and Training Time
+We train all CompilerDream-based methods with 64 CPUs and an RTX-3090 GPU.
+In Section 4.4, we trained CompilerDream on the CodeContests dataset for around 1 day and 20 hours.
+In Sections 4.2 and 4.3, CompilerDream was trained for about 1 day.
+The random search baseline in Figure 3 is conducted with 4 CPUs, which is sufficient since it is a single-thread program.
+The random search baseline in Table 1 utilizes 80 CPUs, as specified in the write-up attached to CompilerGym’s leaderboard (Facebook 2022).
+Our CompilerDream + Guided Search method in Table 1 is tested on the same machine used for training CompilerDream, equipped with 64 CPUs and an RTX-3090 GPU.
+We ensure that sufficient memory is available for all experiments. Memory size should have little impact on the results in all of our experiments.
+B.10 Random Seeds
+All of our experiments that report a min-max range or standard deviation were replicated using three different random seeds.
+These seeds were randomly selected, and the results are generally reproducible with other seeds as well.
 Appendix C Benchmarks
@@
 The number of programs in each dataset after division is detailed in Table 7.
-Table 6: Dataset division of 8 CompilerGym benchmarks.
 Dataset Training Split Validation Split Test Split
@@
 TensorFlow 1,885 50 50
+Table 7: Dataset division of 8 CompilerGym benchmarks.
 AI-Generated Benchmarks
-To further test the generalization ability of our CodeZero agent on different programming languages, we borrow the method from FormAI tihanyi2023formai and generate a dataset containing 50 unique Objective-C programs using GPT-3.5.
+To further test the generalization ability of our CompilerDream agent on different programming languages, we borrow the method from FormAI (Tihanyi et al.
+2023) and generate a dataset containing 50 unique Objective-C programs using GPT-3.5.
 We use the same prompt as FormAI, except that we add an instruction to ask GPT to generate programs that can be directly compiled under Clang version 10.0.0 and do not use ARC (Automatic Reference Counting), to improve the compilation pass rate of generated programs.
@@
 Appendix D Extended Experimental Results
+Dataset Program Pass Sequences by CompilerDream Agent Code size
+O0 CompilerDream Oz (Reduction)
+cBench sha -sroa -gvn -sroa -sroa -instcombine -simplifycfg -sroa -functionattrs -licm -functionattrs -early-cse -simplifycfg -indvars -gvn -simplifycfg -early-cse -functionattrs -reassociate -memcpyopt -gvn -simplifycfg -reassociate -functionattrs -early-cse -instcombine -early-cse -lowerinvoke -simplifycfg -functionattrs -functionattrs -memcpyopt -gvn -simplifycfg -functionattrs -reassociate -early-cse -instcombine -early-cse -strip -strip -early-cse -strip -early-cse -strip -early-cse 799 349 500 (1.43$\times$)
+bzip2 -sroa -gvn -sroa -sroa -simplifycfg -strip -strip -instcombine -simplifycfg -early-cse -simplifycfg -strip -strip -reassociate -reassociate -memcpyopt -jump-threading -memcpyopt -functionattrs -memcpyopt -functionattrs -licm -functionattrs -gvn -simplifycfg -functionattrs -functionattrs -reassociate -memcpyopt -early-cse -jump-threading -memcpyopt -early-cse -strip -strip -instcombine -simplifycfg -jump-threading -jump-threading -early-cse -jump-threading -early-cse -simplifycfg -strip -functionattrs 28748 13565 15946 (1.18$\times$)
+OpenCV #41 -sroa -early-cse -early-cse -simplifycfg -early-cse -early-cse -early-cse -sink -sink -sink -loop-deletion -loop-deletion -loop-deletion -lowerinvoke -simplifycfg -sroa -sroa -sroa -early-cse -early-cse -early-cse -early-cse -lowerinvoke -lowerinvoke -lowerinvoke -globalopt -sroa -instcombine -simplifycfg -early-cse -early-cse -early-cse -sink -globalopt -sroa -sroa -functionattrs -simplifycfg -early-cse -functionattrs -early-cse -sink -functionattrs -functionattrs -functionattrs 28 18 28 (1.56$\times$)
+#9 -sroa -gvn -sroa -early-cse -simplifycfg -strip -strip -loop-deletion -lowerinvoke -simplifycfg -sroa -early-cse -instcombine -simplifycfg -early-cse -memcpyopt -early-cse -strip -strip -strip -functionattrs -instcombine -simplifycfg -jump-threading -simplifycfg -memcpyopt -early-cse -simplifycfg -reassociate -memcpyopt -functionattrs -functionattrs -reassociate -memcpyopt -early-cse -instcombine -simplifycfg -jump-threading -early-cse -jump-threading -strip -simplifycfg -early-cse -simplifycfg -functionattrs 9510 6341 9269 (1.18$\times$)
+TensorFlow #17 -sroa -gvn -early-cse -simplifycfg -early-cse -early-cse -instcombine -early-cse -sink -sink -sink -lowerinvoke -simplifycfg -sroa -early-cse -memcpyopt -early-cse -memcpyopt -instcombine -simplifycfg -early-cse -early-cse -early-cse -simplifycfg -early-cse -early-cse -strip -strip -indvars -simplifycfg -early-cse -memcpyopt -early-cse -instcombine -early-cse -early-cse -simplifycfg -early-cse -memcpyopt -early-cse -early-cse -instcombine -simplifycfg -early-cse -early-cse 5512 4247 5450 (1.28$\times$)
+Table 8: Top performances of the zero-shot CompilerDream agent on individual programs. Full IRs of these programs before and after optimizations are shown in the Supplementary Information.
 D.1 Learning Curves
-The zero-shot test performance of our CodeZero agents during training is shown in Figure 6.
+The zero-shot test performance of our CompilerDream agents during training is shown in Figure 7.
 Note that we validate and test the agent every 10000 environment steps and report the test performance from the checkpoint that achieved the best validation results for comparison among various methods.
-Figure 6: Zero-shot test performance of the CodeZero agent during training. We report mean and standard deviation across three runs.
+Figure 7: Zero-shot test performance of the CompilerDream agent during training. We report mean and standard deviation across three runs.
 D.2 Quantitative Results
 Quantitative results corresponding to Figure 3 in the main text are provided in Table 9.
-Table 7: Quantitative results for code size reduction, corresponding to Figure 3. We report mean and standard deviation across three runs.
--O0 Random Policy
-Random Search
-CodeZero
-(In-domain)
-CodeZero
-(Zero-shot)
-BLAS geomean 0.931 0.821$\pm$0.018 0.960$\pm$0.004 0.993$\pm$0.005 0.991$\pm$0.000
-min 0.707 0.432$\pm$0.106 0.723$\pm$0.076 0.831$\pm$0.148 0.913$\pm$0.000
-max 1.000 1.002$\pm$0.004 1.031$\pm$0.010 1.025$\pm$0.005 1.016$\pm$0.000
-cBench geomean 0.481 0.737$\pm$0.015 0.858$\pm$0.020 N/A 1.036$\pm$0.002
-min 0.299 0.401$\pm$0.097 0.560$\pm$0.030 N/A 0.962$\pm$0.001
-max 0.626 1.150$\pm$0.113 1.298$\pm$0.083 N/A 1.395$\pm$0.054
-CHStone geomean 0.487 0.701$\pm$0.043 1.037$\pm$0.006 N/A 1.094$\pm$0.005
-min 0.402 0.278$\pm$0.108 0.889$\pm$0.044 N/A 1.000$\pm$0.006
-max 0.655 1.038$\pm$0.033 1.304$\pm$0.038 N/A 1.378$\pm$0.021
-Linux geomean 0.988 0.981$\pm$0.004 1.001$\pm$0.000 0.993$\pm$0.000 0.986$\pm$0.005
-min 0.615 0.740$\pm$0.126 1.000$\pm$0.000 0.700$\pm$0.001 0.642$\pm$0.041
-max 1.011 1.001$\pm$0.002 1.011$\pm$0.000 1.011$\pm$0.000 1.009$\pm$0.003
-Mibench geomean 0.389 0.812$\pm$0.022 1.005$\pm$0.001 N/A 1.006$\pm$0.002
-min 0.278 0.352$\pm$0.074 0.858$\pm$0.010 N/A 0.879$\pm$0.003
-max 0.760 1.392$\pm$0.074 1.603$\pm$0.015 N/A 1.588$\pm$0.000
-NPB geomean 0.530 0.833$\pm$0.011 1.074$\pm$0.010 1.075$\pm$0.017 1.108$\pm$0.001
-min 0.191 0.414$\pm$0.043 0.805$\pm$0.104 0.810$\pm$0.082 0.886$\pm$0.000
-max 1.066 1.848$\pm$0.193 2.141$\pm$0.062 2.315$\pm$0.028 2.343$\pm$0.012
-OpenCV geomean 0.981 0.949$\pm$0.011 1.080$\pm$0.001 1.087$\pm$0.007 1.092$\pm$0.000
-min 0.833 0.603$\pm$0.133 0.888$\pm$0.021 0.898$\pm$0.001 0.897$\pm$0.000
-max 1.370 1.409$\pm$0.167 1.571$\pm$0.022 1.635$\pm$0.186 1.556$\pm$0.000
-TensorFlow geomean 0.983 0.912$\pm$0.009 1.006$\pm$0.003 1.032$\pm$0.000 1.032$\pm$0.001
-min 0.927 0.618$\pm$0.046 0.877$\pm$0.009 0.970$\pm$0.010 0.968$\pm$0.011
-max 1.010 1.235$\pm$0.034 1.267$\pm$0.007 1.289$\pm$0.008 1.282$\pm$0.002
-D.3 Data Distribution Visualization
+D.3 Program Case Study
+In Addition to the predicted optimization trajectory showcase in our paper, we present more example of optimization results here.
+Table 8 presents the agent’s top performance outcomes on various benchmark datasets.
+We observe that the agent indeed produces a specialized optimization strategy tailored for each program.
+Additionally, these results also highlight certain passes, such as -sroa (scalar replacement of aggregates), -early-cse(early common subexpression elimination) and -simplifycfg, as particularly effective in code size optimization.
+Figure 8: t-SNE (Van der Maaten and Hinton 2008) visualization of programs from training and test datasets.
+D.4 Data Distribution Visualization
 In Figure 8, we visualize the distribution of our training and test datasets.
@@
 These sequences are subsequently executed on each program, with the resulting Autophase features concatenated to form a feature vector with dimensions of $1000\times 45\times 56$ for every program.
-These comprehensive feature vectors are finally processed using t-SNE van2008visualizing for dimensionality reduction and visualization.
+These comprehensive feature vectors are finally processed using t-SNE (Van der Maaten and Hinton 2008) for dimensionality reduction and visualization.
 Figure 8 illustrates that our training data (denoted as triangles) has a broad coverage of test programs (denoted as circles).
-This contrasts with the Csmith dataset (denoted as crosses) employed in CompilerGym experiments cummins2022compilergym , which shows a significant deviation from real-world applications.
+This contrasts with the Csmith dataset (denoted as crosses) employed in CompilerGym experiments (Cummins et al.
+2022), which shows a significant deviation from real-world applications.
 Nonetheless, our visualization can still not perfectly capture the transferability across datasets.
 For instance, empirical evidence suggests that CodeContests are the most effective in generalizing to OpenCV and TensorFlow, while the visual analysis does not directly imply this.
-Figure 7: t-SNE van2008visualizing visualization of programs from training and test datasets.
+-O0 Random Policy
+Random Search
+Autophase
+CompilerDream
+(In-domain)
+CompilerDream
+(Zero-shot)
+BLAS geomean 0.931 0.821$\pm$0.018 0.960$\pm$0.004 0.993$\pm$0.002 0.993$\pm$0.005 0.991$\pm$0.000
+min 0.707 0.432$\pm$0.106 0.723$\pm$0.076 0.916$\pm$0.004 0.831$\pm$0.148 0.913$\pm$0.000
+max 1.000 1.002$\pm$0.004 1.031$\pm$0.010 1.014$\pm$0.001 1.025$\pm$0.005 1.016$\pm$0.000
+cBench geomean 0.481 0.737$\pm$0.015 0.858$\pm$0.020 1.010$\pm$0.004 N/A 1.036$\pm$0.002
+min 0.299 0.401$\pm$0.097 0.560$\pm$0.030 0.817$\pm$0.109 N/A 0.962$\pm$0.001
+max 0.626 1.150$\pm$0.113 1.298$\pm$0.083 1.405$\pm$0.014 N/A 1.395$\pm$0.054
+CHStone geomean 0.487 0.701$\pm$0.043 1.037$\pm$0.006 1.071$\pm$0.002 N/A 1.094$\pm$0.005
+min 0.402 0.278$\pm$0.108 0.889$\pm$0.044 0.967$\pm$0.027 N/A 1.000$\pm$0.006
+max 0.655 1.038$\pm$0.033 1.304$\pm$0.038 1.313$\pm$0.045 N/A 1.378$\pm$0.021
+Linux geomean 0.988 0.981$\pm$0.004 1.001$\pm$0.000 0.985$\pm$0.001 0.993$\pm$0.000 0.986$\pm$0.005
+min 0.615 0.740$\pm$0.126 1.000$\pm$0.000 0.615$\pm$0.000 0.700$\pm$0.001 0.642$\pm$0.041
+max 1.011 1.001$\pm$0.002 1.011$\pm$0.000 1.009$\pm$0.003 1.011$\pm$0.000 1.009$\pm$0.003
+Mibench geomean 0.389 0.812$\pm$0.022 1.005$\pm$0.001 1.000$\pm$0.007 N/A 1.006$\pm$0.002
+min 0.278 0.352$\pm$0.074 0.858$\pm$0.010 0.833$\pm$0.062 N/A 0.879$\pm$0.003
+max 0.760 1.392$\pm$0.074 1.603$\pm$0.015 1.524$\pm$0.045 N/A 1.588$\pm$0.000
+NPB geomean 0.530 0.833$\pm$0.011 1.074$\pm$0.010 1.071$\pm$0.018 1.075$\pm$0.017 1.108$\pm$0.001
+min 0.191 0.414$\pm$0.043 0.805$\pm$0.104 0.801$\pm$0.070 0.810$\pm$0.082 0.886$\pm$0.000
+max 1.066 1.848$\pm$0.193 2.141$\pm$0.062 2.235$\pm$0.101 2.315$\pm$0.028 2.343$\pm$0.012
+OpenCV geomean 0.981 0.949$\pm$0.011 1.080$\pm$0.001 1.082$\pm$0.004 1.087$\pm$0.007 1.092$\pm$0.000
+min 0.833 0.603$\pm$0.133 0.888$\pm$0.021 0.897$\pm$0.002 0.898$\pm$0.001 0.897$\pm$0.000
+max 1.370 1.409$\pm$0.167 1.571$\pm$0.022 1.620$\pm$0.193 1.635$\pm$0.186 1.556$\pm$0.000
+TensorFlow geomean 0.983 0.912$\pm$0.009 1.006$\pm$0.003 1.010$\pm$0.019 1.032$\pm$0.000 1.032$\pm$0.001
+min 0.927 0.618$\pm$0.046 0.877$\pm$0.009 0.841±0plus-or-minus0$\pm$ 0± 0.134 0.970$\pm$0.010 0.968$\pm$0.011
+max 1.010 1.235$\pm$0.034 1.267$\pm$0.007 1.281$\pm$0.006 1.289$\pm$0.008 1.282$\pm$0.002
+Table 9: Quantitative results for code size reduction, corresponding to Figure 3. We report mean and standard deviation across three runs.

### v3

@@
 CompilerDream: Learning a Compiler World Model for General Code Optimization
-Abstract
+Abstract.
 Effective code optimization in compilers is crucial for computer and software engineering.
 The success of these optimizations primarily depends on the selection and ordering of the optimization passes applied to the code.
-While most compilers rely on a fixed sequence of optimization passes, current methods to find the optimal sequence either employ impractically slow search algorithms or learning methods that struggle to generalize to code unseen during training.
-We introduce CompilerDream, a model-based reinforcement learning approach to general code optimization.
-CompilerDream comprises a compiler world model that accurately simulates the intrinsic properties of optimization passes and an agent trained on this model to produce effective optimization strategies.
-By training on a large-scale program dataset, CompilerDream is equipped to serve as a general code optimizer across various application scenarios and source-code languages.
-Our extensive experiments first highlight CompilerDream’s strong optimization capabilities for autotuning, where it leads the CompilerGym leaderboard.
-More importantly, the zero-shot generalization ability of large-scale trained compiler world model and agent, excels across diverse datasets, surpassing LLVM’s built-in optimizations and other state-of-the-art methods in both settings of value prediction and end-to-end code optimization.
-1 Introduction
+While most compilers rely on fixed pass sequences, current methods to find the optimal sequence for specific programs either employ impractically slow search algorithms or learning methods that struggle to generalize to code unseen during training.
+To address these challenges, we introduce CompilerDream, the first world-model-based approach for general code optimization.
+CompilerDream features a compiler world model with a reward smoothing technique, enabling accurate simulation of optimization processes.
+Built on this model, code optimization agents can then be constructed via value prediction or direct optimization sequence generation.
+Trained on a large-scale program dataset, these agents serve as versatile code optimizers across diverse application scenarios and source-code languages.
+Our extensive experiments highlight CompilerDream’s strong optimization capabilities for autotuning, where it leads the CompilerGym leaderboard.
+More importantly, the zero-shot generalization ability of large-scale trained compiler world model and agent, excels across diverse datasets, surpassing LLVM’s built-in optimizations and state-of-the-art methods in both settings of value prediction and end-to-end code optimization.
+[^1] Code and data available at https://github.com/thuml/CompilerDream.
+World Models, Compiler Optimization, Code Data Mining
+copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736887isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Machine learningccs: Software and its engineering Compilersccs: Computing methodologies Model development and analysis
+KDD Availability Link:
+The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15552746.
+The data of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15549673.
+Figure 1.
+CompilerDream performs code optimization by interacting with the compiler using an agent powered by a world model.
+With strong generalization capabilities, it efficiently optimizes programs from diverse origins.
+[…]
+Diverse source code, including C++, C, Rust, and other languages, is first translated into an intermediate representation (IR) by the compiler frontend.
+CompilerDream, leveraging a world model, optimizes the IR by selecting an optimal pass at each step using an optimization agent.
+The optimized IR is then processed by the compiler backend to generate a high-quality executable file.
+1. Introduction
 Code optimization plays an important role in realizing the full potential of software and hardware.
 Developers desire a universal solution to transform input programs into semantically equivalent but more efficient versions without manual effort.
-To achieve this, compilers usually employ a front-end that translates the source code of any language into an intermediate representation (IR), a back-end that converts IR to the binary code optimizer, and a middle-end optimizer that performs language- and platform-agnostic optimizations over the IR of the original code (Figure 1).
-The optimizer is typically implemented as a series of passes applying transforms on the code.
-The performance of the optimization primarily depends on the selection and order of the optimization passes.
-Standard compilers use a few fixed sets of optimization sequences to enhance specific aspects of program performance.
-For example, -O1, -O2, and -O3 improve for execution speed, and -Os and -Oz for reducing program size.
-Obviously, given the vast diversity of programs and platforms, these off-the-shelf optimization strategies predefined by compiler experts are not optimal for most circumstances, and we can significantly improve the performance of a specific program if we can find a better optimization pass sequence compared to that provided by compilers (Triantafyllis et al.
-2003; Georgiou et al.
-2018).
-Figure 1: CompilerDream agent performs code optimization using a learned policy in a compiler environment.
-Guided by a learned policy, at each step, CompilerDream analyzes the Intermediate Representation (IR) of the program, selects to apply an optimization pass, and receives a reward based on the improvement in optimization metrics.
-Through sequential interactions, the agent aims to maximize cumulative rewards and enhance the final IR’s performance.
-Automatically optimizing the pass sequence is thus crucial in improving the compiler optimization performance.
+Compilers achieve this through a front-end that translates source code into an intermediate representation (IR), a middle-end optimizer that performs language- and platform-agnostic IR optimizations, and a back-end that converts IR to the binary code (Figure 1).
+The optimizer is typically implemented as a series of passes applying transforms on the code, where performance largely depends on the selection and order of these optimization passes.
+Standard compilers use a few fixed sets of optimization sequences to enhance specific aspects of program performance, such as -O1, -O2, and -O3 for execution speed, and -Os and -Oz for program size reduction.
+Obviously, given the vast diversity of programs and platforms, these off-the-shelf strategies predefined by compiler experts are suboptimal for most circumstances.
+Automatically optimizing the pass sequence for specific programs thus can yield significant performance gains over default compiler settings (Triantafyllis et al., 2003; Georgiou et al., 2018).
 To be practical, such an algorithm must produce a satisfactory pass sequence within a reasonable time and handle a wide variety of programs.
 However, current research often fails to meet these requirements simultaneously.
-Search-based methods (Bodin et al.
-1998) ensure near-optimality but require thousands of compilations for each new program, making it too time-consuming for regular use.
-In contrast, machine learning methods can either predict the optimization sequences directly or estimate the outcomes of applying an optimization sequence to guide a search, thereby avoiding the time-consuming process of interacting with the compiler to validate the optimization outcome.
-However, in addition to potentially sacrificing some optimality compared to searching, these learning-based methods face a significant bottleneck in broad generalization across diverse programs that may be out of the training samples.
-A range of machine learning applications (Devlin et al.
-2019; Brown et al.
-2020; Radford et al.
-2021; Kirillov et al.
-2023) have witnessed that training high-capacity models on large-scale datasets can yield unprecedented performance.
-Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al.
-2022).
-However, the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al.
-2020; Shahzad et al.
-2022) or from relatively small training sets comprising only a few hundred programs (Mammadli, Jannesari, and Wolf 2020; Jain et al.
-2022).
-Additionally, the predominant use of model-free reinforcement learning methods with limited capacity may also hinder generalization.
-In this study, we focus on the LLVM (Lattner and Adve 2004) phase ordering problem, a longstanding challenge for compiler research, and propose CompilerDream, a world model-based reinforcement learning approach capable of generating an optimization pass sequence tailored to any particular input program.
-Our model-based method (Hafner et al.
-2023) not only learns an accurate predictive world model of compilers to simulate real compiler executions, whose number can be greatly reduced, but also benefits generalization by learning representations that better capture the structure of the compiler state transitions (Anand et al.
-2022).
-Trained on a carefully curated large-scale dataset of natural programs, the world model accurately captures the dynamics of the compiler optimization process, enabling the optimization agent to generalize to unseen programs and predict superior optimization sequences.
-We demonstrate the effectiveness of CompilerDream on a range of domains (Cummins et al.
-2022), from benchmark suites encompassing fundamental algorithms to production-level open-source programs, including object files from C++ TensorFlow (Abadi et al.
-2016) and OpenCV (Culjak et al.
-2012) library.
-Even without considering generalization, CompilerDream’s strong optimization capabilities place it at the top of the CompilerGym leaderboard for autotuning.
-For general code optimization, the large-scale trained world models of CompilerDream can accurately predict the outcomes of pass sequences on unseen programs, a setting known as value prediction.
-Beyond selecting optimization sequences based on value prediction, the agent can also directly produce optimization sequences in a single trial, yielding more efficient code size reduction compared to both the -Oz flag and state-of-the-art methods.
-The main contributions of this work are three-fold:
-  • We introduce CompilerDream, the first model-based reinforcement learning approach for code optimization.
-  • We first leverage a large-scale dataset, CodeContests, to achieve strong generalization on unseen programs.
-  • Extensive experiments spanning autotuning, value prediction, and end-to-end code optimization, demonstrate the superior performance of CompilerDream.
-2 Related Work
-One of the key challenges for compilation is to determine which code transformations to apply, how to apply them (e.g., using suitable parameters), and in what order.
-This requires effectively searching and evaluating a massive number of possible options, known as iterative compilation (Bodin et al.
-1998) or autotuning (Datta et al.
-2008).
+Search-based methods (Bodin et al., 1998) achieve near-optimality but require thousands of compilations per program to validate the optimization outcomes, making them impractical.
+In contrast, machine learning methods avoid these time-consuming compiler interactions, by either predicting the optimization sequences directly or estimating the outcomes of optimization sequences to guide a search.
+However, these learning-based methods still risk sacrificing some optimality and, more importantly, face a significant bottleneck in broad generalization across diverse programs that may be out of the training samples.
+A range of machine learning applications (Devlin et al., 2019; Brown et al., 2020; Radford et al., 2021; Kirillov et al., 2023) have witnessed that training high-capacity models on large-scale datasets can yield unprecedented performance.
+Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al., 2022), yet the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al., 2020; Shahzad et al., 2022) or from relatively small training sets comprising only a few hundred programs with limited-capacity models (Mammadli et al., 2020; Jain et al., 2022), which hinders generalization.
+This paper focuses on the LLVM (Lattner and Adve, 2004) phase ordering problem, a longstanding challenge for compiler research.
+We propose CompilerDream, a world-model-based approach for general code optimization, capable of handling a wide variety of programs, unlike most prior approaches focusing on narrow, domain-specific ones.
+It learns an accurate predictive world model to simulate compiler executions, forecasting future IR states and optimization metric improvements based on the current IR state and applied pass.
+We believe that employing a world model offers the following advantages for compiler optimization: First, by capturing optimization dynamics, the world model gains generalizable knowledge into our method (Anand et al., 2022), improving the overall generalization performance.
+Second, it replaces costly compiler invocations, significantly reducing computational overheads of search and learning algorithms and facilitating large-scale training.
+Additionally, its high-capacity architecture can extract deeper insights from extensive datasets.
+Specifically, to better adapt the world model for compiler optimization, we closely examine the optimization process and introduce a reward smoothing technique to enhance world model training and improve simulation accuracy.
+Built on this accurate world model, CompilerDream supports various types of optimization agents, including value prediction and reinforcement learning.
+Moreover, we carefully curate a large-scale training dataset of natural programs, enhancing the world model’s and agent’s generalization to unseen programs, enabling our method to predict superior optimization sequences.
+We demonstrate CompilerDream’s effectiveness across a range of program domains (Cummins et al., 2022) and problem scenarios.
+Our test domains include benchmark suites covering fundamental algorithms, as well as production-level open-source programs, such as object files from C++ TensorFlow (Abadi et al., 2016) and OpenCV (Culjak et al., 2012) libraries.
+Even without considering generalization, CompilerDream’s strong optimization capabilities top the CompilerGym leaderboard for autotuning.
+For general code optimization, our large-scale trained world model accurately predicts the outcomes of pass sequences on unseen programs.
+Both types of optimization agents equipped with the world model outperform the built-in -Oz flag and state-of-the-art methods in their respective settings.
+In the value prediction setting, the world model serves as an accurate value predictor, enabling superior action sequence selection.
+In the reinforcement learning setting, the agent trained entirely in the world model directly generates optimization sequences in a single trial, achieving more efficient code across diverse datasets.
+The main contributions of this work are as follows:
+  • We propose CompilerDream, the first world-model-based approach for code optimization.
+  • We introduce a reward smoothing technique that facilitates the application of world models to compiler optimization, benefiting future research in this area.
+  • Our approach supports multiple optimization agents, including value prediction and reinforcement learning agents, unifying different methods in this field.
+  • By leveraging the large-scale CodeContests dataset and the world model’s generalization capability, we achieve strong performance on diverse unseen programs.
+  • Extensive experiments across various program domains and problem scenarios, including autotuning, value prediction, and end-to-end code optimization, show that CompilerDream outperforms state-of-the-art methods.
+2. Related Work
+A key challenge in compilation is determining which code transformations to apply, how to apply them (e.g., using suitable parameters), and in what order.
+This involves effectively searching and evaluating numerous options, a process known as iterative compilation (Bodin et al., 1998) or autotuning (Datta et al., 2008).
 However, this search-based approach only finds a good optimization for one specific program and does not generalize into a compiler strategy.
 This limitation underscores the importance of integrating machine learning techniques.
-Supervised Learning
-Pioneering work has delved into supervised machine learning, adopting two main approaches (Leather and Cummins 2020).
-The first approach requires an extensive search on each training programs to identify the most effective optimization sequence, which then serves as the data labels.
-An early example (Calder et al.
-1997) used a neural network for branch prediction, and one more well-known work is MilepostGCC (Fursin et al.
-2008), a practical attempt to integrate machine learning into a production compiler, GCC.
+Supervised learning.
+Pioneering work has delved into supervised machine learning, adopting two main approaches (Leather and Cummins, 2020).
+The first approach requires an extensive search for each training program to identify the most effective optimization sequence, which then serves as the data labels.
+An early example (Calder et al., 1997) used a neural network for branch prediction, and one more well-known work is MilepostGCC (Fursin et al., 2008), a practical attempt to integrate machine learning into a production compiler, GCC.
 It employs models trained on a large dataset of programs distributed over the Internet.
-The second approach aims to learn a cost or performance function capable of estimating the quality of various compiler options, which enables evaluation of a range of possible options without the need to compile and profile each one (Stephenson et al.
-2003; Luk, Hong, and Kim 2009).
-Reinforcement Learning
-Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the requirement for collecting optimal labeled data (Kulkarni and Cavazos 2012).
-This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al.
-2021), loop transformation (Haj-Ali et al.
-2020; Brauckmann, Goens, and Castrillón 2021), and graph partitioning (Mirhoseini et al.
-2017).
-Several works relevant to us have explored the full optimization pipeline, i.e.
-the LLVM phase ordering problem, including AutoPhase (Haj-Ali et al.
-2020), CORL (Mammadli, Jannesari, and Wolf 2020), and POSET-RL (Jain et al.
-2022).
-Model-based Reinforcement Learning
-While existing work on reinforcement learning for code optimization predominantly focuses on model-free RL methods, our work pioneers the use of an advanced model-based RL approach.
-Model-based RL (Sutton and Barto 2018) can offer advantages in terms of both sample efficiency and generalization.
-Executing and profiling extensive optimization sequences, especially for the runtime metric, can be time-consuming.
-This is further compounded when constructing complex observations, such as control-data flow graphs (Brauckmann et al.
-2020; Cummins et al.
-2021).
-Model-based RL addresses these challenges by learning a world model that approximates state transitions and reward signals of the compiler.
-This is related in spirit to the aforementioned approach of supervised performance models but goes further by also learning a policy through simulating trajectories based on model predictions.
-This approach thus improves sample efficiency not only by eliminating the necessity for a guided search process during deployment but also by reducing trial-and-error interactions with the real compiler environment during training.
-Moreover, as the policy can share the representation with the world model, model learning can act as an auxiliary task and thus aid in learning representations that better capture the structure of the environment and manifest in better generalization of the policy (Anand et al.
-2022; Mazoure et al.
-2022; Ma et al.
-2024).
-Figure 2: Design overview of CompilerDream.
-(a) In the model-based training loop, the CompilerDream agent interacts with the compiler environment across diverse training programs, learns a world model from historical experience, and updates its policy efficiently through model simulations.
-(b) The trained agent can be deployed and is capable of zero-shot generalization to unseen programs, delivering effective optimizations.
-(c) The compiler environment is set up with the input program’s IR, $s_{0}$.
-At each step, upon receiving an optimization pass $a_{t}$, the environment executes the pass internally, resulting in a transformed IR $s_{t+1}$, and provides the agent with relevant IR features $o_{t+1}$ and immediate rewards $r_{t+1}$ based on optimization metric improvements.
-(d) The world model simulates the compiler environment, initiated with an observation, maintaining its internal neural states, and predicting future observations and rewards in response to input actions.
-(e) The policy shares with the world model a generalizable representation that captures the environment’s structure.
-3 Method
-In this section, we first present the mathematical formulation of the code optimization problem (Section 3.1).
-Then we introduce the CompilerDream approach for both compiler world model learning (Section 3.2) and optimization agent learning (Section 3.3).
-Finally, we elaborate on considerations of curating large-scale training dataset (Section 3.4).
-3.1 Phase Ordering Decision Process
-As illustrated in Figure 1, one key problem of compiler optimization is to find the optimal sequence of optimization passes for a given program, which is also referred to as the phase ordering problem.
-It can be naturally formulated as a partially observable Markov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\mu,\mathcal{O},\phi)$ (Sutton and Barto 2018).
+The second approach aims to learn a cost or performance function capable of estimating the quality of various compiler options, which enables the evaluation of possible options without the need to compile and profile each one (Stephenson et al., 2003; Luk et al., 2009).
+Coreset-NVP (Liang et al., 2023) follows this approach and achieves state-of-the-art performance.
+Reinforcement learning.
+Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the need for collecting optimal labeled data (Kulkarni and Cavazos, 2012).
+This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al., 2021), loop transformation (Haj-Ali et al., 2020; Brauckmann et al., 2024), and graph partitioning (Mirhoseini et al., 2017).
+Several works relevant to us, including AutoPhase (Haj-Ali et al., 2020), CORL (Mammadli et al., 2020), and POSET-RL (Jain et al., 2022), have explored the full optimization pipeline, i.e., the phase ordering problem, using model-free RL without incorporating a world model.
+World model.
+A world model (Ha and Schmidhuber, 2018; LeCun, 2022) that approximates state transitions and reward signals is typically used in two ways: (1) it enables simulation of ”unseen” interactions that are unavailable or unaffordable (Janner et al., 2019; Hafner et al., 2020); (2) as an auxiliary learning task, it aids in better representation that captures the underlying structure of the environment (Anand et al., 2022; Mazoure et al., 2022; Ma et al., 2024).
+To the best of our knowledge, we are the first to introduce world models for code optimization.
+In this context, a world model can function as a compiler simulator, approximating IR transformations and eliminating the need for costly execution and profiling of extensive optimization sequences.
+Furthermore, by sharing representations with the world model, the policy can generalize more effectively to unseen programs.
+Thus, model-based agents (Sutton and Barto, 2018) can be implemented, which offer superior sample efficiency and generalization compared to model-free methods.
+3. Method
+This section first defines the code optimization problem and design choices for observation, action, and reward (Section 3.1).
+We then present the design and training technique of CompilerDream’s world model (Section 3.2), two optimization agents (Section 3.3), and considerations for large-scale dataset curation (Section 3.4).
+3.1. Phase Ordering Decision Process
+As illustrated in Figure 1, one key problem of compiler optimization is to find the optimal sequence of optimization passes for a given program, also known as the phase ordering problem.
+It can be naturally formulated as a partially observable Markov decision process (POMDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},r,p,\mu,\mathcal{O},\phi)$ (Sutton and Barto, 2018).
 The state space $\mathcal{S}$ covers all possible Intermediate Representations (IRs), the action space $\mathcal{A}$ comprises individual compiler optimization passes, and the reward function $r$ is defined by the metric being optimized.
@@
 The initial state distribution $\mu\in\Delta(\mathcal{S})$ captures all IRs of interest, which can be approximated via uniform sampling from the training dataset.
-The observation function $\phi:\mathcal{S}\mapsto\mathcal{O}$ maps the underlying IR into the observation space.
-Complex observation spaces offer comprehensive program information, while expert-designed features (Haj-Ali et al.
-2020) embed problem-specific knowledge, potentially enhancing the optimization policy’s generalization by eliminating irrelevant details.
-A code optimization agent guided by policy $\pi$ determines the optimization sequence of an input program through a series of interactions with the compiler environment.
-At each time step $t=0,1,2,\dots$, the agent applies an action $a_{t}$ to current IR based on observation $o_{t}$ and its policy $\pi(a_{t}\mid o_{\leq t})$, which transforms the current state $s_{t}$ to the next state $s_{t+1}=p(s_{t},a_{t})$ and will receive a reward $r_{t}$.
-The goal of the agent is to learn an optimization policy $\pi$ that effectively maximizes the cumulative rewards $G=\sum_{t}r_{t}$, thereby optimizing the performance metrics of the final IR.
-3.2 Learning a Compiler World Model
-We develop our CompilerDream approach by adapting an advanced model-based RL method, Dreamer (Hafner et al.
-2023), as depicted in Figure 2.
-This approach involves learning a predictive world model $(\hat{p}_{\theta},\hat{r}_{\theta})$ of the compiler environment parameterized by $\theta$ that approximates the underlying transition dynamics $p(o_{t+1}|o_{\leq t},a_{\leq t})$ and reward function $r(o_{\leq t},a_{\leq t})$.
+The observation function $\phi:\mathcal{S}\mapsto\mathcal{O}$ maps the underlying IR into the observation space, capturing useful features.
+A code optimization agent determines the optimization sequence for an input program through interactions with the compiler environment.
+At each time step $t=0,1,2,\dots$, the agent applies an action $a_{t}$ to current IR based on observation $o_{t}$, transforming current state to $s_{t+1}=p(s_{t},a_{t})$ and receiving reward $r_{t}$.
+The agent can employ effective strategies, such as search algorithms and parametric neural networks, to select the best actions, as detailed in Section 3.3.
+Under the POMDP formulation, we design the observation and action space as summarized in Table 1.
+The observation features were selected for efficiency and effectiveness.
+In preliminary experiments, we find that complex program features like ProGraML (Cummins et al., 2021) and inst2vec (Ben-Nun et al., 2018) significantly slow down CompilerDream with marginal performance gains.
+In contrast, expert-designed Autophase features (Haj-Ali et al., 2020) leverage domain knowledge, enhancing generalization by filtering irrelevant details.
+We construct the observation by concatenating the 56-dimensional Autophase feature vector with a 42-dimensional action histogram vector, which records the number of times each action has been selected during the current episode.
+Both vectors are normalized for consistency: each Autophase feature is divided by the program’s initial total instruction count, and the action histogram is scaled by the per-episode action limit, which is set to 45.
+We adopt two distinct action spaces in our experiments to align with the baseline methods.
+The full action space consists of all 124 LLVM optimization passes, while the reduced action space derived from Autophase (Haj-Ali et al., 2020) consists of 42 actions.
+Originally, this action space included 45 LLVM passes, but CompilerGym excludes 3 due to updates in the latest LLVM version, leaving 42 actions.
+This reduced action space is widely used and has been shown to be effective in prior studies (Haj-Ali et al., 2020; Cummins et al., 2022).
+The reward function is defined as the normalized change of the optimization metric $C(s)$:
+(1) $r_{t+1}=\frac{C(s_{t})-C(s_{t+1})}{C(s_{0})-C(s_{\mathrm{b}})},$
+where lower $C$ indicates better performance.
+Following prior work on code size reduction, we define $C(s)$ as the IR instruction count.
+$C(s_{\mathrm{b}})$ is the baseline performance achieved by the built-in -Oz flag.
+Table 1. Observation and action space in CompilerDream.
+Aspect Name Description Dim.
+Observation
+Autophase
+(Haj-Ali et al., 2020)
+A feature vector capturing
+various IR code statistics.
+56
+Action
+Histogram
+A vector where each
+dimension represents the
+execution frequency of a
+specific action.
+124
+/42
+Action
+(Sec. 4.2 & 4.3)
+Full LLVM
+Passes
+Full action space with all
+LLVM optimization passes.
+124
+Action
+(Sec. 4.4)
+Autophase
+Passes
+A reduced action space
+derived from Autophase(Haj-Ali et al., 2020).
+42
+Figure 2.
+Design overview of CompilerDream.
+The world model is trained on real episodes generated by the agent, incorporating a reward smoothing process to stabilize training and enhance learning efficiency.
+CompilerDream supports training with RL and value prediction agents, which learn from episodes simulated by the world model.
+[…]
+The agent interacts with a compiler to collect real episodes, which are stored in a replay buffer.
+While training the world model with these episodes, a reward smoothing process is applied to enhance training stability and learning efficiency.
+Once trained, the world model functions as a compiler environment simulator, enabling reinforcement learning (RL) agents and value prediction agents to interact with it and generate simulated episodes.
+These simulated episodes are used to train the agents, ultimately leading to improved performance.
+3.2. Learning a Compiler World Model
+Following the advanced Dreamer approach (Hafner et al., 2023), we build a world model to learn the formulated POMDP process of compiler optimization, as depicted in Figure 2.
+Concretely, we train a model $(\hat{p}_{\theta},\hat{r}_{\theta})$ of the compiler environment parameterized by $\theta$, which approximates the underlying transition dynamics of optimization process $p(o_{t+1}|o_{\leq t},a_{\leq t})$ and the reward function $r(o_{\leq t},a_{\leq t})$.
 The compiler world model simulating the compiler environment is formulated with the following four components:
-Representation model: $z_{t}\sim q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})$ (1)
-Transition model: $\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})$
-Observation decoder: $\hat{o}_{t}\sim p_{\theta}(\hat{o}_{t}\mid z_{t})$
+(2) Representation model: $z_{t}\sim q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t}),$
+Transition model: $\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1}),$
+Observation decoder: $\hat{o}_{t}\sim p_{\theta}(\hat{o}_{t}\mid z_{t}),$
 Reward decoder: $\hat{r}_{t}\sim p_{\theta}(\hat{r}_{t}\mid z_{t}).$
@@
 A representation loss $\mathcal{L}_{\mathrm{repr}}$ is used to train the neural compiler states to accurately reconstruct the observation and reward by two decoders:
-$\mathcal{L}_{\mathrm{repr}}(\theta)\doteq-\ln p_{\theta}(o_{t}|z_{t})-\ln p_{\theta}(r_{t}|z_{t}).$ (2)
-The transition model captures the dynamics of the compiler world model, predicting future neural state ${\hat{z}}_{t}$ directly from $z_{t-1}$ and $a_{t-1}$.
-A prediction loss minimizes the difference between the estimated state $z_{t}$ and the predicted state ${\hat{z}}_{t}$, simultaneously enhancing the transition model’s accuracy in predicting future states and making the state produced by the representation model easier to predict:
-$\mathcal{L}_{\mathrm{pred}}(\theta)\doteq\mathrm{KL}\left[q_{\theta}(z_{t}|z_{t-1},a_{t-1},o_{t})\,\|\,p_{\theta}(\hat{z}_{t}|z_{t-1},a_{t-1})\right].$ (3)
+(3) $\mathcal{L}_{\mathrm{repr}}(\theta)\doteq-\ln p_{\theta}(o_{t}|z_{t})-\ln p_{\theta}(r_{t}|z_{t}).$
+The transition model captures the dynamics of the compiler world model, predicting future neural compiler state ${\hat{z}}_{t}$ directly from $z_{t-1}$ and $a_{t-1}$.
+A prediction loss minimizes the difference between the estimated neural compiler state $z_{t}$ and the predicted compiler state ${\hat{z}}_{t}$, simultaneously enhancing the transition model’s accuracy in predicting future states and making the neural compiler state produced by the representation model easier to predict:
+(4) $\mathcal{L}_{\mathrm{pred}}(\theta)\doteq\mathrm{KL}\left[q_{\theta}(z_{t}|z_{t-1},a_{t-1},o_{t})\,\|\,p_{\theta}(\hat{z}_{t}|z_{t-1},a_{t-1})\right].$
 The overall models are jointly learned by minimizing the sum of the representation loss and the prediction loss.
@@
 Predicted real compiler state $\hat{o}_{\tau+1}$ can be optionally reconstructed by the observation decoder.
+Figure 3. Comparison of the original and smoothed rewards during an episode on a program from the BLAS dataset (Lawson et al., 1979).
+[…]
+Taking a typical episode as an example: out of 45 actions, only 8 receive non-zero rewards before smoothing, and they are scattered throughout the optimization process.
+After smoothing, most rewards become non-zero, and abrupt changes between adjacent rewards are eliminated.
+Moreover, the overall trend remains consistent with the original reward distribution.
 Reward smoothing
-During an optimization sequence, not all passes can effectively transform the code, and the improvements gradually saturate.
-We employ a reward smoothing technique (Lee, Abbeel, and Lee 2024) to mitigate the sparsity and the long tail distribution of rewards during an episode.
-This is achieved through the following:
-$r_{t}^{\prime}\leftarrow\alpha r_{t-1}^{\prime}+(1-\alpha)r_{t},\;\;t=1,2,\dots$ (4)
-with $\alpha\in[0,1)$. Consequently, we train a reward decoder $p_{\theta}(\hat{r}_{t}^{\prime}\mid z_{t})$ to predict the smoothed rewards.
-3.3 Learning a Code Optimization Agent
-We then train a code optimization agent on top of the world model, which consists of an actor and a critic neural networks parameterized on top of the neural compiler state:
+We discover that most optimization passes in an optimization sequence do not change the program IR instruction count, leading to sparse rewards.
+Despite this, these passes are essential as they may modify critical properties, such as instruction order or replacements, that affect the effectiveness of subsequent passes.
+Furthermore, improvements tend to saturate over time, as large non-zero rewards typically appear early in the sequence and diminish as optimization progresses.
+In Figure 3, we present a typical episode of the optimization process: out of 45 actions, only 8 yielded non-zero rewards, whose values decrease as optimization progresses.
+These properties pose challenges for training the agent effectively, as it receives little to no guidance from most actions, even when some are vital to the final result.
+Moreover, the world model may struggle to predict reward values accurately, often defaulting to zero.
+This challenge arises from the significant class imbalance between zero and non-zero rewards, making it difficult to determine both the timing and magnitude of non-zero rewards.
+Therefore, we applied reward smoothing to mitigate their sparsity and long-tailed distribution within an episode. This is achieved by adding an exponential decay to rewards:
+(5) $r_{t}^{\prime}\leftarrow\alpha r_{t-1}^{\prime}+(1-\alpha)r_{t},\;\;t=1,2,\dots$
+with $\alpha\in[0,1)$.
+Consequently, we train a reward decoder $p_{\theta}(\hat{r}_{t}^{\prime}\mid z_{t})$ to predict the smoothed rewards.
+As shown in Figure 3, the smoothed rewards provide non-zero feedback for most steps, making them easier for CompilerDream to learn.
+It is worth noting that the total reward of an episode remains approximately unchanged after smoothing.
+For all $t\geq 1$, its contribution to the total reward after smoothing approximates the original reward value if the horizon is infinite: $\lim_{n\to\infty}(1-\alpha)\sum_{i=1}^{n}\alpha^{i}r_{t}=r_{t}$.
+Since most episodes exhibit a long-tailed reward distribution and we use a relatively small smoothing factor ($\alpha=0.6$), the smoothed total reward is expected to closely match the original total reward.
+3.3. Learning a Code Optimization Agent
+Building on the world model, we implement two agent designs, unifying supervised and reinforcement learning methods for compiler optimization into our world-model-based approaches.
+Value prediction agent
+This kind of agent adopts a classic approach in compiler optimization, employing a heuristic to guide search methods (Liang et al., 2023; Luk et al., 2009; Stephenson et al., 2003).
+It comprises a value prediction model $v_{\theta}$ to estimate the effect of an optimization sequence and a search strategy that leverages this model to identify the best sequence.
+The value of a program state $s$ is defined as the expected cumulative reward of applying an action sequence to the program.
+This aligns with the ratio of the reduced IR instruction count achieved between applying action sequence $\{a_{\tau}\}_{\tau=1}^{m}$ and standard $-Oz$ flag, according to the reward function $r$ defined in Eq.
+(1):
+(6) $v(s_{0},\{a_{\tau}\})=\sum_{\tau=1}^{m}r_{\tau}=\sum_{\tau=1}^{m}\frac{C(s_{\tau-1})-C(s_{\tau})}{C(s_{0})-C(s_{b})}=\frac{C(s_{0})-C(s_{m})}{C(s_{0})-C(s_{b})}.$
+We leverage our world model to build a value prediction model.
+As shown Section 3.2, from an initial observation $o_{0}$, and an action sequence $\{a_{\tau}\}$, our world model can simulate an optimization trajectory $\{\hat{z}_{\tau},\hat{a}_{\tau},\hat{r}_{\tau},\hat{o}_{\tau}\}_{\tau=1}^{m}$.
+Summing up all simulated rewards $\hat{r}_{\tau}$ yields the predicted value:
+(7) $v_{\theta}(s_{0},\{a_{\tau}\})=\sum_{\tau=1}^{m}\hat{r}_{\tau}\approx\sum_{\tau=1}^{m}r_{\tau}$
+The search strategy then seeks the optimal optimization sequence $\{a_{\tau}\}^{*}$ by evaluating candidates using the value prediction model $v_{\theta}$.
+Although various search algorithms are applicable, we employ a simple strategy that enumerates action sequences within a fixed search space (detailed in Section 4.3).
+Reinforcement learning agent
+RL is another popular approach to compiler optimization, learning a policy to select passes sequentially.
+Unlike prior methods that rely on costly compiler interactions, CompilerDream trains RL agents on world model-simulated compiler optimization trajectories, significantly improving efficiency.
+Our RL agent comprises an actor and a critic neural networks, both parameterized based on the neural compiler state:
 […]
-The critic evaluates the cumulative optimization return that the actor can achieve starting from a given neural state, also known as value or return, denoted as $v_{\xi}(\hat{z}_{t})$.
-This is trained by minimizing the difference between the predicted value $v_{\xi}(\hat{z}_{t})$ and the Monte-Carlo return or more advanced bootstrapped $\lambda$-return (Sutton and Barto 2018):
-$\mathcal{L}_{\text{critic}}(\xi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum^{t+H}_{\tau=t}-\log v_{\xi}(V_{\tau}^{\lambda}\mid\hat{z}_{\tau})\right],$ (6)
-$V_{\tau}^{\lambda}\doteq\hat{r}_{\tau}+\gamma\begin{cases}(1-\lambda)v_{\xi}(\hat{z}_{\tau+1})+\lambda V_{\tau+1}^{\lambda}&\text{if}\ \tau<t+H\\ v_{\xi}(\hat{z}_{\tau+1})&\text{if}\ \tau=t+H.\end{cases}$
-The actor produces a policy $\pi_{\psi}$ that predicts the best optimization pass to choose, represented as an action distribution, which is trained to maximize the simulated return through the REINFORCE policy gradient (Williams 1992):
-$(\displaystyle\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big{[}\sum^{t+H}_{\tau=t}\Big{(}$$-\left(V_{\tau}^{\lambda}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}|\hat{z}_{\tau})$
-$-\eta\,\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\Big{)}\Big{]},$ (7)
-where $\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]$ is an entropy regularization which encourages exploration, and $\eta$ is a hyperparameter that adjusts the regularization strength.
-Both the actor and critic are completely trained on world model-simulated compiler optimization trajectories.
-3.4 Data Curation
+Figure 4.
+Results of general code optimization with RL: Code size reduction in terms of IR instruction count over LLVM -Oz under different methods.
+Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
+[…]
+The performance of several methods (-O0, Random Policy, -Oz, Random Search, PPO, In-domain CompilerDream, and Zero-shot CompilerDream) on different benchmarks (BLAS, cBench, CHStone, Linux, MiBench, NPB, OpenCV and TensorFlow).
+The critic evaluates the $\gamma$-discounted value $v_{\xi}(\hat{z}_{t})$ of simulated neural state $\hat{z}_{t}$ under policy $\pi_{\psi}$.
+It is trained by minimizing the difference between the predicted value $v_{\xi}(\hat{z}_{t})$ and the Monte-Carlo or more advanced bootstrapped return (Sutton and Barto, 2018), denoted as $V_{t}$:
+(9) $\mathcal{L}_{\text{critic}}(\xi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum^{t+H}_{\tau=t}-\log v_{\xi}(V_{\tau}\mid\hat{z}_{\tau})\right],$
+The actor produces an optimization policy $\pi_{\psi}$ that predicts an action distribution of the best pass to choose, which is trained to maximize the simulated return through the REINFORCE policy gradient (Williams, 1992) with an entropy regularization (Haarnoja et al., 2019):
+$(\displaystyle\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big{[}\sum^{t+H}_{\tau=t}\Big{(}$$-\left(V_{\tau}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}\mid\hat{z}_{\tau})$
+(10) $-\eta\,\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\Big{)}\Big{]},$
+3.4. Data Curation
 To facilitate that our CompilerDream method can effectively generalize to unseen situations, a concept known as zero-shot generalization, we have identified three critical factors in preparing our training dataset.
 First, the dataset should reflect naturalness.
-During our preliminary experiments, we found that those generated code datasets like Csmith (Yang et al.
-2011) and llvm-stress (Lattner and Adve 2004) provide no benefits or even hurt the generalization to real-world scenarios.
+During our preliminary experiments, we found that those generated code datasets like Csmith (Yang et al., 2011) and llvm-stress (Lattner and Adve, 2004) provide no benefits or even hurt the generalization to real-world scenarios.
 Second, the dataset should be large to prevent the agent from overfitting to a small number of codes and failing to generalize.
@@
 We need data with complex algorithmic logic and potential for optimizations to help the world model and the agent better understand the problem’s intricacies.
-Datasets like AnghaBench (Da Silva et al.
-2021) which consist of millions of human-written codes collected from GitHub, are often too simple to allow for significant optimization improvements.
-Therefore, we choose to construct our training datasets on top of the CodeContests dataset released with AlphaCode (Li et al.
-2022), which consists of over 13,000 problems of coding competition, and each problem, on average, has hundreds of solutions in multiple languages.
+Datasets like AnghaBench (da Silva et al., 2021) which consist of millions of human-written codes collected from GitHub, are often too simple to allow for significant optimization improvements.
+Table 2. Comparison of different training datasets for compiler optimization.
+Dataset
+Name
+Number of
+IR files
+Large-
+scale
+Human-
+written
+Code
+quality
+cBench (Fursin et al., 2007) 23 No Yes High
+Mibench (Guthaus et al., 2001) 40 No Yes Low
+Csmith (Yang et al., 2011) $\infty$ Yes No High
+llvm-stress (Lattner and Adve, 2004) $\infty$ Yes No Low
+AnghaBench (da Silva et al., 2021) $1,041,333$ Yes Yes Low
+CodeContests (Li et al., 2022) $110,240$ Yes Yes High
+Therefore, we choose to construct our training datasets on top of the CodeContests dataset released with AlphaCode (Li et al., 2022), which consists of over 13,000 problems of coding competition, and each problem, on average, has hundreds of solutions in multiple languages.
 We subsample up to ten C++ solutions for each training problem, resulting in 110,240 programs, as our training data, and sample one solution for each of 100 test problems as our validation data.
-4 Experiments
-We evaluate CompilerDream in three different settings and compare its performance with prior state-of-the-art methods.
-We first demonstrate that, without considering generalization, CompilerDream excels as a powerful autotuning method, leading the CompilerGym leaderboard (Section 4.2).
-Then we show that CompilerDream can learn a general compiler world model capable of accurately estimating and selecting pass sequences (Section 4.3) and finally learn a general optimization agent that directly generates a pass sequence tailored to any specific input program.
-(Section 4.4)
-4.1 Evaluation
-Our experiments focus on code size reduction, which benefits applications targeting low-resource hardware such as embedded systems.
-This focus is driven by the practical advantages of code size as a metric: it is cost-effective to construct extensive compilable training and test datasets and to evaluate the optimization performance for code size.
+As shown in Table 2, CodeContests meets all our criteria, whereas other datasets fall short in one or more aspects.
+4. Experiments
+We conducted extensive experiments across various settings, comparing CompilerDream against state-of-the-art methods. Our study aims to answer the following key research questions (RQs):
+  1. RQ 1
+    Optimality: Can CompilerDream’s joint training process of the compiler world model and optimization agent discover superior optimization sequences?
+  2. RQ 2
+    Accuracy: Does CompilerDream’s world model produce accurate simulations of the real optimization process?
+  3. RQ 3
+    Generalization: Is the policy learned with the world model still effective on various unseen programs?
+  4. RQ 4
+    Effectiveness: Are all the techniques employed in CompilerDream effective in enhancing optimization results?
+In the following sections, we first show that CompilerDream excels as a powerful autotuning method, leading the CompilerGym leaderboard by discovering superior optimization sequences and leveraging accurate world model (Section 4.2), addressing RQ 1 and partially RQ 2.
+We then demonstrate that its value prediction agent (Section 4.3) and RL agent (Section 4.4) both outperform state-of-the-art methods in various unseen test datasets, addressing RQ 2 and RQ 3.
+Finally, Section 4.5 presents ablation studies for RQ 4.
+4.1. Evaluation
+Our experiments focus on code size reduction, benefiting applications on low-resource hardware like embedded systems.
+This focus stems from the practical advantages of code size as a metric: It is cost-effective to construct compilable training and test datasets and to evaluate the optimization performance for code size.
 Metric
 To be more robust to outliners, we evaluate the code size optimization results by the geometric mean of IR instruction count reduction of program $s$ in a dataset $\mathcal{D}$:
-$R_{\mathcal{D}}(\pi)=\left(\prod_{s\in\mathcal{D}}\frac{C(s_{Oz})}{C(s_{t}^{\pi})}\right)^{\frac{1}{|\mathcal{D}|}}$ (8)
-where $C$ counts the number of IR-instructions, $s_{Oz}$ is the IR optimized by LLVM -Oz flag, and $s_{t}^{\pi}$ is the final IR produced by the algorithm $\pi$.
-A value above $1$ indicates that the algorithm performs better than -Oz option on average.
+(11) $R_{\mathcal{D}}(\text{agent})=\left(\prod_{s\in\mathcal{D}}\frac{C(s_{b})}{C(s_{\text{agent}})}\right)^{\frac{1}{|\mathcal{D}|}}$
+where $C$ denotes the IR instruction count, $s_{b}$ is the IR state optimized by LLVM’s -Oz flag serving as a baseline, and $s_{\text{agent}}$ is the final IR state produced by agent.
+A value of $R$ above $1$ indicates superior performance compared to LLVM’s -Oz option.
 Benchmarks
-We evaluate our method mainly on benchmarks from the CompilerGym platform (Cummins et al.
-2022): benchmark suites including cBench (Fursin et al.
-2007), CHStone (Hara et al.
-2008), MiBench (Guthaus et al.
-2001), and NASA Parallel Benchmarks (NPB) (Bailey et al.
-1995), in addition to kernels from open source projects including BLAS (Lawson et al.
-1979), Linux, OpenCV (Culjak et al.
-2012), and TensorFlow (Abadi et al.
-2016).
-We disregard benchmarks generated by synthetic program generators (Lattner and Adve 2004; Yang et al.
-2011) since they do not align with real-world scenarios.
-We follow the standard data splits of CompilerGym.
-Additionally, we evaluate CompilerDream on a large-scale dataset named FormAI (Tihanyi et al.
-2023).
+We evaluate our method mainly on benchmarks from the CompilerGym platform (Cummins et al., 2022): benchmark suites including cBench (Fursin et al., 2007), CHStone (Hara et al., 2008), MiBench (Guthaus et al., 2001), and NASA Parallel Benchmarks (NPB) (Bailey et al., 1995), as well as kernels from open source projects such as BLAS (Lawson et al., 1979), Linux, OpenCV (Culjak et al., 2012), and TensorFlow (Abadi et al., 2016).
+Synthetic benchmarks from program generators (Lattner and Adve, 2004; Yang et al., 2011) are excluded as they lack real-world relevance.
+We adhere to the standard data splits of CompilerGym.
+For benchmarks with a total number of programs more than 100, we use the first 50 programs as the test set, the following 50 programs as the validation set, and all of the rest as the training set.
+These training and validation sets are only used for in-domain training.
+The datasets comprising fewer than 100 programs are not applicable for in-domain training; instead, all their programs are allocated to the test set.
+The number of programs in each dataset after division is detailed in Table 3.
+Table 3. Dataset division of 8 CompilerGym benchmarks.
+Benchmark Training Split Validation Split Test Split
+BLAS 200 50 50
+cBench N/A N/A 23
+CHStone N/A N/A 12
+Linux 13,794 50 50
+MiBench N/A N/A 40
+NPB 22 50 50
+OpenCV 342 50 50
+TensorFlow 1,885 50 50
+Additionally, we evaluate CompilerDream on a large-scale dataset named FormAI (Tihanyi et al., 2023).
 FormAI comprises a vast collection of AI-generated C programs with diverse functionalities and coding styles.
@@
 Inspired by FormAI, we also constructed a dataset of 50 Objective-C programs generated by a Large Language Model to further evaluate CompilerDream’s ability to generalize to different programming languages.
-All details on benchmarks can be found in Appendix C.
-Figure 3: Generel code optimization results: Code size reduction in terms of IR instruction count over LLVM -Oz under different methods.
-Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
-4.2 Autotuning: CompilerGym Leaderboard
+Details can be found in Appendix B.8.
+4.2. Autotuning: CompilerGym Leaderboard
+Table 4. Autotuning results on cBench, i.e., the CompilerGym leaderboard (Facebook, 2022), where learning-based methods report only inference time following leaderboard convention.
 Method Walltime
@@
 GATv2 + DD-PPO 258.1s 1.047$\times$
-Table 1: Autotuning results: CompilerDream’s code size reduction and time consumption on cBench dataset compared with other methods from CompilerGym leaderboard (Facebook 2022).
-Results for methods other than ours are taken directly from the leaderboard, where learning-based methods report only inference time as wall time.
-We first use CompilerDream as an autotuning method to verify that its training process has sufficient capability to discover high-quality optimization sequences, supporting our goal of training an effective general code optimization agent.
+We first validate CompilerDream as an autotuning method to demonstrate its ability to discover high-quality optimization sequences, supporting our goal of achieving superior code optimization.
 Implementation
-We target the task of CompilerGym leaderboard (Facebook 2022), optimizing pass sequences across 23 cBench programs.
-CompilerDream is trained for 25 hours, averaging about one hour per program—a time budget comparable to search methods.
-After training, besides evaluating single-trial optimization performance (more in Section 4.4), we test CompilerDream+Guided Search, following the leaderboard’s top method.
-This approach uses the agent’s policy $\pi_{\psi}(a_{t}|s_{t})$ to guide random search and limits search time to 1 minute per program (details in Appendix B.7).
+We target the CompilerGym leaderboard task (Facebook, 2022), optimizing pass sequences for 23 cBench programs.
+Following the common setup adopted by other methods on the leaderboard, we train CompilerDream using our RL agent on all cBench programs except ghostscript, which is excluded due to its large size that would significantly slow down training.
+The action space is set to the full action space of all 124 actions in LLVM.
+CompilerDream is trained for 25 hours, averaging about one hour per program—comparable to the 3600-second wall time of the random search algorithm.
+Beyond evaluating single-trial optimization performance, we test CompilerDream+Guided Search, inspired by the leaderboard’s leading approach.
+This method leverages the RL agent’s policy $\pi_{\psi}(a_{t}|s_{t})$ to guide random search, limiting search time to 1 minute per program (details in Appendix B.6).
 Results
-As shown in Table 1, CompilerDream’s policy achieves an average 1.068×1.068$\times$1.068 × code size reduction on cBench with a single trial after training, surpassing random search methods that require significantly more time—even when accounting for CompilerDream’s training time.
-This highlights the effectiveness of CompilerDream’s model-based RL training process in finding superior optimization passes, making it a promising component in building a general code optimizer.
-Moreover, when combined with guided search, CompilerDream outperforms the top method on the leaderboard while using less wall time.
-4.3 General Value Prediction
-We then demonstrate that the large-scale trained world model in CompilerDream can generalize to accurately simulate the optimization process of unseen input programs, thereby supporting the training of a general optimization agent.
-To illustrate this, we utilize our trained world model to estimate the outcome of applying an optimization sequence, a task named value prediction.
+As shown in Table 4, our RL agent trained on world model simulations achieves an average 1.068×1.068$\times$1.068 × code size reduction on cBench in a single trial, surpassing random search methods despite their longer execution times, even when accounting for CompilerDream’s training time.
+This demonstrates the world model’s ability to accurately simulate optimization processes and support effective agent learning.
+Furthermore, when combined with guided search, CompilerDream outperforms the top leaderboard method while using less wall time.
+4.3. General Value Prediction
+In this section, we demonstrate that CompilerDream’s large-scale trained world model accurately simulates the optimization process of unseen programs, enabling the construction of an excellent value prediction agent that surpasses state-of-the-art methods.
 Implementation
-We adopt the state-of-the-art Coreset-NVP method (Liang et al.
-2023), which evaluates pass sequences within a fixed candidate set (the coreset with 625 sequence prefixes) and selects the best prefix via compiler validation.
-Coreset-NVP employs a neural network to directly predict the normalized value of sequences, which is learned via supervised learning.
-Our Coreset-CompilerDream method leverages a world model trained on CodeContests, to simulate the optimization process and accumulate predicted rewards as the value.
-Both our method and the baseline method Coreset-NVP are tested on four datasets orthogonal to their training sets.
-The number of compiler validation calls is set the same across methods to ensure a fair comparison.
-Additional implementation details are available in Appendix B.8.
+To ensure a fair comparison of the value prediction capabilities, we adopt the setting of the state-of-the-art Coreset-NVP method (Liang et al., 2023), which evaluates pass sequences from a fixed set of 50 distinct sequences (the core set).
+We train our world model to build a value prediction agent (detailed in Section 3.3) that predicts the value of applying sequences from the same core set.
+We follow the same search strategy as Coreset-NVP, enumerating the prefixes of roughly 3 or 4 sequences with the highest predicted value and selecting the best by compiler validation.
+Both methods are evaluated on four datasets distinct from their training sets, with an identical number of compiler validation calls to ensure a fair comparison.
+Additional details are provided in Appendix B.7.
+Table 5. General value prediction results: The geometric mean of code size reduction achieved by the best sequences selected by different methods across 4 datasets.
 Method Dataset
@@
 Coreset-Oracle 1.041 1.106 1.020 1.159
-Table 2: General value prediction results: The geometric mean of code size reduction achieved by the best sequences selected by different methods across 4 datasets.
-Results of Coreset-NVP and Coreset-Oracle is reported from Coreset-NVP’s original paper (Liang et al.
-2023).
 Results
-Table 2 shows the geometric mean reduction achieved by different methods.
-Despite the test datasets being distinct from the training set, CompilerDream outperforms Coreset-NVP across three datasets, highlighting its strong compiler modeling and generalization capabilities.
+Table 5 presents the geometric mean reduction achieved by different methods.
+Despite being tested on datasets distinct from the training set, CompilerDream surpasses Coreset-NVP on three datasets, demonstrating superior compiler simulation and generalization capabilities.
 On MiBench, CompilerDream performs nearly optimally compared to the Oracle baseline, which uses a brute-force search and serves as the problem’s upper bound.
 Both our method and Coreset-NVP achieve an average reduction of $1.101$ on CHStone, close to the upper bound of $1.06$.
-4.4 General Code Optimization
-We finally aim to demonstrate that CompilerDream can learn a general optimization agent, capable of end-to-end generating effective optimization sequences for a wide variety of programs unseen during training.
-This scenario represents a typical daily use case where the optimization algorithm has limited time to compute the best solution for each of the diverse programs.
-Consequently, all methods generate only one optimization sequence per program in the test datasets unless otherwise specified.
-Figure 4: Histograms of large-scale evaluations comparing CompilerDream and PPO on the FormAI dataset.
-Figure 5: A comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model.
+4.4. General Code Optimization with RL
+We finally demonstrate the ability of CompilerDream’s reinforcement learning agent, which can generate optimization sequences end-to-end in a single trial for a wide variety of programs unseen during training.
+This scenario mirrors real-world use cases where the optimization algorithm has limited time to compute the best sequence for arbitrary programs.
+Therefore, all methods in this section generate only one optimization sequence per program in the test datasets unless otherwise specified.
+Figure 5.
+A comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model.
 The learned world model accurately captures the variations of program features and optimization metrics.
-Figure 6: Analysis.
+[…]
+This figure is the comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model.
+Both trajectories start from the “cBench/stringsearch2” program and follow the same sequence of optimization passes (-sora, -early-cse, -simplifycfg, $\cdots$).
+Evaluating key metrics such as reward and intermediate representation (IR) features (BranchCount, BlockLow, TotalBlocks, $\cdots$), the world model’s predictions closely align with the real environment.
+After the first three passes, the mean squared error (MSE) remains low at 0.00071, 0.00088, and 0.00092, demonstrating the model’s accuracy in capturing optimization dynamics.
+Figure 6.
+Analysis.
 Evaluations of different RL algorithms: (a) Learning curves of various RL algorithms, measured by the geometric mean of code size reduction on the CodeContests validation set.
 A Gaussian filter ($\sigma=2.0$) is applied to enhance the visualization of trends.
-(b) Zero-shot generalization capabilities of different RL algorithms on various test benchmark datasets.
+(b) Generalization capabilities of different RL algorithms on various test datasets.
 Effect of training dataset: (c) Test performance of CompilerDream trained on CodeContests and Csmith.
 Bars indicate the standard deviation.
+[…]
+(a)The performance of CompilerDream steadily and consistently improves throughout training, and achieves a validation code size reduction of 1.02 after about $4\times 10^{4}$ environment steps.
+In contrast, other methods (PPO, A2C, IMPALA, DQN, and APEX) all achieve reductions of less than 0.98 at the same stage.
+To reach a reduction of 1, PPO needs nearly $2\times 10^{5}$ steps–roughly six times the steps required by CompilerDream.
+The remaining four methods exhibit unstable performance and fail to exceed a reduction of 1 even after $10^{6}$ steps.
+(b) The generalization capabilities of different RL algorithms (CompilerDream, PPO, A2C, IMPALA, DQN, and APEX) on five test datasets unseen during training (cBench, CHStone, NPB, OpenCV, and TensorFlow).
+CompilerDream outperforms all other methods across all five datasets, demonstrating a clear advantage.
+PPO consistently ranks second, though with a noticeable performance gap compared to CompilerDream.
+The remaining four methods perform poorly, particularly on the cBench and CHStone datasets.
+(c) The impact of the training dataset is demonstrated by comparing CompilerDream trained on two datasets: CodeContests and Csmith.
+Across all five benchmarks (cBench, CHStone, NPB, OpenCV, and TensorFlow), CompilerDream trained on CodeContests consistently outperforms its counterpart trained on Csmith.
+Notably, the performance gap even reaches a geometric mean code size reduction difference of 0.2 on the CHStone benchmark.
+Figure 7. Histograms of large-scale evaluations comparing CompilerDream and PPO on the FormAI dataset.
+[…]
+The performance of CompilerDream and PPO was evaluated on the large-scale FormAI test.
+For both methods, the code size reduction for most test cases (approximately $5\times 10^{4}$) falls within the range of $0.8\textasciitilde{}1.2$.
+As the reduction magnitude increases or decreases beyond this range, the number of corresponding cases gradually decreases.
+Comparing CompilerDream with PPO, we observe that CompilerDream optimizes more cases with higher code size reduction while producing fewer cases with worsened results.
+Notably, CompilerDream achieves a reduction of more than nine times for some cases, surpassing all previous studies.
 Implementation
-We compare CompilerDream with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method.
-The random search conducts hundreds of trials within a time budget similar to CompilerDream’s.
+We train CompilerDream with its RL agent, and compare it with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method (Haj-Ali et al., 2020) using Proximal Policy Optimization (PPO) (Schulman et al., 2017).
+The random search conducts hundreds of trials within a time budget similar to our RL agent.
 The -Oz flag represents LLVM’s highest level of code size optimization while the -O0 flag represents no optimization.
-The learning-based method using Proximal Policy Optimization (PPO) (Schulman et al.
-2017) and CompilerDream are trained on the large-scale CodeContests dataset and zero-shot generalized to unseen test programs.
-Following the state-of-the-art approach, all methods operate within a limited action space.
+The learning-based method using PPO and CompilerDream are trained on the large-scale CodeContests dataset and zero-shot generalized to unseen test programs.
+Following the state-of-the-art approach, all methods use the reduced action space (Table 1).
 Results
-Figure 3 shows CompilerDream’s performance in reducing code size, measured by the geometric mean of IR instruction count reduction.
-Without specific in-domain training, CompilerDream outperforms -Oz in all but two test benchmarks in a single trial and surpasses PPO in all but the BLAS datasets.
-It also exceeds random search within a similar time budget on most datasets, except Linux.
-The minimal performance differences on the BLAS and Linux datasets suggest they are already highly optimized.
-Moreover, CompilerDream’s zero-shot generalization matches or exceeds in-domain training.
-This advantage is especially notable in the NPB dataset, where data sparsity challenges in-domain agents, yet CompilerDream achieves an additional $3\%$ code size reduction.
-CompilerDream also generalizes well to new languages, evidenced in its performances on the BLAS and NPB datasets of Fortran programs.
-Further, on the AI-generated Objective-C dataset, CompilerDream achieves an average code size reduction of 1.027$\times$, reaching up to 2.87$\times$ in certain instances.
+Figure 4 presents the code size reduction achieved by CompilerDream’s RL agent, measured by the geometric mean of IR instruction count reduction.
+Without in-domain training, CompilerDream surpasses -Oz in all but two benchmarks in a single trial and consistently outperforms PPO, except on the BLAS datasets.
+It also outperforms random search on most datasets within a comparable time budget, except for Linux.
+The minimal performance differences across various methods on BLAS and Linux suggest these datasets are already highly optimized.
+Moreover, CompilerDream’s zero-shot generalization matches or surpasses in-domain training.
+This advantage is particularly evident in the NPB dataset, where data sparsity limits in-domain agents, yet CompilerDream achieves an additional 3% code size reduction.
+Its robust generalization extends to new languages, as shown by results on the Fortran-based BLAS and NPB datasets.
+On the AI-generated Objective-C dataset, CompilerDream achieves an average code size reduction of 1.027$\times$, reaching up to 2.87$\times$ in certain test cases.
 On the large-scale FormAI test set (Figure 7), CompilerDream outperforms PPO, matching or outperforming -Oz on more programs and achieving higher optimization levels, demonstrating its superior and consistent performance.
-4.5 Analysis
+4.5. Analysis
 Comparison with model-free methods
-We further evaluate the sample efficiency and zero-shot generalization abilities of CompilerDream based on the world model algorithm, Dreamer, against a range of model-free counterparts, including PPO (Schulman et al.
-2017), DQN (Mnih et al.
-2015), A2C (Mnih et al.
-2016), APEX (Horgan et al.
-2018), and IMPALA (Espeholt et al.
-2018; Schulman et al.
-2017).
-Figure 6a illustrates that while PPO is the strongest model-free baseline, CompilerDream outperforms it, learning an order of magnitude faster with fewer compiler interactions.
-As Figure 6b demonstrates, CompilerDream also excels in generalizing to unseen benchmarks, supporting our hypothesis that world model-based agents better capture the environment’s structure and enhance zero-shot generalization.
+We further evaluate the sample efficiency and zero-shot generalization abilities of our model-based CompilerDream (with RL agent) against various model-free counterparts, including PPO (Schulman et al., 2017), DQN (Mnih et al., 2015), A2C (Mnih et al., 2016), APEX (Horgan et al., 2018), and IMPALA (Espeholt et al., 2018; Schulman et al., 2017).
+Figure 6a illustrates that while PPO is the strongest model-free baseline, our RL agent trained on world model simulation learns an order of magnitude faster with fewer compiler interactions.
+Figure 6b further highlights its superior generalization to unseen benchmarks, supporting our claim that world model-based agents better capture the dynamics of compiler optimization and enhance generalization to unseen datasets.
+Comparison with model-based methods
+To demonstrate the necessity and effectiveness of learning a compiler world model with deep representations of the compiler.
+We compare our approach with a classical model-based RL method, MBPO (Janner et al., 2019), under the RL agent setting described in Section 4.4.
+MBPO adopts standard MLPs for its dynamics model and does not share a latent representation space with the actor or critic.
+For a fair comparison, we apply reward smoothing to MBPO using the same smoothing factor as in our method.
+The results are shown in Table 6.
+While MBPO outperforms the best model-free baseline, PPO, on some tasks, it consistently underperforms CompilerDream, except on MiBench, where the performance gap across all methods remains marginal.
+These results suggest that model-based RL alone is insufficient, and that both a more expressive world model and a shared latent space are crucial for better optimization performance and generalization to unseen programs.
+Table 6. Comparison between CompilerDream with MBPO and PPO for general code optimization.
+Dataset PPO MBPO CompilerDream
+BLAS 0.993 0.988 0.991
+cBench 1.010 1.020 1.036
+CHStone 1.071 1.066 1.094
+MiBench 0.985 0.988 0.986
+NPB 1.000 0.996 1.006
+Linux 1.071 1.069 1.108
+OpenCV 1.082 0.998 1.092
+TensorFlow 1.010 0.996 1.032
+Effect of reward smoothing
+To evaluate the effectiveness of the reward smoothing technique described in Section 3.2, we compare the performance of the RL agent trained with the complete CompilerDream method against a variant without reward smoothing.
+Table 7 presents these results alongside PPO for reference.
+The results demonstrate that reward smoothing consistently enhances optimization performance across all datasets and enables CompilerDream to outperform PPO on MiBench, Linux, and OpenCV.
+Table 7. Comparison between CompilerDream with and without reward smoothing for general code optimization.
+Dataset PPO
+Ours w/o
+Reward Smoothing
+Ours w/
+Reward Smoothing
+BLAS 0.993 0.987 0.991
+cBench 1.010 1.021 1.036
+CHStone 1.071 1.076 1.094
+MiBench 1.000 0.995 1.006
+NPB 1.071 1.093 1.108
+Linux 0.985 0.980 0.986
+OpenCV 1.082 1.080 1.092
+TensorFlow 1.010 1.022 1.032
 Effect of training dataset
-To assess the impact of the CodeContests dataset on the generalization ability, we compared it with the commonly used Csmith (Yang et al.
-2011) dataset, a large LLVM IR dataset generated by rules.
-We train CompilerDream on both datasets and the results shown in Figure 6 indicate that CompilerDream trained on CodeContests significantly outperforms the Csmith-trained version across all five test datasets.
-This advantage is particularly evident on the manually curated cBench (Fursin et al.
-2007) and CHStone (Hara et al.
-2008) datasets, demonstrating that the CodeContests-trained model can generalize more effectively to human-written programs.
+To assess the impact of the CodeContests dataset on the generalization ability, we compared it with the commonly used Csmith (Yang et al., 2011) dataset, a large LLVM IR dataset generated by rules.
+We train CompilerDream with RL agent on both datasets and the results shown in Figure 6 indicate that CompilerDream trained on CodeContests significantly outperforms the Csmith-trained version across all five test datasets.
+This advantage is particularly evident in the manually curated cBench (Fursin et al., 2007) and CHStone (Hara et al., 2008) datasets, demonstrating that the CodeContests-trained model can generalize more effectively to human-written programs.
 Program showcase
@@
 This instance exemplifies the capability of our learned compiler world model to serve as a viable alternative for a real compiler environment in training code optimization agents.
-5 Discussion
-We aim to tackle the major challenge of zero-shot generalization that arises when applying learning-based code optimization.
-We thus introduce the CompilerDream approach that leverages the sample efficiency and generalization capabilities of world model-based reinforcement learning, as well as the strengths of a diverse, high-quality training program dataset.
-Our results have shown that CompilerDream, with its powerful generalization ability, attains optimization performance that outperforms existing methods, in multiple problem scenarios under the challenging area of phase ordering.
+5. Discussion
+We aim to address the major challenge of generalization in learning-based code optimization by introducing the CompilerDream approach, which leverages the simulation and generalization capabilities of a world model.
+By incorporating a reward smoothing technique and a large-scale training program dataset, we enable effective world model training for compiler optimization tasks.
+Our method supports two types of optimization agents: a value prediction agent and an RL agent.
+Experimental results demonstrate that CompilerDream achieves superior optimization performance across diverse problem scenarios and program datasets, surpassing built-in compiler optimization flags and state-of-the-art methods.
+Limitation
+Although our method can naturally extend to optimization objectives such as execution time and object file size reduction, we focus solely on code size optimization for scalability and stability, as discussed in Section 4.1.
+Execution time measurements often exhibit high variance and biases due to hardware and operating system states, introducing significant noise that additional sampling cannot mitigate.
+Furthermore, existing open-source frameworks, such as CompilerGym, lack robust support for accurate runtime measurement, making it challenging to obtain reliable signals for effective learning.
+As a result, we only focus on code size reduction for this paper.
+Future Works
 There is substantial scope for further exploration, including expansion of the training dataset, scaling up the compiler world model, optimizing multiple objectives like execution time, and enriching feature and action spaces with deeper expert knowledge or large language models.
+Additionally, our approach could support more types of optimization agents, such as search agents leveraging advanced tree search algorithms (Schrittwieser et al., 2020) based on world model simulations.
+6. Acknowledgments
+This work was supported by the National Natural Science Foundation of China (62021002).
 Appendix A Extended Related Work
@@
 Machine learning techniques introduced in Section 2 require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering.
-The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni, Dubach, and O’Boyle 2014; Haj-Ali et al.
-2020), designed by expert intuitions.
+The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni et al., 2014; Haj-Ali et al., 2020), designed by expert intuitions.
 Numerous studies have aimed to reduce the cost of feature design.
-Following the success of word2vec embeddings in natural language processing (Mikolov et al.
-2013), methods like code2vec (Alon et al.
-2019), inst2vec (Ben-Nun, Jakobovits, and Hoefler 2018), and IR2vec (VenkataKeerthy et al.
-2020) represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR).
-The surge in deep learning has enabled feeding raw information such as AST (Alon et al.
-2018), control-data flow graphs (CFG) (Brauckmann et al.
-2020; Cummins et al.
-2021), and code token sequences (Cummins et al.
-2017) into powerful deep neural networks, capable of learning useful representations end-to-end.
+Following the success of word2vec embeddings in natural language processing (Mikolov et al., 2013), methods like code2vec (Alon et al., 2019), inst2vec (Ben-Nun et al., 2018), and IR2vec (VenkataKeerthy et al., 2020) represent programs as distributed vectors that capture syntactic and semantic information from the abstract syntax tree (AST) or intermediate representation (IR).
+The surge in deep learning has enabled feeding raw information such as AST (Alon et al., 2018), control-data flow graphs (CFG) (Brauckmann et al., 2020; Cummins et al., 2021), and code token sequences (Cummins et al., 2017) into powerful deep neural networks, capable of learning useful representations end-to-end.
 Large Language Model
-With the rise of large language models (LLMs), recent work has introduced LLMs into compiler optimization problems, demonstrating remarkable effectiveness in predicting optimal optimization sequences and comprehending the optimization process (Cummins et al.
-2023).
+With the rise of large language models (LLMs), recent work has introduced LLMs into compiler optimization problems, demonstrating remarkable effectiveness in predicting optimal optimization sequences and comprehending the optimization process (Cummins et al., 2023).
 LLMs can directly analyze the intermediate representation of code in textual form, enabling a better understanding of the code’s intricate details.
@@
 Platforms that expose the compiler as a playground for AI experiments have significantly reduced the entry barriers to intelligent compiler research.
-OpenTuner (Ansel et al.
-2014), and YaCoS (Filho, Rodriguez, and da Silva 2018) serve as autotuning frameworks with a range of compiler optimization techniques.
-Our experiments utilize CompilerGym (Cummins et al.
-2022), which offers user-friendly interfaces for researchers to interact with compilers in a reinforcement learning manner.
+OpenTuner (Ansel et al., 2014), and YaCoS (Filho et al., 2018) serve as autotuning frameworks with a range of compiler optimization techniques.
+Our experiments utilize CompilerGym (Cummins et al., 2022), which offers user-friendly interfaces for researchers to interact with compilers in a reinforcement learning manner.
 We are optimistic that the future release of our compiler world model and code optimization agents, in conjunction with these platforms, can have a democratizing effect on applying AI techniques to compiler optimizations.
 Appendix B Implementation Details
-B.1 Compiler Environment
-Our experiments are conducted on the CompilerGym platform (Cummins et al. 2022), version 0.2.5, with LLVM-10.0.0 integration.
-B.2 Features and Actions
+B.1. Compiler Environment
+Our experiments are conducted on the CompilerGym platform (Cummins et al., 2022), version 0.2.5, with LLVM-10.0.0 integration.
+B.2. Details of Features and Actions
+As shown in Table 1, we adpot the 56-dimensional Autophase feature and a reduced action space from Autophase (Haj-Ali et al., 2020).
+A detailed list of the Autophase features and actions space is available in Table 8 and Table 9.
+Table 8. Descriptions of 56-dimension Autophase features (adapted from (Cummins et al., 2022)).
 Index Name Description
@@
 55 testUnary Number of Unary operations
-Table 3: Descriptions of 56-dimension Autophase features (adapted from (Cummins et al. 2022)).
-We extract program features following the approach of Autophase (Haj-Ali et al.
-2020).
-We use a 56-dimension Autophase feature vector concatenated with a 42-dimension action histogram vector as the observation.
-The Autophase feature contains various statistics of the IR code, with each dimension thoroughly explained in Table 3.
-The action histogram vector contains the counts of actions already taken by the agent within the current episode.
-Both vectors are normalized to ensure that their values fall within a reasonable range.
-Specifically, each element of the Autophase vector is divided by the total instruction count of the program, whereas the action histogram vector is normalized by the total action count per episode, set as 45.
-We choose to use this feature extraction method because it is both efficient and effective.
-In our preliminary experiments, we find that more complex feature extraction methods, such as ProGraML (Cummins et al.
-2021) and inst2vec (Ben-Nun, Jakobovits, and Hoefler 2018), significantly slow down the training and inference of CompilerDream, with only marginal improvements in performance and generalization.
-We employ two distinct action spaces in our experiments to align with the baseline methods.
-In Sections 4.2 and Section 4.3, we utilize the full action space comprising all 124 optimization passes in LLVM.
-In Section 4.4 and Section 4.5, CompilerDream and the baseline methods are trained and tested on an action space derived from Autophase (Haj-Ali et al.
-2020).
-The original version of this action space included 45 LLVM optimization passes.
-However, CompilerGym excludes 3 actions due to updates in the latest LLVM version.
-Thus, we utilize a total of 42 actions.
-Table 4 presents the names of the LLVM optimization passes corresponding to these 42 actions.
-Using this limited action space is a prevailing practice, and it has been proved to be effective in prior works(Haj-Ali et al.
-2020; Cummins et al.
-2022).
-It also makes the training easier since the model can focus on the most effective optimization passes.
+Table 9. A list of LLVM transformation passes selected in Autophase’s action space.
 Index Name Index Name Index Name
@@
 13 -inline 27 -loweratomic 41 -tailcallelim
-Table 4: A list of LLVM transformation passes selected in Autophase’s action space.
-B.3 Reward Function
-The reward function is defined as the normalized change of the optimization metric $C(s)$:
-$r_{t+1}=\frac{C(s_{t})-C(s_{t+1})}{C(s_{0})-C(s_{\mathrm{b}})},$ (9)
-where a lower $C$ indicates better performance.
-In line with other works focused on optimizing code size, we define $C(s)$ as the IR instruction count at compiler state $s$.
-The $C(s_{\mathrm{b}})$ stands for the performance of a baseline policy, specifically the built-in -Oz flag in our experiments.
-A total reward greater than 1 means that the optimization sequence performs better than the baseline policy.
-B.4 Hyperparameters
-The hyperparameters for our DreamerV3 agent implementation are outlined in Table 5.
-For hyperparameters not specified, we use the same value as the original DreamerV3 (Hafner et al.
-2023).
-Most of the listed hyperparameter values are directly taken from DreamerV2 (Hafner et al.
-2021) or previous works, with minimal tuning.
+B.3. Hyperparameters
+CompilerDream
+The hyperparameters for our world model and agent implementation are outlined in Table 10.
+For hyperparameters not specified, we use the same value as the DreamerV3 (Hafner et al., 2023).
+Most of the listed hyperparameter values are directly taken from DreamerV2 (Hafner et al., 2021) or previous works, with minimal tuning.
 However, the loss scales are carefully tuned, as the observation loss is relatively small compared to the reward loss in our setting, and the balance between these losses has a significant impact on CompilerDream’s performance.
 All experiments share the same set of hyperparameters unless otherwise specified.
+Model-based Baseline
+For MBPO in Section 4.5, the dynamics model is a 4-layer MLP with a hidden size of 1024.
+We adopt the implementation from https://github.com/Xingyu-Lin/mbpo_pytorch, with most hyperparameters left unchanged.
+The modified hyperparameters are listed in Table 12.
+Table 10. Hyperparameters in our experiments.
 Hyperparameter Value
@@
 Replay buffer capacity $2\times 10^{6}$
-Reward smoothing $\alpha$ (Lee, Abbeel, and Lee 2024) 0.6
+Reward smoothing $\alpha$ (Lee et al., 2024) 0.6
 Training frequency Every 5 environment steps
@@
 Gradient clipping 100
-Table 5: Hyperparameters of DreamerV3 in our experiments.
-B.5 Training Model-free Baselines
-We use RLlib (Liang et al.
-2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al.
-2017) (Haj-Ali et al.
-2020), A2C (Mnih et al.
-2016), IMPALA (Espeholt et al.
-2018), APEX (Horgan et al.
-2018), and DQN (Mnih et al.
-2015).
-All the experiments are conducted with 10 CPUs and an RTX-3090 GPU.
-For the PPO baseline in Section 4.4, we trained it for around 7 hours since we observed overfit to CodeContests when training for a longer time.
-Other model-free methods are trained for at least 10 hours.
-The training process uses 5 workers to interact with the environment and use 4 evaluation workers to evaluate the checkpoint on the validation set.
-We use default hyperparameters of algorithms in RLlib following the CompilerGym platform (Cummins et al.
-2022), except that we have carefully tuned the hyperparameters for PPO, as listed in Table 6.
+Model-free Baselines
+We use RLlib (Liang et al., 2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al., 2017) (Haj-Ali et al., 2020), A2C (Mnih et al., 2016), IMPALA (Espeholt et al., 2018), APEX (Horgan et al., 2018), and DQN (Mnih et al., 2015).
+We use default hyperparameters of algorithms in RLlib following the CompilerGym platform (Cummins et al., 2022), except that we have carefully tuned the hyperparameters for PPO, as listed in Table 11.
 We explored roughly 2 or 3 values for each hyperparameter listed and selected the set of hyperparameters that yielded the best performance on the validation set.
 The results of the PPO baseline in Section 4.5 and Figure 6 are exactly the same as those in Section 4.4 and Figure 4.
+Table 11. Hyperparameters for the PPO baseline, well-tuned on our dataset to be deviating from the default value in RLlib.
 Hyperparameters Value
@@
 weight_decay 1e-6
-Table 6: Hyperparameters for the PPO baseline, well-tuned on our dataset to be deviating from the default value in RLlib.
-B.6 Autotuning
-In Section 4.2, we extensively trained CompilerDream on the cBench dataset to evaluate its effectiveness as an autotuning method in comparison to other approaches on CompilerGym’s leaderboard.
-Similar to other studies, we trained CompilerDream using all programs in cBench, excluding the one called ghostscript due to its considerable size, which would significantly hinder the training process.
-The action space is set to the full set of 124 actions in LLVM.
-All other training details remained unchanged.
-B.7 Guided Search
-The guided search strategy we use in Section 4.2 followed the design in the PPO+Guided Search baseline.
-During the search, actions are sampled from the agent’s learned policy $\pi_{\psi}(a_{t}|s_{t})$ up to 45 steps, at which point a new episode begins.
-Code size reductions after applying each action are all recorded to calculate the maximum.
-To enhance exploration, we add a uniform distribution to the policy, implemented by a $5\%$ chance of selecting a uniformly random action instead of sampling from the policy.
-At each step, the code size reduction achieved so far is recorded for later comparison, and the elapsed time is monitored to ensure that the search is completed within the 1-minute time limit per benchmark.
-Compared to the guided search in PPO+Guided Search (in Table 1), we use a horizon of 45 steps instead of 200 steps.
-Additionally, PPO+Guided Search includes an extra search of 500 steps on the best sequence found, which we do not include, resulting in our wall time being a few seconds shorter than theirs.
-B.8 Value Estimation
-Here, we provide additional details about the experiment in Section 4.3.
-The baseline method, Coreset-NVP (Liang et al.
-2023), trains a value prediction model that scores sequences from a fixed set of optimization sequences, referred to as the core set, which is obtained through an extensive search on their training set.
-For each program to optimize, it enumerates the first 45 passes in the top-scored sequences and validates their effectiveness through actual compilation.
-Specifically, sequences are processed in descending order of scores, executing each pass within a sequence.
-The compiler state (i.e., the IR) resets to its initial state after all passes in a sequence have been enumerated, before moving on to the next sequence.
-The code size reduction achieved after each pass is recorded, and the pass, along with its corresponding sequence prefix that yields the highest code size reduction, is selected.
-The core set in Coreset-NVP contains 50 sequences comprising a total of 625 passes, resulting in an average sequence length of 12.5 passes.
-Thus, roughly 3 or 4 sequences are attempted per benchmark, given the limit of 45 passes.
-These passes correspond to the optimization actions defined in Section 3.1, with a total of 79 different optimization passes included.
-For our experiment, we set the action space to the full set of all 124 optimization passes available in LLVM when training our world model.
-Focusing on predicting the target value of cumulative reward after executing each optimization pass in the core set, we replace CompilerDream’s optimization agent with a policy that sequentially outputs the sequences from the core set, allowing the world model to focus on the problem dynamics driven by these sequences.
-This setup is also comparable to the baseline, as Coreset-NVP employs a value predictor that outputs a 50-dimension vector corresponding to the normalized value of each sequence in the core set.
-That prediction model is obtained through supervised training.
-We train our world model using the CodeContests dataset introduced in Section 3.4, consistent with the experiments in Section 4.4.
-During training, we increase the weight of the reward loss to 100.0, while keeping the other hyperparameters unchanged.
-To align with the methodology used in Coreset-NVP, we validate the top 45 sequence prefixes of each program with the highest predicted values through actual compilation and select the best-performing sequence.
-The world model is trained using three different random seeds, and the average results are reported in Table 2.
-B.9 Hardware and Training Time
+Table 12. Modified hyperparameters in our MBPO baseline.
+Hyperparameters Value
+MBPO target_update_interval 10
+lr 1e-3
+rollout_batch_size 50000
+rollout_max_length 45
+init_exploration_steps 3600
+B.4. Hardware and Training Time
+CompilerDream
 We train all CompilerDream-based methods with 64 CPUs and an RTX-3090 GPU.
@@
 In Sections 4.2 and 4.3, CompilerDream was trained for about 1 day.
+Random Search
 The random search baseline in Figure 4 is conducted with 4 CPUs, which is sufficient since it is a single-thread program.
-The random search baseline in Table 1 utilizes 80 CPUs, as specified in the write-up attached to CompilerGym’s leaderboard (Facebook 2022).
+The random search baseline in Table 4 utilizes 80 CPUs, as specified in the write-up attached to CompilerGym’s leaderboard (Facebook, 2022).
 Our CompilerDream + Guided Search method in Table 4 is tested on the same machine used for training CompilerDream, equipped with 64 CPUs and an RTX-3090 GPU.
-We ensure that sufficient memory is available for all experiments. Memory size should have little impact on the results in all of our experiments.
-B.10 Random Seeds
-All of our experiments that report a min-max range or standard deviation were replicated using three different random seeds.
-These seeds were randomly selected, and the results are generally reproducible with other seeds as well.
-Appendix C Benchmarks
-CompilerGym Benchmarks
-In our study, we select eight benchmarks for zero-shot test and in-domain training: BLAS, cBench, CHStone, Linux, MiBench, NPB, OpenCV, and TensorFlow.
-These benchmarks are part of the built-in datasets provided by CompilerGym version 0.2.5.
-For benchmarks with a total number of programs more than 100, we use the first 50 programs as the test set, the following 50 programs as the validation set, and all of the rest as the training set.
-These training and validation sets are only used for in-domain training.
-The datasets comprising fewer than 100 programs are not applicable for in-domain training; instead, all their programs are allocated to the test set.
-The number of programs in each dataset after division is detailed in Table 7.
-Dataset Training Split Validation Split Test Split
-BLAS 200 50 50
-cBench 23 N/A N/A
-CHStone 12 N/A N/A
-Linux 13,794 50 50
-MiBench 40 N/A N/A
-NPB 22 50 50
-OpenCV 342 50 50
-TensorFlow 1,885 50 50
-Table 7: Dataset division of 8 CompilerGym benchmarks.
-AI-Generated Benchmarks
-To further test the generalization ability of our CompilerDream agent on different programming languages, we borrow the method from FormAI (Tihanyi et al.
-2023) and generate a dataset containing 50 unique Objective-C programs using GPT-3.5.
+Model-free and Model-based baselines
+The hardware setting is the same as used for training CompilerDream.
+For the PPO baseline in Section 4.4, we train for about 7 hours, as longer training leads to overfitting on CodeContests.
+Other model-free methods in Section 4.4 are trained for at least 10 hours.
+We use 5 workers for environment interaction and 4 evaluation workers to assess checkpoints on the validation set.
+In Section 4.5, we train MBPO for approximately 10 hours, as its evaluation score converges within this time.
+B.5. Random Seeds
+All experiments reporting a min-max range or standard deviation are conducted with three different random seeds.
+The results in Table 5, Table 6, and Table 7 are also averaged over three runs with different seeds.
+The seeds are randomly selected, and the results are generally consistent across other random seeds as well.
+B.6. Detail of Guided Search in Autotuning
+Our guided search in Section 4.2 follows the PPO+Guided Search design.
+Actions are sampled from the learned policy for up to 45 steps per episode, recording code size reductions after each step to track the maximum.
+To encourage exploration, 5% of actions are sampled uniformly at random.
+We record the best reduction so far and monitor elapsed time to stay within the 1-minute budget per benchmark.
+Unlike PPO+Guided Search (Table 4), which uses a 200-step horizon and an extra 500-step search on the best sequence, we omit the latter and use only 45 steps, making our wall time slightly shorter.
+B.7. Detail of Value Prediction Experiment
+We provide additional details for the experiment in Section 4.3.
+The baseline, Coreset-NVP(Liang et al., 2023), trains a value model to score sequences from a fixed core set, obtained via extensive search on its training set.
+For each program, it evaluates the top-scored sequences in order, executing up to 45 passes across them.
+After each full sequence, the IR resets to the initial state.
+Code size reductions after each pass are tracked, and the best-performing pass and prefix are selected.
+The core set includes 50 sequences (625 passes total, 12.5 passes each on average), so around 3–4 sequences are tried per benchmark.
+The 79 passes used correspond to the optimization actions in Section 3.1.
+In our setup, the action space includes all 124 LLVM passes.
+To train the world model to predict cumulative reward values for the core set passes, we apply each core set sequence to training programs and collect trajectories.
+We use the CodeContests dataset (Section 3.4), consistent with the experiments in Section 4.4.
+Reward loss weight is set to 100.0, with other hyperparameters unchanged.
+For evaluation, we compile the top 45 predicted prefixes per program and report the best.
+B.8. AI-Generated Benchmarks
+To further test the generalization ability of our CompilerDream agent on different programming languages, we borrow the method from FormAI (Tihanyi et al., 2023) and generate a dataset containing 50 unique Objective-C programs using GPT-3.5.
 We use the same prompt as FormAI, except that we add an instruction to ask GPT to generate programs that can be directly compiled under Clang version 10.0.0 and do not use ARC (Automatic Reference Counting), to improve the compilation pass rate of generated programs.
 We compile the generated programs using Clang without including any third-party libraries, and all programs that cannot pass compilation are discarded.
-Appendix D Extended Experimental Results
+Appendix C Extended Experimental Results
+Table 13. Top performances of the zero-shot CompilerDream RL agent on individual programs.
 Dataset Program Pass Sequences by CompilerDream Agent Code size
@@
 TensorFlow #17 -sroa -gvn -early-cse -simplifycfg -early-cse -early-cse -instcombine -early-cse -sink -sink -sink -lowerinvoke -simplifycfg -sroa -early-cse -memcpyopt -early-cse -memcpyopt -instcombine -simplifycfg -early-cse -early-cse -early-cse -simplifycfg -early-cse -early-cse -strip -strip -indvars -simplifycfg -early-cse -memcpyopt -early-cse -instcombine -early-cse -early-cse -simplifycfg -early-cse -memcpyopt -early-cse -early-cse -instcombine -simplifycfg -early-cse -early-cse 5512 4247 5450 (1.28$\times$)
-Table 8: Top performances of the zero-shot CompilerDream agent on individual programs. Full IRs of these programs before and after optimizations are shown in the Supplementary Information.
-D.1 Learning Curves
-The zero-shot test performance of our CompilerDream agents during training is shown in Figure 7.
+C.1. Learning Curves
+The performance of our CompilerDream with RL agents during training is shown in Figure 8.
 Note that we validate and test the agent every 10000 environment steps and report the test performance from the checkpoint that achieved the best validation results for comparison among various methods.
-Figure 7: Zero-shot test performance of the CompilerDream agent during training. We report mean and standard deviation across three runs.
-D.2 Quantitative Results
-Quantitative results corresponding to Figure 3 in the main text are provided in Table 9.
-D.3 Program Case Study
-In Addition to the predicted optimization trajectory showcase in our paper, we present more example of optimization results here.
-Table 8 presents the agent’s top performance outcomes on various benchmark datasets.
-We observe that the agent indeed produces a specialized optimization strategy tailored for each program.
-Additionally, these results also highlight certain passes, such as -sroa (scalar replacement of aggregates), -early-cse(early common subexpression elimination) and -simplifycfg, as particularly effective in code size optimization.
-Figure 8: t-SNE (Van der Maaten and Hinton 2008) visualization of programs from training and test datasets.
-D.4 Data Distribution Visualization
-In Figure 8, we visualize the distribution of our training and test datasets.
-To accurately represent the dynamic behavior of programs, we randomly select 1000 action sequences, each with a length of 45, from our action space.
-These sequences are subsequently executed on each program, with the resulting Autophase features concatenated to form a feature vector with dimensions of $1000\times 45\times 56$ for every program.
-These comprehensive feature vectors are finally processed using t-SNE (Van der Maaten and Hinton 2008) for dimensionality reduction and visualization.
-Figure 8 illustrates that our training data (denoted as triangles) has a broad coverage of test programs (denoted as circles).
-This contrasts with the Csmith dataset (denoted as crosses) employed in CompilerGym experiments (Cummins et al.
-2022), which shows a significant deviation from real-world applications.
-Nonetheless, our visualization can still not perfectly capture the transferability across datasets.
-For instance, empirical evidence suggests that CodeContests are the most effective in generalizing to OpenCV and TensorFlow, while the visual analysis does not directly imply this.
+Figure 8. Zero-shot test performance of the CompilerDream agent during training. We report mean and standard deviation across three runs.
+[…]
+The learning curves for 7 out of the 8 datasets saturate quickly, while performance on the Linux dataset remains almost constant throughout the training process.
+Table 14. Quantitative results for code size reduction, corresponding to Figure 4. We report mean and standard deviation across three runs.
 -O0 Random Policy
@@
 max 1.010 1.235$\pm$0.034 1.267$\pm$0.007 1.281$\pm$0.006 1.289$\pm$0.008 1.282$\pm$0.002
-Table 9: Quantitative results for code size reduction, corresponding to Figure 3. We report mean and standard deviation across three runs.
+C.2. Quantitative Results
+Quantitative results corresponding to Figure 3 in the main text are provided in Table 14.
+C.3. Program Case Study
+In Addition to the predicted optimization trajectory showcased in our paper, we present more examples of optimization results here.
+Table 13 presents the RL agent’s top performance outcomes on various benchmark datasets.
+We observe that the agent indeed produces a specialized optimization strategy tailored for each program.
+Additionally, these results also highlight certain passes, such as -sroa (scalar replacement of aggregates), -early-cse(early common subexpression elimination), and -simplifycfg, as particularly effective in code size optimization.
+Figure 9. t-SNE (Van der Maaten and Hinton, 2008) visualization of programs from training and test datasets.
+[…]
+In each dataset, the majority of data points are concentrated within a few clusters, with a smaller proportion dispersed across various regions.
+Notably, the data distribution in the CodeContest dataset overlaps with that of most other datasets.
+C.4. Data Distribution Visualization
+In Figure 9, we visualize the distribution of our training and test datasets.
+To accurately represent the dynamic behavior of programs, we randomly select 1000 action sequences, each with a length of 45, from our action space.
+These sequences are subsequently executed on each program, with the resulting Autophase features concatenated to form a feature vector with dimensions of $1000\times 45\times 56$ for every program.
+These comprehensive feature vectors are finally processed using t-SNE (Van der Maaten and Hinton, 2008) for dimensionality reduction and visualization.
+Figure 9 illustrates that our training data (denoted as triangles) has a broad coverage of test programs (denoted as circles).
+This contrasts with the Csmith dataset (denoted as crosses) employed in CompilerGym experiments (Cummins et al., 2022), which shows a significant deviation from real-world applications.
+Nonetheless, our visualization can still not perfectly capture the transferability across datasets.
+For instance, empirical evidence suggests that CodeContests are the most effective in generalizing to OpenCV and TensorFlow, while the visual analysis does not directly imply this.

### v4

@@
 With strong generalization capabilities, it efficiently optimizes programs from diverse origins.
-[…]
 Diverse source code, including C++, C, Rust, and other languages, is first translated into an intermediate representation (IR) by the compiler frontend.
@@
 A range of machine learning applications (Devlin et al., 2019; Brown et al., 2020; Radford et al., 2021; Kirillov et al., 2023) have witnessed that training high-capacity models on large-scale datasets can yield unprecedented performance.
-Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al., 2022), yet the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al., 2020; Shahzad et al., 2022) or from relatively small training sets comprising only a few hundred programs with limited-capacity models (Mammadli et al., 2020; Jain et al., 2022), which hinders generalization.
+Prior studies on compiler optimization also suggest that training on large datasets with diverse programs could be beneficial (Cummins et al., 2022), yet the prevalent practice in this field is still to learn optimization strategies in a per-program manner (Haj-Ali et al., 2020b; Shahzad et al., 2022) or from relatively small training sets comprising only a few hundred programs with limited-capacity models (Mammadli et al., 2020; Jain et al., 2022), which hinders generalization.
 This paper focuses on the LLVM (Lattner and Adve, 2004) phase ordering problem, a longstanding challenge for compiler research.
@@
 Recent advancements have seen reinforcement learning (RL) techniques making strides in compiler optimization, circumventing the need for collecting optimal labeled data (Kulkarni and Cavazos, 2012).
-This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al., 2021), loop transformation (Haj-Ali et al., 2020; Brauckmann et al., 2024), and graph partitioning (Mirhoseini et al., 2017).
-Several works relevant to us, including AutoPhase (Haj-Ali et al., 2020), CORL (Mammadli et al., 2020), and POSET-RL (Jain et al., 2022), have explored the full optimization pipeline, i.e., the phase ordering problem, using model-free RL without incorporating a world model.
+This technique has been applied to optimize individual compilation heuristics, such as inlining (Trofin et al., 2021), loop transformation (Haj-Ali et al., 2020a; Brauckmann et al., 2024), and graph partitioning (Mirhoseini et al., 2017).
+Several works relevant to us, including AutoPhase (Haj-Ali et al., 2020b), CORL (Mammadli et al., 2020), and POSET-RL (Jain et al., 2022), have explored the full optimization pipeline, i.e., the phase ordering problem, using model-free RL without incorporating a world model.
 World model.
@@
 In preliminary experiments, we find that complex program features like ProGraML (Cummins et al., 2021) and inst2vec (Ben-Nun et al., 2018) significantly slow down CompilerDream with marginal performance gains.
-In contrast, expert-designed Autophase features (Haj-Ali et al., 2020) leverage domain knowledge, enhancing generalization by filtering irrelevant details.
+In contrast, expert-designed Autophase features (Haj-Ali et al., 2020b) leverage domain knowledge, enhancing generalization by filtering irrelevant details.
 We construct the observation by concatenating the 56-dimensional Autophase feature vector with a 42-dimensional action histogram vector, which records the number of times each action has been selected during the current episode.
@@
 We adopt two distinct action spaces in our experiments to align with the baseline methods.
-The full action space consists of all 124 LLVM optimization passes, while the reduced action space derived from Autophase (Haj-Ali et al., 2020) consists of 42 actions.
+The full action space consists of all 124 LLVM optimization passes, while the reduced action space derived from Autophase (Haj-Ali et al., 2020b) consists of 42 actions.
 Originally, this action space included 45 LLVM passes, but CompilerGym excludes 3 due to updates in the latest LLVM version, leaving 42 actions.
-This reduced action space is widely used and has been shown to be effective in prior studies (Haj-Ali et al., 2020; Cummins et al., 2022).
+This reduced action space is widely used and has been shown to be effective in prior studies (Haj-Ali et al., 2020b; Cummins et al., 2022).
 The reward function is defined as the normalized change of the optimization metric $C(s)$:
@@
 Autophase
-(Haj-Ali et al., 2020)
+(Haj-Ali et al., 2020b)
 A feature vector capturing
@@
 A reduced action space
-derived from Autophase(Haj-Ali et al., 2020).
+derived from Autophase(Haj-Ali et al., 2020b).
 42
@@
 CompilerDream supports training with RL and value prediction agents, which learn from episodes simulated by the world model.
-[…]
 The agent interacts with a compiler to collect real episodes, which are stored in a replay buffer.
@@
 Figure 3. Comparison of the original and smoothed rewards during an episode on a program from the BLAS dataset (Lawson et al., 1979).
-[…]
 Taking a typical episode as an example: out of 45 actions, only 8 receive non-zero rewards before smoothing, and they are scattered throughout the optimization process.
@@
 The value of a program state $s$ is defined as the expected cumulative reward of applying an action sequence to the program.
-This aligns with the ratio of the reduced IR instruction count achieved between applying action sequence $\{a_{\tau}\}_{\tau=1}^{m}$ and standard $-Oz$ flag, according to the reward function $r$ defined in Eq.
+This aligns with the ratio of the reduced IR instruction count achieved between applying action sequence $\{a_{\tau}\}_{\tau=1}^{m}$ and standard Oz-Oz flag, according to the reward function $r$ defined in Eq.
 (1):
@@
 Our RL agent comprises an actor and a critic neural networks, both parameterized based on the neural compiler state:
-[…]
+(8)  $\begin{split}&\text{ Actor: }\hat{a}_{t}\sim\pi_{\psi}\left(\hat{a}_{t}\mid\hat{z}_{t}\right)\\ &\text{ Critic: }v_{\xi}\left(\hat{z}_{t}\right)\approx\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum\nolimits_{\tau\geq t}\gamma^{\tau-t}\hat{r}_{\tau}\right].\end{split}$
 Figure 4.
@@
 Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
-[…]
 The performance of several methods (-O0, Random Policy, -Oz, Random Search, PPO, In-domain CompilerDream, and Zero-shot CompilerDream) on different benchmarks (BLAS, cBench, CHStone, Linux, MiBench, NPB, OpenCV and TensorFlow).
@@
 The actor produces an optimization policy $\pi_{\psi}$ that predicts an action distribution of the best pass to choose, which is trained to maximize the simulated return through the REINFORCE policy gradient (Williams, 1992) with an entropy regularization (Haarnoja et al., 2019):
-$(\displaystyle\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big{[}\sum^{t+H}_{\tau=t}\Big{(}$$-\left(V_{\tau}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}\mid\hat{z}_{\tau})$
+$\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big[\sum^{t+H}_{\tau=t}\Big($ $-\left(V_{\tau}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}\mid\hat{z}_{\tau})$
 (10) $-\eta\,\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\Big)\Big],$
@@
     + Guided Search
-60.8s 1.073$\times$
-PPO + Guided Search 69.8s 1.070$\times$
-CompilerDream 2.9s 1.068$\times$
-Random Search ($t$=10800) 10,512.4s 1.062$\times$
-Random Search ($t$=3600) 3,630.8s 1.061$\times$
-Greedy Search 169.2s 1.055$\times$
-GATv2 + DD-PPO 258.1s 1.047$\times$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 We first validate CompilerDream as an autotuning method to demonstrate its ability to discover high-quality optimization sequences, supporting our goal of achieving superior code optimization.
@@
 Results
-As shown in Table 4, our RL agent trained on world model simulations achieves an average 1.068×1.068$\times$1.068 × code size reduction on cBench in a single trial, surpassing random search methods despite their longer execution times, even when accounting for CompilerDream’s training time.
+As shown in Table 4, our RL agent trained on world model simulations achieves an average $1.068\times$ code size reduction on cBench in a single trial, surpassing random search methods despite their longer execution times, even when accounting for CompilerDream’s training time.
 This demonstrates the world model’s ability to accurately simulate optimization processes and support effective agent learning.
@@
 The learned world model accurately captures the variations of program features and optimization metrics.
-[…]
 This figure is the comparison between a ground-truth code optimization trajectory and an imagined trajectory by a learned compiler world model.
@@
 […]
-(a)The performance of CompilerDream steadily and consistently improves throughout training, and achieves a validation code size reduction of 1.02 after about $4\times 10^{4}$ environment steps.
-In contrast, other methods (PPO, A2C, IMPALA, DQN, and APEX) all achieve reductions of less than 0.98 at the same stage.
-To reach a reduction of 1, PPO needs nearly $2\times 10^{5}$ steps–roughly six times the steps required by CompilerDream.
-The remaining four methods exhibit unstable performance and fail to exceed a reduction of 1 even after $10^{6}$ steps.
-(b) The generalization capabilities of different RL algorithms (CompilerDream, PPO, A2C, IMPALA, DQN, and APEX) on five test datasets unseen during training (cBench, CHStone, NPB, OpenCV, and TensorFlow).
-CompilerDream outperforms all other methods across all five datasets, demonstrating a clear advantage.
-PPO consistently ranks second, though with a noticeable performance gap compared to CompilerDream.
-The remaining four methods perform poorly, particularly on the cBench and CHStone datasets.
-(c) The impact of the training dataset is demonstrated by comparing CompilerDream trained on two datasets: CodeContests and Csmith.
-Across all five benchmarks (cBench, CHStone, NPB, OpenCV, and TensorFlow), CompilerDream trained on CodeContests consistently outperforms its counterpart trained on Csmith.
-Notably, the performance gap even reaches a geometric mean code size reduction difference of 0.2 on the CHStone benchmark.
 Figure 7. Histograms of large-scale evaluations comparing CompilerDream and PPO on the FormAI dataset.
 […]
-The performance of CompilerDream and PPO was evaluated on the large-scale FormAI test.
-For both methods, the code size reduction for most test cases (approximately $5\times 10^{4}$) falls within the range of $0.8\textasciitilde{}1.2$.
-As the reduction magnitude increases or decreases beyond this range, the number of corresponding cases gradually decreases.
-Comparing CompilerDream with PPO, we observe that CompilerDream optimizes more cases with higher code size reduction while producing fewer cases with worsened results.
-Notably, CompilerDream achieves a reduction of more than nine times for some cases, surpassing all previous studies.
 Implementation
-We train CompilerDream with its RL agent, and compare it with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method (Haj-Ali et al., 2020) using Proximal Policy Optimization (PPO) (Schulman et al., 2017).
+We train CompilerDream with its RL agent, and compare it with the following baselines: a random pass sequence, an autotuning approach using random search, LLVM’s -O0 and -Oz flag, and a state-of-the-art learning-based method (Haj-Ali et al., 2020b) using Proximal Policy Optimization (PPO) (Schulman et al., 2017).
 The random search conducts hundreds of trials within a time budget similar to our RL agent.
@@
 The minimal performance differences across various methods on BLAS and Linux suggest these datasets are already highly optimized.
-Moreover, CompilerDream’s zero-shot generalization matches or surpasses in-domain training.
-This advantage is particularly evident in the NPB dataset, where data sparsity limits in-domain agents, yet CompilerDream achieves an additional 3% code size reduction.
-Its robust generalization extends to new languages, as shown by results on the Fortran-based BLAS and NPB datasets.
-On the AI-generated Objective-C dataset, CompilerDream achieves an average code size reduction of 1.027$\times$, reaching up to 2.87$\times$ in certain test cases.
+[…]
 On the large-scale FormAI test set (Figure 7), CompilerDream outperforms PPO, matching or outperforming -Oz on more programs and achieving higher optimization levels, demonstrating its superior and consistent performance.
@@
 CHStone 1.071 1.066 1.094
-MiBench 0.985 0.988 0.986
-NPB 1.000 0.996 1.006
-Linux 1.071 1.069 1.108
+Linux 0.985 0.988 0.986
+MiBench 1.000 0.996 1.006
+NPB 1.071 1.069 1.108
 OpenCV 1.082 0.998 1.092
@@
 Machine learning techniques introduced in Section 2 require crafting high-quality features that capture the important characteristics of programs, a process known as feature engineering.
-The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni et al., 2014; Haj-Ali et al., 2020), designed by expert intuitions.
+The most prevalent feature vectors are based on the frequencies of various types of instructions or patterns within the programs (Magni et al., 2014; Haj-Ali et al., 2020b), designed by expert intuitions.
 Numerous studies have aimed to reduce the cost of feature design.
@@
 B.2. Details of Features and Actions
-As shown in Table 1, we adpot the 56-dimensional Autophase feature and a reduced action space from Autophase (Haj-Ali et al., 2020).
+As shown in Table 1, we adpot the 56-dimensional Autophase feature and a reduced action space from Autophase (Haj-Ali et al., 2020b).
 A detailed list of the Autophase features and actions space is available in Table 8 and Table 9.
@@
 Model-free Baselines
-We use RLlib (Liang et al., 2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al., 2017) (Haj-Ali et al., 2020), A2C (Mnih et al., 2016), IMPALA (Espeholt et al., 2018), APEX (Horgan et al., 2018), and DQN (Mnih et al., 2015).
+We use RLlib (Liang et al., 2018) to train and test model-free reinforcement learning algorithms including PPO (Schulman et al., 2017) (Haj-Ali et al., 2020b), A2C (Mnih et al., 2016), IMPALA (Espeholt et al., 2018), APEX (Horgan et al., 2018), and DQN (Mnih et al., 2015).
 We use default hyperparameters of algorithms in RLlib following the CompilerGym platform (Cummins et al., 2022), except that we have carefully tuned the hyperparameters for PPO, as listed in Table 11.
@@
 O0 CompilerDream Oz (Reduction)
-cBench sha -sroa -gvn -sroa -sroa -instcombine -simplifycfg -sroa -functionattrs -licm -functionattrs -early-cse -simplifycfg -indvars -gvn -simplifycfg -early-cse -functionattrs -reassociate -memcpyopt -gvn -simplifycfg -reassociate -functionattrs -early-cse -instcombine -early-cse -lowerinvoke -simplifycfg -functionattrs -functionattrs -memcpyopt -gvn -simplifycfg -functionattrs -reassociate -early-cse -instcombine -early-cse -strip -strip -early-cse -strip -early-cse -strip -early-cse 799 349 500 (1.43$\times$)
-bzip2 -sroa -gvn -sroa -sroa -simplifycfg -strip -strip -instcombine -simplifycfg -early-cse -simplifycfg -strip -strip -reassociate -reassociate -memcpyopt -jump-threading -memcpyopt -functionattrs -memcpyopt -functionattrs -licm -functionattrs -gvn -simplifycfg -functionattrs -functionattrs -reassociate -memcpyopt -early-cse -jump-threading -memcpyopt -early-cse -strip -strip -instcombine -simplifycfg -jump-threading -jump-threading -early-cse -jump-threading -early-cse -simplifycfg -strip -functionattrs 28748 13565 15946 (1.18$\times$)
-OpenCV #41 -sroa -early-cse -early-cse -simplifycfg -early-cse -early-cse -early-cse -sink -sink -sink -loop-deletion -loop-deletion -loop-deletion -lowerinvoke -simplifycfg -sroa -sroa -sroa -early-cse -early-cse -early-cse -early-cse -lowerinvoke -lowerinvoke -lowerinvoke -globalopt -sroa -instcombine -simplifycfg -early-cse -early-cse -early-cse -sink -globalopt -sroa -sroa -functionattrs -simplifycfg -early-cse -functionattrs -early-cse -sink -functionattrs -functionattrs -functionattrs 28 18 28 (1.56$\times$)
-#9 -sroa -gvn -sroa -early-cse -simplifycfg -strip -strip -loop-deletion -lowerinvoke -simplifycfg -sroa -early-cse -instcombine -simplifycfg -early-cse -memcpyopt -early-cse -strip -strip -strip -functionattrs -instcombine -simplifycfg -jump-threading -simplifycfg -memcpyopt -early-cse -simplifycfg -reassociate -memcpyopt -functionattrs -functionattrs -reassociate -memcpyopt -early-cse -instcombine -simplifycfg -jump-threading -early-cse -jump-threading -strip -simplifycfg -early-cse -simplifycfg -functionattrs 9510 6341 9269 (1.18$\times$)
-TensorFlow #17 -sroa -gvn -early-cse -simplifycfg -early-cse -early-cse -instcombine -early-cse -sink -sink -sink -lowerinvoke -simplifycfg -sroa -early-cse -memcpyopt -early-cse -memcpyopt -instcombine -simplifycfg -early-cse -early-cse -early-cse -simplifycfg -early-cse -early-cse -strip -strip -indvars -simplifycfg -early-cse -memcpyopt -early-cse -instcombine -early-cse -early-cse -simplifycfg -early-cse -memcpyopt -early-cse -early-cse -instcombine -simplifycfg -early-cse -early-cse 5512 4247 5450 (1.28$\times$)
+[…]
+[…]
+[…]
+[…]
+[…]
 C.1. Learning Curves
@@
 Figure 8. Zero-shot test performance of the CompilerDream agent during training. We report mean and standard deviation across three runs.
-[…]
 The learning curves for 7 out of the 8 datasets saturate quickly, while performance on the Linux dataset remains almost constant throughout the training process.
@@
 (Zero-shot)
-BLAS geomean 0.931 0.821$\pm$0.018 0.960$\pm$0.004 0.993$\pm$0.002 0.993$\pm$0.005 0.991$\pm$0.000
-min 0.707 0.432$\pm$0.106 0.723$\pm$0.076 0.916$\pm$0.004 0.831$\pm$0.148 0.913$\pm$0.000
-max 1.000 1.002$\pm$0.004 1.031$\pm$0.010 1.014$\pm$0.001 1.025$\pm$0.005 1.016$\pm$0.000
-cBench geomean 0.481 0.737$\pm$0.015 0.858$\pm$0.020 1.010$\pm$0.004 N/A 1.036$\pm$0.002
-min 0.299 0.401$\pm$0.097 0.560$\pm$0.030 0.817$\pm$0.109 N/A 0.962$\pm$0.001
-max 0.626 1.150$\pm$0.113 1.298$\pm$0.083 1.405$\pm$0.014 N/A 1.395$\pm$0.054
-CHStone geomean 0.487 0.701$\pm$0.043 1.037$\pm$0.006 1.071$\pm$0.002 N/A 1.094$\pm$0.005
-min 0.402 0.278$\pm$0.108 0.889$\pm$0.044 0.967$\pm$0.027 N/A 1.000$\pm$0.006
-max 0.655 1.038$\pm$0.033 1.304$\pm$0.038 1.313$\pm$0.045 N/A 1.378$\pm$0.021
-Linux geomean 0.988 0.981$\pm$0.004 1.001$\pm$0.000 0.985$\pm$0.001 0.993$\pm$0.000 0.986$\pm$0.005
-min 0.615 0.740$\pm$0.126 1.000$\pm$0.000 0.615$\pm$0.000 0.700$\pm$0.001 0.642$\pm$0.041
-max 1.011 1.001$\pm$0.002 1.011$\pm$0.000 1.009$\pm$0.003 1.011$\pm$0.000 1.009$\pm$0.003
-Mibench geomean 0.389 0.812$\pm$0.022 1.005$\pm$0.001 1.000$\pm$0.007 N/A 1.006$\pm$0.002
-min 0.278 0.352$\pm$0.074 0.858$\pm$0.010 0.833$\pm$0.062 N/A 0.879$\pm$0.003
-max 0.760 1.392$\pm$0.074 1.603$\pm$0.015 1.524$\pm$0.045 N/A 1.588$\pm$0.000
-NPB geomean 0.530 0.833$\pm$0.011 1.074$\pm$0.010 1.071$\pm$0.018 1.075$\pm$0.017 1.108$\pm$0.001
-min 0.191 0.414$\pm$0.043 0.805$\pm$0.104 0.801$\pm$0.070 0.810$\pm$0.082 0.886$\pm$0.000
-max 1.066 1.848$\pm$0.193 2.141$\pm$0.062 2.235$\pm$0.101 2.315$\pm$0.028 2.343$\pm$0.012
-OpenCV geomean 0.981 0.949$\pm$0.011 1.080$\pm$0.001 1.082$\pm$0.004 1.087$\pm$0.007 1.092$\pm$0.000
-min 0.833 0.603$\pm$0.133 0.888$\pm$0.021 0.897$\pm$0.002 0.898$\pm$0.001 0.897$\pm$0.000
-max 1.370 1.409$\pm$0.167 1.571$\pm$0.022 1.620$\pm$0.193 1.635$\pm$0.186 1.556$\pm$0.000
-TensorFlow geomean 0.983 0.912$\pm$0.009 1.006$\pm$0.003 1.010$\pm$0.019 1.032$\pm$0.000 1.032$\pm$0.001
-min 0.927 0.618$\pm$0.046 0.877$\pm$0.009 0.841±0plus-or-minus0$\pm$ 0± 0.134 0.970$\pm$0.010 0.968$\pm$0.011
-max 1.010 1.235$\pm$0.034 1.267$\pm$0.007 1.281$\pm$0.006 1.289$\pm$0.008 1.282$\pm$0.002
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 C.2. Quantitative Results
@@
 Figure 9. t-SNE (Van der Maaten and Hinton, 2008) visualization of programs from training and test datasets.
-[…]
 In each dataset, the majority of data points are concentrated within a few clusters, with a smaller proportion dispersed across various regions.